A Complete Bibliography of Publications in Computer Physics Communications: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.10

Title word cross-reference

(a, a′) [SR73, SR73]. (a, b) [SMD71, SR73]. (a, bγ) [SR73]. (a, bγ − γ) [SR73]. (a, bσ) [SMD71]. (a, bσ − σ) [SMD71]. (a, γ) [SR73]. (a, γ − γ) [SR73]. (η, ρ) [BDG^{+}72]. (np)^{q} [CLS^{+}70]. + [CL79]. 1 [RCL75]. 1/2 [Hei72, KF71, SDH74]. $15.00 [Sut72]. $15.20 [Chr75a, Gos78, Mor78b]. 2 [BPW71, Bun76]. 3 [Bun76, CSC74]. $35.95/Dfl [Hib79]. 3j [SG76]. 3n [Sha70]. 6j [SG76]. $7.70 [Cam72]. = [CL79]. ^{+} [BS77]. _{2} [CL79, DST75, Mag75, RS76]. A [Kac77]. α [KPH78]. b_{n} [BS77]. d [Chi73b]. η [BFSG74]. f [AM74]. F_{0}(x) [SC71]. h [CMS77]. H_{n}^{(1)}(z) [AM79]. H_{n}^{(2)}(z) [AM79]. Hx = eSx [Nas74]. i [Hub70a]. I = 32 [DG75]. J [Cha74, Cha74, Rei79, Hub70a, Hub70b, Sha70]. J_{0} [AM77]. J_{1} [AM77]. j_{l}(r) [Del79b]. j_{n} [AM78]. J_{n}(z) [AM79]. J_{ν}(x) [Cam79]. jj [Gra72, PGB78]. L^{2} [Rei73, Rei79]. λ [Chi72]. LU [KPP79]. N [KK71, CN73, Mor78a]. O(3) [MPS75]. P_{L}^{M}(x) [Del79a]. R [BBC^{+}74, BBL^{+}78, BBL79, Bur73, Cla78, LWSS79, Rob72]. R^{k} [Lew70a]. ρ [BFSG74]. S [Cug73]. S = 1/2 [MGTD73]. S_{n} [SM78c]. SU(2) [Sha70]. SU(3) [Bra78a, Bra78b, CN73]. SU_{3} [AD73]. → [CL79, vdM72]. U(3) [MPS75]. X [EG74]. Y [EG74]. Y_{2}^{0} [Hir73]. Y_{4}^{0} [Hir73]. y_{n} [AM78]. Y_{n}(z) [AM79]. Y_{ν}(x) [Cam79]. Z [Lew70b, Lew79].

-coupling [Gra72, PGB78]. -D [CSC74]. -expansion [Lew70b, Lew79]. -matrices [Chi72]. -matrix [BBC^{+}74, BBL^{+}78, BBL79, Bur73, Cla78, LWSS79, Rei79, Rob72]. -particle [Mor78a]. -spin [Hub70a]. -stability [CMS77]. -state [Cug73].

01 [BPW72c].

1 [CRL78, Mag78]. 10.25 [Bra72c]. 10.50/$22.00 [Del77]. 10.75 [McN71]. 11-15 [Ano78t]. 12.15 [Bee73]. 15 [DTE74]. 165 [HRR75]. 17.20/$29.00 [Gra78]. 18.70 [Nor73b]. 1975 [Ano74d, Ano74e]. 1977 [Ano76h]. 1978 [Ano79f]. 1p [MNJL71]. 1p-shell [MNJL71].

2 [CHW^{+}78, CW79b]. 2.15s [Hoa70]. 2.50 [Pat73]. 22 [Smi77].

370 [HRR75]. 370/165 [HRR75].

5.50 [May74, Moo72, War73]. 50 [RR74].

6.75 [Smi72]. 6.95 [Lai74]. 6000 [Str74]. 6000/7000 [Str74]. 6500 [HRL75].

7.00 [Zac74]. 7000 [Str74].

8.15 [Bur72a].

9.50 [AF72b]. 90.00 [Hib79].

A-Three [Aue78]. AAKF [FS75b]. AAKL [Sax77b]. AAKP [Sax77a, Sax78]. Abel [FS75a, FC74]. absorption [All72, CHMM73, Mor72]. Academic [Bra72c, Bur72a, McN71, Smi72, War73]. accelerator [Kon78b]. according [Wil70]. accuracy [TS75b]. Accurate [AM79, CS78]. achieved [BH72]. ACQN [NA72]. acquisition [Rus73]. ACRL [BTS74]. acrz [God78]. Activities [Ano75c]. Adams [Moo79b]. Adaptation [CT79, GH76, Hib74b, Hib74a, MRW76, Sax77a, Sax78, Deu79c, God78, Hub70d, MW70, MW71a, MW71b, MW71c, MT79]. adapted [BTS74, FS75b, NA72, Sax77b]. adaptive [Gen72]. additive [Joh73]. Adiabatic [RCL75, Hen75, NW79]. adjuster [Han78f]. Adriatic [Tom73]. advance [Bun76]. adverse [LM71]. afterglows [LNS76]. aid [FC71, Zhi72]. aided [KS78]. air [Goo74, Sta72]. Alan [Del77]. Alder [Bra72c, McN71]. Alfven [WBG79]. algebra [Jon74]. Algebraic [Vel72, AT75, BH79, Ner78, Str74]. algebras [Bos76a, Bos76b]. Algol [KPPR76, Jab73, KPPR76, SMD71]. algorithm [BB70, Bos76a, Chi72, Klu72b, PW73, Tob79]. Algorithms [Cha76]. aligned [RSL^{+}78]. alkali [Moo71c]. alkali-like [Moo71c]. all-purpose [vM76a, vM79a]. Allowed [Hub70a]. allowing [KK75]. alloys [BP77]. alpha [SDH74]. ALTRAN [FP73]. amorphous [KH78, RM76a, RM76b, RM76c]. amplification [RS76]. amplitude [CSP^{+}79]. Amplitudes [Moo71c, GPS75, PS76, PS77, Rei73]. analyser [CHMM73]. analysing [KE72, KE74, VDM78, War78, Zhi72]. Analysis [BB74, BGM79, Cro72, HEL71, Sal75, AK76, AWG71, BJ76, BCV^{+}72, CMS77, CT71, De 75a, GCCB73, JR75, KN73, NM73, PS78, RR76, Sut72, VG74, WW74, WW72, Zlo79, vMG73, vM75c, vM75a, vM78a, vM79b, Hib79]. Analytic [CN73, CSP^{+}79, CC79, Goo74, Tro73, Lab79c]. Analytical [TYCT73]. analyticity [Nut73]. analyze [AZS^{+}79]. angle [Deu79a, Deu79b, Deu79c, Deu79d, SK79]. Angular [TTW79b, Tam70, BRT78, BGMP79, Gra73a, Gra76b, Hib70, Hib71, Hib74b, Hib74a, Hub70a, Lab79b, Lan70, RV78, RSL^{+}78, SDH74, TTW79a, Wil71]. anisotropic [LB76, Pet76a, ST72]. Anna [AK76]. annotation [Bur70b]. Announcement [Ano73a, Ano74a, Ano76a, Ano79a, Ano79b]. anomalous [TT78]. antenna [FM79]. Application [Ano72a, FC74, KKN^{+}72, KSSA78, Wil72, BPZ73, BS75a, EJ77, Ner78, TYCT73]. Applications [BH79, Sou73, BP73, Cam70, Dem72, Kil79, Nad76, PB73, TA72, War73]. Applied [Mor78b, TSD76]. Approach [Cal72, BGT76, Yos78, Gra78]. approximant [Bur72a]. approximants [GMR75a, GMR75b, Sta76, War73]. approximate [CE75, Ort72]. Approximation [Jab73, ATV76, Bok78, Cha75, GPS75, Hen75, NK78, PS76, PS77, TCR71, Lab79d]. approximations [Lab79c]. Arbitrary [Sha70, Deu79c, Deu79d, Hub71, KK77]. architecture [Pen78]. arcs [FCW78, SF78]. area [BS74]. Argument [AM77, Köl72a, AM78, AM79, Cam79]. arising [Jes79]. AROSA [RR76]. AROVMI [AF72a, AF74]. array [Nas74]. artificial [BD72]. aspects [AAB75, GJ76]. assignments [SDH74]. assistance [Sut72]. assisted [AWG71]. Associated [Bra73, Del79a]. astronomical [BCV^{+}72]. astronomy [Sch78, Wil72]. astrophysics [Tho72]. asymmetry [DG75, PI78]. asymptotic [Chi73a, Nor73a]. asymptotically [AAJ76]. Athene [CRL78]. atom [AT75, AV78, CSW78, DT79, Joa73, NM73, Nes79, Rei73, Rin79, ST73, TYCT73, Yat71]. atom-molecule [DT79]. Atomic [Ano79f, Bra72c, Moo72, MR73, BBC^{+}74, BBL^{+}78, BB77, Bur72c, Bur73, CCR79, CLS^{+}70, Cro72, DSK72, DSK73, DHN79, DH71, EJN74, Fro73, GPS75, Gra73a, Gra76b, Gra79, Hib70, Hib71, Hib74b, Hib74a, Kla71, Lab79a, Lab79b, Lew70b, Moo71c, Nut73, PS76, PS77, RFSG72a, Rei79, Rob72, WB74]. Atomique [Ano75c]. atoms [Aur77, BBH^{+}77, CCR76, GH78, Gra78, Lew79, LCW71, Mal70, Nus70, Per73, Sil78a, Sil78b, Wil78, WS79]. Author [Ano72b, Ano72c, Ano72d, Ano73b, Ano73c, Ano74b, Ano74c, Ano75b, Ano75a, Ano76b, Ano76c, Ano78a, Ano78b, Ano78c, Ano79c, Ano79d, Ano79e, Ano70h, Ano71d]. authors [Ano74o, Ano76m, Ano77i]. Automatic [EGMS72, PPR72, VG74, vMG73, BCV^{+}72, BEK^{+}72, CMW74, vM75c]. automation [Pen78]. averaging [vM79b]. axial [PI78]. axially [Cam75b, DEW78, Ugn78]. axisymmetric [KS73, KS75c].

B [Bra72c, McN71]. Back [Vio72]. backward [MT79]. bag [CH74]. Baker [Bur72a]. Band [Gru75, Hof71, HM71a, HM71b, HM71c]. bands [HRB72, MW70, MW71a, MW71b, MW71c, WC78]. bandstructure [De 75b]. bank [GCCB73]. Barbara [SMD71]. base [Rob79c]. based [Hof74, KK77, Kon78b, Mat72, WBG79]. bases [MPS75]. basic [CW79a, Sut72]. basis [AT75, AN79, Rei73, ST73]. batch [CMW74]. BATMAN [CMW74]. BCC [MR77]. be [CSP^{+}79, dJ78]. beam [IL74]. beam-foil [IL74]. beams [LL72]. Beeler [Nor73b]. behaviour [CIT73, CL79, Wes76]. Berztiss [Smi72]. Bessel [AM77, AM78, AM79, Cam79, Del79b, SZ79]. between [BGMP79, DEW78, DF76, FT71, Gia76, Jam70, Mal70, Per73]. Bicubic [Cha75, CM79]. binary [BP77, Gol70, MR77]. binding [Cug73]. biological [Cro72]. BNDPKG [WC78]. Board [Ano71a, Ano72g, Ano72h, Ano72i, Ano73f, Ano73g, Ano74h, Ano74i, Ano75g, Ano75h, Ano76i, Ano76j, Ano77b, Ano78i, Ano78j, Ano78k, Ano79j, Ano79k]. body [Ahl79, Ass75, Bra72b, Bra72a, HIM73, Kel79, Mue71, Sil78a, Sil78b, TYCT73, Wil78, WS79]. Boltz [TSD76]. bombardment [MM74, MRW76]. bond [NW79]. Book [AF72b, Ano79f, Bee73, Bra72c, Bur72a, Cam72, Chr75a, Del77, Gos78, Gra78, Hib79, Hoa70, Lai74, May74, McN71, Moo72, Mor78b, Nor73b, Pat73, Smi72, Smi77, Sut72, War73, Zac74]. Born [Hol79, Moi73, NK78, TCR71]. Bound [DHD74, LR71, Bat74, DeV76, Fro73, LW75, SS74, SS76]. boundary [Cam75b, De 77, O’C76, RD79]. brackets [Dob79, FT75, LJ71, SG72]. Brandt [Hib79]. Breit [BGMP79]. Bremsstrahlung [BRT78]. brief [Fos72]. Brody [LJ71]. bubble [BSP71, BD72, BEK^{+}72]. bulk [Mor72]. Butterworths [Hoa70].

C [Hoa70, Nor73b]. C.P.C. [Ano70b]. CACI [Smi77]. caesium [ÖW70]. Cain [Zac74]. Calcomp [CA70]. Calcul [Ano75c]. calculate [BGMP79, BBC^{+}74, BBL^{+}78, BTS74, Bur70c, DHN79, DH71, Duf77, EM74, EM76, God78, Gra73a, Gra73b, Gra76b, Hib71, Hib74b, Hib74a, Hib75, Jam70, KF71, Klo75a, MT79, NA72, PV74, RBHW76, SM78a, Ste79, TSD76, dJ78]. calculated [Gue74, Hen75]. calculating [BRT78, Bat74, BTS73, FT71, Gia76, Gru79, Hib70, Hof71, HRB72, Hof74, Isl73, Jon74, KM74, MRCL76, MW70, MW71a, MW71b, MW71c, MS78, MT76, PGB78, RR74, RSL^{+}78, Sei73, Tal76, Yat71]. Calculation [BM76, BCE79, Cal72, CHR75b, GMR75a, JP78, KM79b, LLC76, MRW76, Mik74, Moo72, PS74, Pie73, Pre79a, RFSG72a, Sal78, SS71, AN79, AAJ76, All72, And70, BKE79, CLS^{+}70, Cop74, Cop75a, EJN74, FK73, Fer75, FFH76, GPS75, GMR75b, Han78c, Han78e, Hol79, JPB78, Kac77, Mar73, PS76, PS77, RFSG72b, Ran72, Rei73, RS74, RS75, Sta76, TT78, WC78, ZDN78, vdB74]. Calculations [Dob79, FT75, AT75, AFH71, AF72a, AF74, Ahl79, AAH72, Aue78, BC79, Bet75, BP75, BBL79, CB78, Con79, Cop75b, Des70, Dic79, Ell75, FB75, FKU^{+}76, Gra79, Han78a, Jad75, Kel79, KK77, Ner78, Nes79, Net75, Nut73, RR76, Rob72, TL74, Yos78, vdM72]. calibration [Wil72]. CAMAC [Kol78, Taf75b]. Cambridge [Pat73]. Canterbury [GMR75a]. cards [Fem71]. Carla [SM78a]. Carlo [Chr74b, Cop74, Cop75a, Cop75b, DS75, Jad75, Kas75, KVW70, Mor78a, MR77, Per73, Ran72, RR74, Sta72, SK79, Wil72]. CARS [SLH78]. cascade [AV78]. cascades [Ran72, RR74, SV79]. case [GPS75, PS77, RM76a, RM76b, RM76c]. Castor [CW79b]. catalogue [Haw73]. cavity [Kon78b]. CDC [HRL75, LS79, Str74]. cell [JP78]. center [Yur78]. Centre [Ano75c]. certain [Hol79]. CFP [Bra78b]. CFPJJ [Gra72]. CFPJJ-Fractional [Gra72]. chain [KSSA78]. chamber [BSP71, BD72, BEK^{+}72, EGMS72]. chambers [BCV^{+}72]. change [Fem71]. changes [KM79a]. channel [HLS75, NK78, SDH74]. channel-spin- [SDH74]. channels [Com78, RSA74]. characteristics [Dem72, Dem78]. charge [DE72, Mal70, NW79, vdB74]. charge-exchange [vdB74]. charged [CJ71]. charges [Ano74d, Ano74e, Ano76h]. Charlotte [Gra78]. Chebyshev [Del79a, Del79b, She74, She76]. checking [Hib74a, Sax78]. checks [NM73]. chemical [OBLR73, Rob79a, Rob79b, Rob79c]. chemistry [FH74, Ner78]. chief [Hoa70]. chloride [ÖW70]. CINDY [SR73]. CIV3 [GH76, Hib75]. Classical [BPW71, BPW72a, BPW72b, BBH^{+}77, CMS77, Per73]. Clebsch [RB75, SM78c]. Clebsch-Gordan [SM78c]. Close [Sea73]. closed [CW79a, Sin70, Tro73]. clouds [BS77]. cluster [Mue71]. clusters [KM79b]. CO [DST75, Mag75, RS76]. COCHASE [HLS75]. code [Aue78, CHW^{+}78, CAR74, CRL78, CW79b, CH74, DTE74, DST75, DH75, ETH73, ET76, Fem71, FB75, GA74, HLS75, Kon78b, LNS76, NK78, RR76, RSA74, SM78a, Tal76, TSD76, WBG79, vM75c, FSR78]. codes [DSK72, DSK73, HW77, Rin79]. codes-V103 [HW77]. Codnum [Fem71]. coefficient [Bur70c, Gra73b]. Coefficients [Hub70b, RB75, SG76, AD73, AM74, BGMP79, CN73, Chi73b, DHN79, DW72, DH71, Gra72, Gra73a, Gra76b, JS72, Joh73, LB76, MRCL76, Nus70, PR75, RV78, Rob79c, RSL^{+}78, SM78c, SF76, Tam70, TS75a, TSD76, Tob79, Wil71, Zoh72]. coherent [SDH74]. cold [BS77, FM79, SR78]. Colin [Pat73]. collect [vM78b]. colliding [MS78]. collinear [CL79]. Collision [Moo72, Sar70, BKE79, BC79, Cla78, Con79, CSW78, Dic79, Hof74, Hub73, McC75, NM73, NA72, OS71, Sal78, Yat71]. collisions [BBH^{+}77, GPS75, JB77, Joa73, KVW70, LWSS79, Par73, Per73, PS76, PS77]. Collrad [Tal76]. Colour [BP75]. column [TT78]. combination [Wil70]. Combinatorial [BB77]. Comment [Pay76]. committee [Ano72n]. committees [Ano76o]. common [Klu72b]. Communications [Ano74d, Ano74e, Ano75d, Ano76d, Ano77a, Ano78d, Ano78e, AF72b]. compact [KPH78, Nas74]. compared [CT79]. comparing [ABMT78]. comparison [AMB75, GCCB73, ZKU79]. Compilation [Haw73]. Complete [Ell75, Han78a, KK77]. Complex [AM77, Köl72a, TTW79a, TTW79b, AM79, KF71, Nus70, SF76, dJ78]. component [CMS77, O’C76, Sta72]. compound [SM78a, SMD71, SR73, SDH74]. compound-nuclear [SDH74]. compounds [KN79b, KN79a]. comprising [RD79]. Compton [SS76]. Computation [BS74, BS75b, BS77, Cug73, Deu79b, Deu79d, HIM73, Hei72, JVS72, Kno70, Köl72b, Lac76, LJ71, LB76, Mal70, MBW76, Pay72, SMD71, SR73, SR78, AL79, BGT76, BKK76, Blo72, BB77, BH79, CS78, DG75, DE72, Del77, Dem78, FCW78, Han78c, JS72, Joh73, Klu72a, Lab79a, Pet76a, RD79, Ste76, Zoh72]. Computational [Ano79f, Bra72c, BH72, DT79, GJ76, GA74, Hib79, McN71, Ned72, RC72, Ano79f, Bee73, Lai74, May74]. computations [KS75a, Kla71, Lab79c, MR77]. compute [Chi73a, Chi73b, WB74]. computed [Klu72b, vdSVSH73]. Computer [AF72b, AWG71, BJ76, BK72, BS75a, CT71, Chr74a, DBB71, Gro72, KS78, KM79a, MS72, MGTD73, PW73, Pen78, Rap74, Ryc78, Sch73, Sch72, ST72, Tho72, Tor73, WD77, WP73, BB74, CHW^{+}78, CLS^{+}70, Daw72, EJ77, EM74, EM76, Gio79, Han78a, Han78b, Han78c, Han78e, Haw73, JB77, Kac77, LL72, LB78, PR75, Rob71, RM76a, RM76b, RM76c, Tom73, WDM77, Zac72, Zhi72, Ano74d, Ano74e, Ano75d, Ano76d, Ano77a, Ano78d, Ano78e, Sut72, Hoa70]. Computer-assisted [AWG71]. Computer-drawn [BS75a]. computer-simulated [WDM77]. computers [BTS74, BS75a, Bur72c, Cas72, Cha72, DTE74, ETH73, NA72, Pas72, Sou73, Taf75a, Web73, vM75c]. Computing [Aig72, Bos76b, Köl72a, Mor78b, O’C76, Han78d, Haz72, McC75, Nad75, Rou72, Tel73, TW79, Vio72]. concepts [BPZ73]. concerning [Bou72]. conditioned [TA72]. conditions [Cam75b, De 77, LM71]. Condon [Tob79]. conduction [Fos72]. conductivity [RM76a, RM76b, RM76c]. Conference [Ano79f, Cam72, Chr75a, Ano78p, Gos78]. configuration [AN79, BCE79, CLS^{+}70, Fro70, Fro72, Fro78, Hib74a, Hib75]. configurations [BDD78, Cha74]. confinement [BMS72]. conjugate [KPP79]. connected [Rap74]. considerations [Lab79b]. consistency [NM73]. consistent [CCR76, Des70, LCW71]. constant [Han78f]. constants [BCE79, JP78]. Construction [HM71c]. contained [BM76]. containing [MR73, Rob79c]. containment [CHW^{+}78, CHR75b]. Content [Ano76e, Rob79a]. Contents [Ano70a, Ano72e, Ano72f, Ano73d, Ano73e, Ano74f, Ano74g, Ano75f, Ano75e, Ano76f, Ano78f, Ano78g, Ano78h, Ano79g, Ano79h, Ano79i, Ano71e]. continua [Rei79]. continuity [Bor76]. continuous [CCR79]. Continuum [AT75, Bat74, BBC^{+}74, BBL^{+}78, BFKD77, BS75b, DF76]. Contribution [Daw72, KS75a]. contributions [Mue71, TS75a]. control [BP73, CR74, HRL75, KM79a, Rus73]. controlled [TS75b]. Convection [Gra76a, PW72, MPW73]. conventions [BPW72c]. convergence [CH75, Hub70d]. convergent [DESL79]. Conversion [Gol70, DHN79, DH71, PR75]. Converter [Sal76]. Cooley [Gue74]. coordinate [BP75, CB77, De 75a]. coordinates [Han78a, Hug71, MR73]. core [Bat74, CCR79, CHR75b, Hir73, HH75, LS79]. Coriolis [Kac77]. coronal [Duf77]. correction [Deu79a, Deu79b, Deu79c, Deu79d, Fro71, Han78b, Rem72]. corrections [DHN79, EJN74]. correlated [WD77]. correlation [MRCL76, RSL^{+}78, WDM77]. correlations [CSP^{+}79, JB77, JR75]. corresponding [Sch74]. cosine [TS75b]. Coulomb [AAJ76, AL79, BDG^{+}72, BFSG74, Bar76, DF76, FB75, Köl72b, LJLC72, PS76, RR76, RSA74, SS71, SF76, TTW79b]. Coulomb-nuclear [FB75]. Counting [Aur77]. coupled [Ahl79, BG72, Cha73, Che73, Chi73a, Cla78, Com78, CSW78, GPS75, Hir73, HH75, HLS75, Hub70a, NK78, PS76, RSA74]. coupling [BCE79, CN73, Cha74, Cla78, DW72, Gra72, Hub70a, Hub70b, KK77, Nus70, PGB78, Sea73, Tam70, Wil71]. CPC [Ano76g, Ano76h]. create [Rob79c]. criteria [DSK72, VD74]. critical [CSC74, Fro73]. cross [All72, BKE79, Bet75, BTS73, BTS74, Com78, CLS^{+}70, Hei72, Hen75, LJLC72, McC75, MMM74, MT79, PV74, PS77, Sar72, Sar78, SMD71, SR73, Yos78, vdB74]. cross-collision [McC75]. cross-section [Bet75, Yos78]. cross-sections [PS77, vdB74]. crossings [Vio72]. crystal [Ell75, Han78d, PS74, Pre79a, Ste71]. crystalline [Han78b, Han78c, Han78e, Ryc78]. crystallographic [CA70]. crystals [De 75a, Han78a, Han78f, LB76, SF76, WD77, WDM77, Wor72]. cubic [AAMB79, BDD78, Bok78, LB76, WD77, WDM77, vM76b]. curve [HE73, vM75b, vM76a, vM79a]. curve-fitting [vM76a, vM79a]. curves [AMB75, RS74, Tel73, dJ78]. CW [Smi78]. cylinder [Cop75b]. cylindrical [Hug71, Jad75, Pie73, ST72, TT78]. cylindrically [SF78].

D [Chr75a, CL79, May74, Bun76, CSC74, RCL75]. Dalton [Bee73]. damage [CE75, Tor73]. data [AAMB79, AZS^{+}79, CSP^{+}79, CM79, EJN74, GCCB73, Han78b, Han78c, Han78e, Hib74a, Hib79, IL74, KM74, KN73, Moo79a, Mor72, NM73, PB73, Rob79c, Rus73, Sal76, Sax78, Ugn78, Yat71, vM76b, vM78a, Smi72]. Database [Smi77]. DATSTOR [Rob79c]. David [Lai74]. Davydov [AFH71]. Day [Pat73]. DBLCON [War78]. DCSCH3 [FC71]. DCSCH4 [FC71]. decay [FC71, IL74]. deck [Fem71]. decomposition [KPP79, Zlo78]. decomposition-conjugate [KPP79]. defect [KM79b]. defects [DBB71, LB76, Nor73b]. deffered [Fro71]. deformation [KN79b, PI78]. deformation-dipole [KN79b]. deformations [Hir73]. Deformed [HH75]. degenerate [MW70, MW71a, MW71b, MW71c]. degradation [LNV78]. degrees [CIT73]. delay [CE75]. Delphi [Coh71]. Delsqphi [CH71a]. dense [FS73]. densities [Rob79b, Tal76]. density [Chr74a, Chr74b, Cra72]. dependent [Has76, LW75, Mag78, MPW73, Wei72]. deposition [MM74, MRW76, PSBG73]. Depth [MM74, MRW76]. Derivation [HE71]. derivatives [GMP76, JP78]. description [Coh71, WP73]. Design [Gli78, VD74, DSK72, Web73]. designed [DTE74]. Desy [ADF^{+}72]. detecting [CSP^{+}79, Sta72]. detectors [OVQT^{+}76, SK79, Wil72]. determinantal [Lab79a]. Determination [Cam75a, WDM77, CSC74, DG75]. Development [Str79, Zac72, Cas72, Sar78]. developments [BC79, GJ76]. device [BMS72]. diagonal [GMR75b]. Diagrammatic [Sil78a, Sil78b, Wil78, WS79]. Diagrams [Cal72, CW79a, GVV79, KK77, PW73, WS79, WP73]. diatomic [BB74, FT71, Gia76, Sal78, SLH78, Sin70, Tel73]. Diehl [Sut72]. dielectrics [LVPG78]. Difference [MPW73, Fro71, HH79, Mor76, PPR72]. Differential [LJLC72, BJ76, BTS73, BTS74, Cha73, CSW78, Dem72, EJ77, Jes79, May74, Par73, PPR72, PS77, SMD71, SR73, Ugn78]. differing [Fun79a, Fun79b]. diffraction [Gru79, Han78b, Han78c, Han78e, vM76b]. diffractive [Kas75]. Diffusion [Fos72, CRL78, De 77, WD77, WDM77]. diffusive [WBG79]. Digamma [Köl72a]. dilute [KSSA78, OS71, Tho79]. dimensional [ABG^{+}75, AK76, BG72, Blo72, Cho70, CH71a, CAR74, CRL78, CW79b, Dem78, DH75, EG74, Has76, LL72, LNS76, MPW73, RR74, RM76a, RM76b, RM76c, WBG79, Zlo79]. dimensionally [GVV79]. dimensions [Bos76b, CKM^{+}73]. dipolar [LVPG78]. Dipole [DF76, Hib75, KN79b]. DIRAC [DSK73, Des75, DSK72]. Direct [BG72, Klu72a, FS75a]. directional [MRCL76, RSL^{+}78]. Dirichlet [Cam75b]. discharge [DE72]. discharges [DEW78]. discontinuities [Moo79a]. discrete [AN79, CCR79]. discretization [Rei79, RD79]. discussion [FJK^{+}72]. dislocations [Pet76a]. disorder [MR77]. dispersion [Han78d]. displacement [JB77]. display [Kar78]. distored [DE72]. Distorted [TCR71, vM77]. distortion [HE73]. distribution [BB74, BRT78, Han78c, Han78e, Lan70, MM74, MRW76, Pie73, RSL^{+}78]. distributions [Rou72, Sch74, SDH74, TSD76]. DM [Sut72]. DM57 [Mor78b]. documentation [BPW72c]. Domus [Zlo79]. dose [Hub73]. dosimetry [Hub71]. Double [She76, BCE79, HEL71, HE71, SMD71]. double-differential [SMD71]. Doublet [VG74, KK75]. drawn [BS75a]. drift [De 77]. drift-diffusion [De 77]. drifting [FM79]. drop [PI78]. Durham [GCCB73]. during [OBLR73]. DWBA [But78, PV74, TL74, TUWA79]. DWPI [EM76, Fun79a]. dynamic [CHR75b, DSK72, DSK73, Rin79]. dynamical [CIT73, MS72]. dynamics [CKM^{+}73, KN79b, KN79a, MAK^{+}72, NW79, Tom73, WW74, ZS77, Cam72, Chr75a, Gos78]. dynamo [Moo79b].

Ecsimpact [LS79]. ed [AF72b, Cam72, War73]. edge [Pet76a]. edited [Bee73, Chr75a, Gos78, Nor73b]. edition [BPW72a, BPW72b]. editor [Day77]. Editorial [Ano71a, Ano72g, Bur70a, Bur70b, Bur72b, Bur74, Ano72h, Ano72i, Ano73f, Ano73g, Ano74h, Ano74i, Ano75g, Ano75h, Ano76i, Ano76j, Ano77b, Ano78i, Ano78j, Ano78k, Ano79j, Ano79k]. Editors [Ano74j, Ano76d, Ano77a, Ano78d, Ano78e]. eds [Bra72c, Bur72a, McN71, Mor78b, Sut72, Zac74]. effect [DG75, DT79, Kac77, TT78, Vio72, WS74]. effective [HIM73, LR72, MGTD73]. effective-range [LR72]. effects [Cop74, Cop75a, Cop75b, Ell75, GH78, Sei73]. efficient [CHMM73, KS75a]. eigenenergies [AL79]. eigenfunctions [AAJ76]. eigenproblem [Nas74]. Eigenstates [MPS75]. eigensystem [Gar74]. eigenvalue [CB77, FH74, Gru75, MAK^{+}72]. eigenvalues [AAJ76, Gue74, MPS75]. Eikonal [Joa73]. Einstein [Pay72]. EISPACK [Gar74]. elastic [All72, Com78, EM74, Fun79b, KF71, LB76, MS78, Pet76a, Smi70b, TTW79a, Yat71]. elastic-scattering [Com78]. electric [God78, Hib75, LJLC72, ÖW70, SS71]. electric-dipole [Hib75]. electrical [Isl73, RM76a, RM76b, RM76c]. electrodes [DEW78]. electrodynamics [Rem72]. Electromagnetic [BF71, Bun76]. Electron [MMM74, PSBG73, Sin70, AT75, Ahl79, BC79, BDD78, BBL79, Chi73a, CLS^{+}70, CSW78, FT71, Gia76, GH78, Haw73, Hib74b, Joa73, Lan70, Lew79, MS78, NM73, Nes79, Rei73, Sal78, ST73, Sta72, SV79, TYCT73, TSD76, Web73, WB74, Yat71]. electron- [Web73]. electron-atom [AT75, CSW78, Joa73, NM73, Nes79, Rei73, ST73, TYCT73, Yat71]. electron-molecule [BBL79]. electron-photon [SV79]. electronic [BB74, FS73, HRB72, JVS72, Rei79, WC78]. electronics [Kol78]. electrons [AM74, BJ76, BGMP79, BB77, Chi73b, Gra72, Moo71c, Per73, SS74, SS76]. electrostatic [Bun76]. electrotransport [OBLR73]. element [ATV76, BGT76, KS75b, KS73, KS75c, Kon78b, Mor76, PB72b, TT78]. elementary [GCCB73]. elements [Ano78p, ABG^{+}75, Bra78a, Cha74, Cla78, Gla78, Hod76, Jam70, Jon74, Klo75b, Lab79a, Lew79, McC75, Nil72, Par73, PGB78, Rob73, Sal78, SS71, Sax77a, Sax77b, Sax78, FS75b]. Ellips [Ber77]. ellipsometer [Ber77]. elliptic [EJ77, Jes79, Jon72]. Elsasser [Moo79b]. emission [Duf77]. emitters [Pie73]. emphasis [vdSVSH73]. energies [Lew70b, Mag75, PI78, Sil78b, Wil70, Wil78, WB74]. Energy [AFH71, AF72a, AF74, CJ71, KK77, BGMP79, Böc75, CMS77, Cug73, FJK^{+}72, HIM73, HRB72, JPB78, KVW70, Kol78, KK75, LB76, LLC76, MM74, MRW76, Nil72, PSBG73, RFSG72a, Ran72, RR74, Sko74, VD74, WC78, Zab73]. Energy-level [AF74]. Energy-loss [CJ71]. engineers [Del77]. England [Ano79f]. Entire [BDG^{+}72]. environment [For78]. EPR [Ste76]. EPS [Ano79f]. equal [Jab73]. equation [Bel72, CE75, Cha75, EG74, FS75a, FC74, HH79, HD78, Hug71, Kil79, LR71, Mar73, Mor73, RA78, Ste79, TSD76, Wei72]. equations [BG72, Bor76, Cha73, Che73, Chi73a, CSW78, De 77, Dem72, EJ77, GPS75, Has76, HLS75, Jes79, Jon72, Par73, PPR72, PS76, TA72, WBG79, May74]. EQUIL [RCL75]. equilibria [BGT76, Lac76]. equilibrium [Bet75, BM76, CRL78, Dem78, Joh78, Mik74, RD79, RCL75]. equivalent [AM74, BB77, Chi73b, Gra72]. Erratum [Ano70f, Ano70g, Ano70c, Ano70d, Ano70e, Ano71b, Ano71c, Ano72j, Ano72k, Ano73j, Ano73l, Ano73k, Ano73h, Ano73i, Ano74k, Ano74l, Ano74m, Ano74n, Ano75i, Ano75j, Ano75k, Ano75l, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano76k, Ano76l, Ano77c, Ano77d, Ano77e, Ano77f, Ano77h, Ano77g, Ano78l, Ano78m, Ano78n, Ano78o, Ano79l, Ano79o, Ano79m, Ano79n]. Error [RR76]. Errors [Gue74, JR75]. especially [Aue78]. establishing [FC71]. estimation [Joh78, Pac72]. Europe [Aig72]. European [Bis75]. Européen [Ano75c]. eV [vdM72]. eV-region [vdM72]. EVA2 [BPW72a]. evaluate [Cha74, GH76, Rob73]. Evaluation [BFKD77, SG76, FP73, GVV79, HKK73, KK71, Kon78a, Moo79b, Mue71, Sax77a, SC71, ST73, Str74, TS75b, Wil71]. EVAR [BPW72b]. event [LB78]. events [BD72, BCV^{+}72, BEK^{+}72]. evolution [BMS72]. ex [LS79]. ex-core [LS79]. Exact [DG75, FKU^{+}76, FP73, Gol78, KH78, TL74, TUWA79, Mar73]. examples [Bog77]. exchange [AT75, Par73, ST73, vdB74]. excitation [Kas75, KK75, LJLC72, MMM74, RR76, RSA74, SS71, TSD76]. excited [Lab79b, Mal70, Tal76, WS79]. exemplified [BEK^{+}72]. expansion [Del79a, Lew70b, Lew79, Mor73, Ort72, Sil78a, Sil78b, Wil78, WS79]. Expansions [She76, She74]. experiment [VD74]. experimental [ABMT78, BJ76, BGM79, IL74]. experiments [Ald72, Bou72, CIT73, Cop74, Cop75a, Cop75b, Haz72, Ugn78]. Explicit [Dem72, RB75]. explosions [CHR75b]. Exponential [RA78, Dem78]. Exponential-fitting [RA78]. expressions [EK72, Str74]. extended [AF74, DBB71, KE74]. extensive [Goo74, KS75a, Sta72]. External [Wor72, BS75a]. extraction [IL74, Mor72]. extrapolation [CC79]. extrapolations [CSP^{+}79, CC79].

F [Bee73, CL79]. facilities [HR74b]. factor [Han78b, Isl73, SS74, SS76, TUWA79, WDM77]. factors [BS77, DW72, Fun79a, Han78d]. Faddeev [Che73]. Faddeev-type [Che73]. Faradaic [BGM79]. fast [CHW^{+}78, Cha75, Tel73, CH71b, FSR78, Mar76]. fast-neutron [Cha75]. Fernbach [Bra72c, McN71]. ferromagnetism [Joh78]. Feshbach [Smi70a]. few [CIT73]. Feynman [Cal72, GVV79, PW73, WP73]. Field [Mat72, BM76, CCR76, Des70, Ell75, Klu72b, KK77, KS73, KS75c, LCW71, O’C76, ÖW70, Ste71]. fields [Ano74j, BS75a, DE72, Mik74, PW72, Pet76a, TSD76]. FIFPC [FSR78]. fifth [Gos78]. file [PB73]. file-handling [PB73]. films [Hub70c, Hub70d]. filtering [Pac72]. final [Bra72b, Bra72a]. finding [CSP^{+}79]. Fine [Sar72, Sar78]. Finite [ATV76, HH79, Mor76, TT78, Ano78p, ABG^{+}75, BGT76, FKU^{+}76, Hod76, KS75b, KH78, KS73, KS75c, Kon78b, NK78, PV74, PPR72, TL74, TUWA79]. Finite-difference [HH79]. finite-element [KS75b]. finite-range [FKU^{+}76, PV74]. First [Hub73, GMP76, JP78]. Fischer [Gra78]. fit [Han78f, VG74]. fits [KL72]. Fitting [Com78, CH75, Dem78, EK72, O’C74a, RA78, Ste71, Wil70, WS74, vM75b, vM76a, vM79a]. fixed [Bat74]. flexible [WS74, vM76a]. floods [Tom73]. flow [Blo72, FCW78]. flows [Dem78, RD79, TW79]. Fluck [Sut72]. Fluid [Cam72, Chr75a, Gos78, RC72]. flux [SF78]. Fock [Des70, Gra78, Des75, AAH72, ABH^{+}72, Bat74, CCR76, CCR79, Fro70, Fro72, Fro78, Lab79c]. foil [IL74]. Fokker [FSR78]. force [Han78a, Han78f]. forces [LR71, LR72]. FORDOC [BPW72c]. forecasting [Tom73]. Form [SS74, Ano72l, BM76, Fun79a, Isl73, Sal76, SS76, TUWA79]. form-factor [Isl73]. FORMAC [McC75, Pay72]. FORMAC73 [Pay76]. formal [HKK73]. formalism [CSC74, Hof71]. format [Sal76]. formation [Tsc76]. formula [BK74, Jab73, KB73]. formulae [Sei73]. Formulas [RB75]. formulation [CN73]. FORTRAN [BPW72c, Bar77, Ber77, CR74, PS78, Sal76, SR73, Zoh72, vM79b, AD73, MW71a, MW71b, PV74, RV78, RCL75, SMD71, SDH74, vM75c, vM76b, vM77, vM78b, vM78a, Pat73]. Fortran-IV [SDH74]. Fortschritte [Sut72]. four [Mue71]. four-body [Mue71]. Four67 [CH71b]. Fourgen [Mar76]. Fourier [CH71b, Mar76, Col77, CT79, PS78, TS75b]. Fourth [Chr75a, CS78, Joh73, WS79]. fourth-order [CS78, WS79]. Fox [AF72b]. Fractional [AM74, Gra72, Chi73b, Hub70b, MNJL71]. frames [SM78b]. Fredholm [Rei73]. freedom [CIT73]. Freint [Gru79]. frequencies [Han78f]. Fresnel [Gru79]. Freudenthal [KB73]. frictional [Has76]. Froese [Gra78]. Frozen [CCR79]. full [CHR75b]. Function [Köl72a, Bar77, Cha75, Del79b, Deu79b, Deu79c, Deu79d, HD78, JR75, KSSA78, LB76, TYCT73, dJ78]. Functions [AM77, BDG^{+}72, SM78b, She76, TTW79b, AT75, And70, AM78, AM79, BFSG74, Cam79, Cug73, DF76, FS75c, Han78c, Han78e, Hib75, Jam70, KM74, Lab79d, Ort72, PB72a, Rob70, RBHW76, Sal78, She74, ST73, SG72, Ste79, Yur78, vM79a]. Furi [Bar77]. fusion [CAR74].

G [Bur72a, Zac74]. Gamma [Hub71, Köl72a, AWG71, BCV^{+}72, FC71, Hub73, MRCL76, RSL^{+}78, SS74, SS76, Wil72]. gamma-gamma [MRCL76, RSL^{+}78]. Gamma-radiation [Hub71]. gamma-ray [AWG71, FC71, Hub73, Wil72]. Gammel [Bur72a]. gas [OS71, Tho79, TW79]. gaseous [DE72, McC75, PSBG73]. gases [BJ76, TSD76]. Gatchina [KKN^{+}72]. gauges [Chr74a, Chr74b]. Gaunt [Moo79b]. Gaussian [AAH72, Mat72, VG74, War78]. Gaussians [vM75b]. Gehlen [Nor73b]. general [And70, Aue78, BBC^{+}74, BBL^{+}78, Bur70c, Cha73, DS75, Fro78, Gra73a, Gra73b, Hib70, Hib71, Hib74b, Hib74a, Hib75, Klo75a, Pay76, VD74, WS74, vM75a, PW73, Sil78a]. general-purpose [vM75a]. Generalisation [Chi72]. Generalized [SG72, CB77, Dob79, FH74, MP74, Zoh72]. generate [CW79a, Rob70, Rob79a]. generated [BK72]. generating [BB70, Bos76a]. Generation [SM78c, BD72, Han78a, KVW70, PW73, Rap74, WP73]. generator [Jad75, KPP79, KCA76, Mar76, Lab79a, Lab79b]. generators [GR78]. Geomagnetic [Klu72b, KL72, Moo79b]. geometries [HKK73]. geometry [Cop75a, Cop75b, EG74, Hub71, ST72]. Geophysics [Pek72]. GeV [RR74]. given [Rob79a]. Glowcode [LNS76]. Glowinski [Mor78b]. Gordan [RB75, SM78c]. Goudsmit [Vio72]. gradient [Klu72b, KPP79, Mat72]. Gram [O’C74a]. graphic [Kar78]. Graphical [Cra72, MMK73, Rou72]. graphs [Rap74]. Graves [War73]. Graves-Morris [War73]. gravitional [CH74]. Green [HD78, LB76, RBHW76, TYCT73]. ground [And70, KH78]. ground-state [And70]. Group [WW72, CB77, HM71b, NW79, SM78b, WW74]. group-coordinate [CB77]. Group-theoretical [WW72, WW74]. groups [ADO78, BK72, KSSA78, KK77, Net75, TS78]. growth [DE72, DEW78, KM79b]. Grundlagen [Sut72]. guide [AD73, CSP^{+}79, CC79, Hoa70]. gun [Web73].

H [CL79, Sut72, BS77, vdM72]. hadron [RR74]. hadronic [Ran72]. hadrons [KVW70]. half [Nil72]. half-lives [Nil72]. Hamiltonian [Hib74b, HM71c, MGTD73, Ste76]. handle [Fun79b]. handler [Taf75b]. handling [PB73]. hardware [EJ77, Kol78]. harmonic [Dob79, FT75, SG72]. Harmonics [Bra73]. Hartree [Gra78, AAH72, ABH^{+}72, Bat74, CCR76, CCR79, Des70, Fro70, Fro72, Fro78, Lab79c]. Hauser [Smi70a]. HBr [vdM72]. heat [Fos72]. heavy [Aue78, But78, CJ71, Dic79, FKU^{+}76, Nil72, Par73, TL74, TUWA79]. heavy-ion [Aue78, FKU^{+}76, TL74, TUWA79]. Heidelberg [BEK^{+}72]. height [AK76, Deu79a, Deu79b, Deu79c, Deu79d, vM79b]. Heisenberg [KSSA78]. helium [HH79, Kno70]. Herman [Bee73]. Hermitian [Nas74]. heuristic [KM79a, Pas72]. Hierarchical [Haz72]. High [MMK73, SC71, Böc75, Cra72, FJK^{+}72, HIM73, KVW70, Kol78, LL72, Ran72, RR74, SM78a, SDH74, Sut72, VD74]. high-density [Cra72]. high-energy [Böc75, FJK^{+}72, RR74, VD74]. high-lying [SDH74]. high-resolution [Sut72]. Higher [Moi73, OC78, DHN79, TS75a]. higher-order [DHN79]. histogram [BEM79]. histograms [vM78b]. HLISP [KSSA78]. Holland [Hib79]. Holt [Cam72]. homogeneous [KS79, TT78]. homonuclear [SLH78]. horizontal [Cop75b]. hot [FM79]. hydrodynamical [FM79]. hydrogen [BBH^{+}77, Mal70, Tal76]. hydrogen-like [Tal76]. Hydrogenic [Lew70a, Jam70, Moo71c]. hydromagnetic [Wes76]. Hymnia [Gru75]. hyperbolic [Dem72]. hyperfine [BCE79, GH76].

Ian [Smi77]. IBM [BTS74, HRR75, NA72]. ideal [HW77, Lac76]. identical [BB70]. Identification [OVQT^{+}76]. II [BP73, Bra78a, CSP^{+}79, DSK73, Deu79b, Han78c, Han78d, HM71b, KN79a, KPPR76, Lab79b, MW71b, MW71c, PS77, Rob79b, RM76b, Sil78b, WP73]. III [Deu79c, Han78e, Han78f, HM71c, Lab79c, PS76, Rob79c, RM76c, SMD71, Wil78]. ill [TA72]. ill-conditioned [TA72]. ILTHII [Sal75]. ILUCG [KPP79]. immersed [FM79]. Impact [Cas72, Cha72, Bur72c, GPS75, LS79, MMM74, PS76, PS77, CSW78]. impedance [BGM79]. implementation [Pel78, DSK73]. importance [KVW70]. Improper [Ned72]. improve [Taf75a]. Improved [WW74, CH75, Fro72]. including [DHN79, EJN74, Ell75]. Inclusion [Hib74b, Hib74a, Sax77a, Sax78]. inclusive [MT76, MT79]. incoherent [SR78]. incomplete [KPP79]. incorporate [Fun79a]. increase [Hub70d]. index [Ano70h, Ano70i, Ano71d, Ano71f, Ano72b, Ano72c, Ano72d, Ano72o, Ano72p, Ano73b, Ano73c, Ano73n, Ano73o, Ano74b, Ano74c, Ano74p, Ano74q, Ano75b, Ano75a, Ano75s, Ano75r, Ano76b, Ano76c, Ano76p, Ano76q, Ano78a, Ano78b, Ano78c, Ano78r, Ano78s, Ano78t, Ano79c, Ano79d, Ano79e, Ano79p, Ano79q, Ano79r]. Individual [Ano72l]. induced [FKU^{+}76, TL74, TUWA79]. induction [Klu72b]. industrial [Cas72]. inelastic [BKE79, Bog77, DT79, EM76, Fun79a, Han78d, Hof74, LWSS79, SR78]. inequalities [KS79]. infinite [MAK^{+}72]. information [DSK72, DSK73, HM71b, Rob79c]. infrared [MW70, MW71a, MW71b, MW71c]. inhomogeneous [Cha73]. initial [CR74]. initial-value [CR74]. Inner [TS78, KB73]. input [Sax78]. instabilities [Wes76]. instruction [KS78]. Instructions [Ano74o, Ano76m, Ano77i]. instrument [Han78b]. instrumentation [HR74a]. integer [AM78, AM79, KCA76]. integral [BTS73, BTS74, CE75, De 77, FS75a, FC74, HD78, Jab73, Mor78a, MR73, Ste79, Yur78]. integrals [AT75, BDD78, BFKD77, DF76, FP73, Gio79, Gol78, Hib70, Hib71, Hib74b, Hib74a, Hol79, Lew70a, Mat72, Moo79b, NA72, OS71, ST73, SK79, WB74, Yur78, ZDN78]. integration [BG72, Gru79, Kil79, Mor78a, PS76, PB72b, GPS75]. integro [CSW78]. integro-differential [CSW78]. intensities [FK73, Fer75, Hof71, Hof74]. Intensity [Deu79a, Deu79c, BB74, BRT78, HE71, IL74, Lan70, LL72, ÖW70, RS74, RS75]. Inter [BDD78]. Inter-electron [BDD78]. interaction [AN79, ATV76, BGMP79, Bog77, BCE79, DT79, Ell75, FT71, Gia76, Hib75, Jam70, Joh73, Klo75a, Sax77b, vdSVSH73]. interactions [BKK76, Ste79, vM78b]. Interactive [Rou72, VDM78, AZS^{+}79, AK76]. Interatomic [Nor73b]. interelectronic [MR73]. interfacing [Zac74]. interference [FB75]. interferometer [HE71]. interior [CC79]. Intermediate [Cla78, Rus73]. Intermediate-level [Rus73]. internal [DHN79, DH71, PR75]. International [Ano78p, Cam72, Chr75a, Gos78]. interpolation [AAMB79, CM79, Moo79a]. interpret [vM76b]. Interscience [Moo72]. interstellar [BS77]. intrinsic [ABH^{+}72]. introduction [Rob74]. invariant [Jab73, Mor78a]. Inversion [FS75a, FC74, Hof74, LL79]. investigation [MS72]. investigations [SSZ73]. involved [Hei72]. involving [BFKD77, WS79]. Ion [FSR78, Aue78, But78, FKU^{+}76, KN79b, MM74, MRW76, TL74, TUWA79, Web73]. ion-gun [Web73]. ionic [SF76]. ionisation [Mag78]. ionographic [OVQT^{+}76]. ions [LCW71, Sko74, Ste76, Tal76]. Irreducible [Wor73, Bos76b, KSSA78, Net75]. Ising [CSC74, Joh78, KH78, vdSVSH73]. isoelectronic [HH79, Kno70]. isolated [Bou72]. Isometric [Cho70]. isospin [Hub70b]. iteration [Cam75a]. IV [Deu79d, Lab79d, NW79, SR73, SDH74, WS79]. IX [McN71].

J [AF72b, Bur72a, Gos78, May74, Mor78b, Nor73b]. Jaffee [Nor73b]. jajom [Sar78]. January [Ano74d, Ano74e, Ano76h]. Jennings [Del77]. Job [BP73].

John [Del77, Gra78, Lai74, May74]. Jr [Bur72a]. jump [Moo79a]. junction [Blo72].

Kac [Cha76]. Kahane [Chi72]. KBr [vdM72]. Kellerhals [Sut72]. Kenneth [Moo72]. Kinematic [CH75]. Kinematical [WMC71]. Kinematics [Ass75, Bra72b, Bra72a]. kinetic [Joh78]. kinetics [DST75]. Koehler [Bee73]. Kosfeld [Sut72]. Kramers [Col77, CT79, KN73]. Kronig [Col77, CT79, KN73].

L [Bur72a, Mor78b]. labelling [MPS75]. laboratory [Haz72, KS78, Pen78]. LACC [Kon78a]. ladder [Sil78b]. Lambert [May74]. Landau [Sch74]. language [FP73, Pel78]. languages [Hoa70, Rus73, Hoa70]. Lanthanide [Ste71]. Laplace [HD78]. Laplacian [O’C76]. large [AZS^{+}79, AAB75, BPZ73, BP73, CB77, PB73]. laser [CAR74, CW79b, DST75, DH75, RS76, Smi78]. lateral [MRW76]. Lattice [KN79b, KN79a, NW79, Nor73b, HM71a, MR77, ÖW70, WW74, WW72]. lattices [CMS77, Zab73]. Laue [CA70, Pre79a, Pre79b]. layer [RD79, RS74]. layer-by-layer [RS74]. layered [Chr74b]. layers [Sko74]. LCGO [WC78]. learning [KM79a]. least [Bok78, Han78f, LM71, O’C74a, Pac72, RFSG72b, WS74, vM75b]. least-squares [RFSG72b, vM75b]. Lecture [Cam72, Mor78b]. LEED [Hof71, Hof74, RS74, RS75]. Legendre [Bra73, Del79a]. lengths [JP78]. level [AFH71, AF72a, AF74, Cro72, RFSG72a, RFSG72b, Rus73, Taf75b, Vio72, Wil70]. levels [SDH74]. LIANA [Smi70a]. library [Ano70b, Ano72a, Ano74d, Ano74e, Ano75d, Ano76g, Ano76h]. Lie [Bos76a, Bos76b]. lifetime [KE72, KE74, VDM78, War78]. lifetimes [IL74]. ligand [KK77]. light [HE73, RS76]. like [Moo71c, Tal76]. limit [BGMP79]. limited [EJ77, Han78c, Han78e]. Limits [JPB78]. line [ADF^{+}72, BS75b, Cra72, Duf77, GPS75, Jon74, KKN^{+}72, KK75, PS77, Rus73, Sch78, Ugn78]. Linear [HW77, Ano78p, Jes79, KSSA78, KS79, Kon78b, TA72, Wes76]. linearized [McC75]. lines [BS77, Jon72]. linked [WS79]. Lions [Mor78b]. Liquid [PI78, Isl73, Sch73]. LISP [Cam70]. List [Ano72m, Ano76d, Ano76n, Ano78d, Ano78e, Ano77a]. lives [Nil72]. load [KM79a]. local [Ste79]. Logarithm [Köl72a]. London [Smi77, Zac74]. Lorentz [Jab73]. Lorentz-invariant [Jab73]. Lorentzians [vM75b]. loss [CJ71, Sko74]. low [BGMP79]. LS [Cla78]. Lustig [Sut72]. lying [SDH74].

M [Bra72c, Cam72, Hoa70, McN71, KKN^{+}72]. M1 [BF71]. Machine [Zan72]. machines [LS79]. macromolecules [LNV78]. MACSYMA [BKK76]. made [HE71, RD79]. Madelung [JP78, ÖW70]. magnetic [BM76, Gol70, Hub70c, Hub70d, Klu72a, Mik74, PW72, vM78a]. magnetohydrodynamic [ABG^{+}75]. magnetoplasma [ST72]. magnetospheres [Pet76b]. magnetotelluric [KS75b]. making [HR74a, SDH74]. Mandy [SMD71]. manipulation [BEM79, FP73]. Manual [Str74, EK73]. Many [Ahl79, Kel79, vdSVSH73, GH78, Lew79, Sil78a, Sil78b, TYCT73, Wil78, WS79]. many-body [Sil78a, Sil78b, TYCT73, Wil78, WS79]. many-electron [Lew79]. Many-particle [vdSVSH73]. mappings [BS74]. masses [Jab73]. matching [BSP71, Hof71]. materials [Han78b, Han78c, Han78e]. mathematical [Sou73, Zhi72]. matrices [BTS73, CB77, Chi72, Cho70, Han78a, Hen75, MAK^{+}72, Sar70, Sar72, Sar78]. Matrix [FS75b, Hod76, Nes73, Nes79, BBC^{+}74, BBL^{+}78, BC72, Bra78a, BBL79, Bur73, Cha74, Cla78, Gar74, Gla78, Gru75, HM71c, Jam70, Jon74, KSSA78, Klo75a, Klo75b, Lab79a, Lew79, LWSS79, McC75, Nas74, Par73, PB72b, PGB78, Rei79, Rob72, Rob73, Sal78, SS71, Sax77a, Sax77b, Sax78, Sta76, Hof74, Del77]. matter [RR74, Sko74]. may [dJ78]. MCBP [BGMP79]. MCHF [FS75c]. measured [Rou72]. measurement [EGMS72]. measurements [BGM79, HR74a]. mechanical [BTS73, RSA74]. mechanics [Ald72, Ano78p, LL79, RC72]. media [Chr74a, Chr74b, FS73, KK75]. medium [vM75c]. medium-scale [vM75c]. MEDUSA [CAR74]. Melting [CKM^{+}73]. memory [LS79]. metal [FS73]. metal-non-metal [FS73]. metals [DBB71, Isl73]. meteorology [Cha72]. method [ABG^{+}75, BBC^{+}74, BBL^{+}78, BCV^{+}72, Bur73, BCE79, CB77, Col77, CT79, CH75, DHD74, Dem72, Dem78, EG74, FS75a, Fro71, Gra78, Gue74, Hof74, HD78, Jon72, Jon74, KPH78, KVW70, Kla71, Kon78b, KPP79, LB78, Moo79a, Nes73, Ort72, Rob72, RS74, RM76a, RM76b, RM76c, Tel73, TW79, WC78, Wil72]. methodological [Sta72]. Methods [Cam72, Chr75a, Fle72, Mor78b, BH72, Cal73, Fos72, Gos78, Hib79, Kil79, LWSS79, May74, Moi73, MPW73, Mor76, Nor73a, Per73, PPR72, RA78, TTW79a, TYCT73, Zhi72, Bra72c, McN71]. MeV [RR74]. MHD [BGT76, GJ76, HW77, Lac76, RCL75, WBG79]. microcrystals [CKM^{+}73]. microscopic [FKU^{+}76, Fun79a, vM78b]. migrating [WD77]. millimeter [Smi78]. Minicomputer [Zac74]. minicomputers [VG74, vMG73]. Miniconsult [Zac74]. minimization [JR75, Tro73]. Minuit [JR75]. mixed [BF71, RSL^{+}78]. mixing [KF71, MRCL76]. mixture [PSBG73, Zlo78]. mixtures [McC75, Tho79]. mode [Han78a]. model [AFH71, AF72a, AF74, Aue78, Bet75, CSC74, CHMM73, DTE74, ETH73, GCCB73, GMP76, GA74, Hof74, Joh78, Kas75, KN79b, KN79a, MR77, NW79, PI78, vdSVSH73, Smi78]. Modelling [JB77, KS75b, Pet76b, Rob79a, Rob79b, Rob79c]. models [Ahl79, Klu72b, LW75, Sou73]. modes [Han78d, Wor72]. Modification [Bar76, Cop75a, Cop75b]. Modified [Bra73, O’C74a]. modulation [Ber77]. module [RCL75]. Moléculaire [Ano75c]. Molecular [Ano79f, BKE79, BFKD77, Bra72c, Con79, CKM^{+}73, Cro72, GPS75, MAK^{+}72, PS76, PS77, Rei79, Wor72, ZS77]. molecule [BBL79, DT79, FT71, Gia76]. molecules [BB74, Han78a, Han78f, Han78d, Sal78, SLH78, Sil78a, Sil78b, Sin70, Wil78, WS79]. Moment [Rei79]. Momenta [TTW79b]. momentum [ET76, Gra73a, Gra76b, Hib70, Hib71, Hib74b, Hib74a, Hub70a, Jab73, LW75, RV78, TTW79a, Tam70, Wil71]. momentum-space [Jab73]. monoenergetic [Lan70]. Monte [Chr74b, Cop74, Cop75a, Cop75b, DS75, Jad75, Kas75, KVW70, Mor78a, MR77, Per73, Ran72, RR74, Sta72, SK79, Wil72]. Morris [War73]. Morse [Zac74, PB72a]. Moshinsky [LJ71]. Mössbauer [CT71, Gro72, MGTD73, WS74]. motion [BPW71, BPW72a, BPW72b, Gio79, WDM77]. Mu [CL79]. multi [Fro70, Fro72, Fro78, KKN^{+}72, PB72b]. multi-configuration [Fro70, Fro72, Fro78]. multi-particle [PB72b]. multi-users [KKN^{+}72]. multichannel [vM79b]. multiconfiguration [Des75]. multiconfigurational [Lab79c]. multidimensional [Gen72, ZDN78]. multiple [Cop74, Cop75a, Cop75b, CH74]. multiplicity [ADO78, BK74, KB73, TS78]. multiplier [Wor73]. multipolarities [RSL^{+}78]. multipoles [OC78]. Multistate [GPS75, PS76, PS77]. Multivariate [O’C74a]. Muon [Rin79]. Muonic [AV78, Rin79]. muonic-atom [Rin79]. mutual [BS75a].

N [Bee73]. N.F.E. [Isl73]. nanosecond [RS76]. narrow [EJ77]. Nearly [Mar73]. necessary [Han78c]. Néel [Hub70c, Hub70d]. neuronal [Sou73]. neutron [Cha75, Cop74, Cop75a, Cop75b, Fun79a, Fun79b, Han78b, Han78d, SR78]. Newnes [Hoa70]. Newnes-Butterworths [Hoa70]. Nilsson [BF71, Hir73]. NMR [Sut72, WDM77, vM77, Sut72]. no [PV74]. no-recoil [PV74]. Non [Dic79, SS76, Wes76, Ano78p, Bou72, Dem78, FS73, Han78b, Han78c, Han78e, Lab79a, RD79, Ryc78, ST73, War78, WB74]. non-crystalline [Han78b, Han78c, Han78e, Ryc78]. non-equilibrium [Dem78, RD79]. non-exchange [ST73]. non-Gaussian [War78]. non-isolated [Bou72]. Non-linear [Wes76, Ano78p]. Non-reactive [Dic79]. Non-relativistic [SS76, Lab79a, WB74]. noncircular [BGT76]. Nonlinear [Pet76a, BKK76, HW77, LM71, MS72, SSZ73, Zab73, Zhi72]. nonrelativistic [BRT78]. Normal [De 75a, Han78a, Han78d]. normalised [PB72a]. normalization [Han78b]. North [Hib79]. North-Holland [Hib79]. notation [FS75b, Lab79a, Sax77b]. Note [BB79]. Notes [Cam72, Mor78b]. notice [Ano70f, Ano73j, Ano73k, Ano74l, Ano74m, Ano74n, Ano75j, Ano75k, Ano75l, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano76k, Ano76l, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano78l, Ano78m, Ano78n, Ano78o, Ano79l, Ano79m, Ano79n]. notices [Ano70g, Ano71b, Ano71c, Ano72k, Ano73l, Ano77h, Ano79o]. Nottingham [Ano79f]. nozzle [FCW78]. Nuclear [Nil72, Sko74, Smi70b, AAH72, ABH^{+}72, AAB75, Bet75, DG75, DTE74, ETH73, FB75, Hei72, KKN^{+}72, KM79a, LW75, OVQT^{+}76, Sei73, SMD71, SR73, SDH74, Smi70a, Sou73, TCR71, TL74, TUWA79, WMC71]. nucleation [KM79b]. nuclei [BF71, EM74, EM76, Hen75, Hod76, Kac77, PI78, RSL^{+}78]. nucleon [DHD74, FKU^{+}76, LR71, LR72, SDH74]. nucleons [Hub70a]. nucleus [Fun79a, Fun79b, SM78a]. null [LB78]. null-event [LB78]. Nullijn [dJ78]. number [Aur77, EJ77, Fem71, GR78, RFSG72b, Rob79b]. numbers [Sch74]. Numerical [Ald72, Bel72, Bor76, BMS72, Bou72, Cam72, CIT73, CSC74, Chr75a, Col77, CT79, De 77, FS73, HKK73, KL72, MAK^{+}72, Moo79b, Net75, Nor73a, Pet76b, Rob79a, Rob79b, Rob79c, Wei72, BG72, CMS77, CH74, EG74, For78, Fos72, FS75c, Gos78, Gra78, HD78, Jon72, Kil79, Lab79d, PB72b, RA78, Rob70, SSZ73, SZ79, TS75b, TYCT73, ZDN78]. Numerov [Fro71, Gue74].

objects [BB70]. observables [MT76]. observations [HE71, O’C76]. observed [Han78f]. obtain [Han78b, Han78f]. odd [BF71, Kac77, MR73]. odd- [Kac77]. off [GMR75b]. off-diagonal [GMR75b]. OLYMPUS [CR74, HRL75, Rob74, HR74b, HRR75]. On-line [Sch78, ADF^{+}72, KKN^{+}72, Rus73]. on-resonance [SDH74]. One [Has76, ABG^{+}75, AK76, BG72, CMS77, Cha74, CAR74, CRL78, DHD74, Dem78, Hib74b, Klo75b, Lew79, LNS76, MS78, PGB78, Rob73, Sal78, WB74, Yur78, dJ78, RM76a]. one-center [Yur78]. One-dimensional [Has76, ABG^{+}75, AK76, BG72, CAR74, CRL78, Dem78, LNS76, RM76a]. one-electron [Hib74b, Lew79, MS78, Sal78, WB74]. one-particle [Cha74, Klo75b, PGB78, Rob73]. only [Hei72, O’C76]. open [BDD78]. operators [Cha74, CW79a, Gla78, Lew79, McC75, MPS75, PGB78, Rob73]. OPIT [BP73, PB73]. optical [All72, Aue78, DTE74, ETH73, ET76, GMP76, Has76, TTW79a]. optically [KK75]. optics [Haw73, SSZ73]. optimal [CC79, Ste76]. optime [EK72, EK73]. optimization [Fle72]. orbit [Ell75, HE71, Klo75a, Sax77b, Tro73]. orbital [Cal73, HE73, Rob70]. orbitals [AT75, Bat74, BFKD77, Lab79c, Lab79d, Mat72, ST73]. orbits [Hir73]. order [Ano72l, AM78, AM79, BP77, BCE79, Cam79, Cha73, CS78, Dem72, DHN79, Lew70b, MR77, OC78, Sil78b, Wil78, WS79]. order-disorder [MR77]. Ordinary [Bra73, May74]. organization [Sil78a]. Organizing [Ano72n, Ano76o]. orientations [Pre79a]. oriented [Zlo78]. oscillations [Mik74]. Oscillator [FS75c, Dob79, FT75, God78, Hib75, SG72]. other [TS75a]. outer [BK74]. output [Cra72]. Overall [Coh71]. overlap [DF76]. overlapping [vM75b].

P [Gos78, Nor73b, Sut72, War73, Zac74]. package [AMB75, ABMT78, Cam75b, CH71b, CR74, Gar74, Gru75, HRL75, HRR75, KS75b, KS73, KS75c, Mag78, WC78, Yur78, vdB74]. packages [BPW72c]. packet [Has76]. Padé [Bur72a, Moi73, Sta76, War73]. pages [Sut72]. pair [Ahl79, Han78c, Han78e]. pairwise [Joh73]. pairwise-additive [Joh73]. Paker [Zac74]. Palmer [Smi77]. Panel [FJK^{+}72]. parallel [DEW78]. paramagnetic [Ell75]. parameter [DG75, GPS75, GMP76, JR75, Klu72a, KL72, PS76, PS77, Smi70b]. Parameterisation [For78]. parameters [Cam75a, Han78c, KF71, LR72, Pac72, PR75, Ste76, WDM77, WMC71]. parametric [Kla71]. parentage [AM74, Chi73b, Gra72, Hub70b, MNJL71]. parity [SDH74]. Part [Fer75, Hib74b, CSP^{+}79, CC79]. partial [CHMM73, EJ77, Jes79, PPR72]. partial-wave [CHMM73]. partially [RSL^{+}78]. participants [Ano72m, Ano76n]. particle [ATV76, Bun76, Cha74, Dic79, GCCB73, Hir73, JB77, Klo75b, Lew70b, MT76, Mor78a, Par73, PR75, PB72b, PGB78, Rob73, vdSVSH73, Zan72]. particle-solid [JB77]. particles [BPW71, BPW72a, BPW72b, BS75a, CN73, Cha73, CJ71, Hei72, Jab73, KK71, KF71]. partition [KSSA78]. parts [Jam70]. PATIWEN [FB75]. Pattern [ZS77]. patterns [Pre79a, Pre79b, RS75, Zan72, vM77]. PDP [DTE74]. PDP-15 [DTE74]. peak [vMG73, vM75a]. PEARL [Pel78]. penetrating [Sko74]. perform [vM79b]. performance [Smi78, Taf75a]. peripheral [PB72b]. permutation [KSSA78]. permutations [BB70]. Perturbation [CB78, Ahl79, BCE79, Kel79, Kil79, MS78, PW73, RS74, Sil78a, Sil78b, Wil78, WS79, WP73]. pervaded [Mik74]. Peter [Hoa70]. Phase [NM73, BB79, Jad75, KK71, KF71, MS78, Mor78a, PB72b, Ste79]. phenomena [Sch72]. phonon [Han78d, SR78]. photoelectric [KM74]. photographs [BSP71, CA70]. photoionization [CLS^{+}70, Jon74, Kel79]. photon [SV79]. photons [Chr74a, Chr74b]. Physical [Bis75, Cam70, Nad76, TA72, DSK72]. Physics [Ano74d, Ano74e, Ano75d, Ano76d, Ano77a, Ano78d, Ano78e, Ano79f, Cam72, Mor78b, Böc75, BS75a, Bur72c, Bur73, FJK^{+}72, KKN^{+}72, Kol78, McN71, Nad75, Ned72, Pas72, VD74, Zac72, Zan72, Zhi72, Bee73, Bra72c, Lai74, McN71, Bur72a]. pic [DH75]. pictures [EGMS72]. pion [Fun79a, Fun79b]. pion-nucleus [Fun79a, Fun79b]. pions [EM74, EM76, ET76]. Pipit [ET76]. PIRK [Fun79b, EM74]. pitfalls [Haz72]. place [Rob79a]. Planck [FSR78]. Plane [BDG^{+}72, DEW78]. plant [KM79a]. PLASKEM [Rob79b]. Plasma [McN71, BGT76, BMS72, CMS77, Daw72, FM79, FC74, LNS76, MS72, OC78, Rob79a, Rob79b, Rob79c, TT78]. plasmas [ATV76]. Plenum [Bee73, Nor73b]. Plot [Pre79b, CA70]. plotter [Moo71a, Moo71b]. plotting [AMB75, ABMT78, Cra72, MMK73]. point [CSC74, FM79, KK77, LB76]. points [AMB75, AAMB79, CC79]. Poisson [CH71a, Hug71]. Poisson-solver [CH71a]. POL [AZS^{+}79]. polarisation [Lan70]. polarization [Ber77, CS78, Sei73]. polarization-modulation [Ber77]. polarizations [Hei72]. Polarized [Cal73]. poles [CSP^{+}79]. poloidal [BM76]. polycrystal [SR78]. polylogarithms [MP74]. polymers [And70]. Polymol [And70]. polynomial [Del79a]. Polynomials [Bra73, Ort72]. Polytechnic [AF72b]. polytropic [BM76, Mik74]. population [SM78a, Tal76]. portability [Böc75]. portable [Day77]. position [KE74, War78]. positions [FFH76]. positive [Bos76a]. positron [KE72, VDM78]. positronfit [VDM78, KE72, KE74, War78]. Potential [Haz72, AAJ76, All72, Cug73, DT79, ET76, FT71, FM79, Gia76, Has76, Hir73, HH75, Kla71, Klu72b, Nil72, ÖW70, TTW79a, Tel73]. potential-energy [Nil72]. Potentials [Nor73b, AAH72, AL79, Joh73, LW75, PS74]. Potter [Lai74]. powder [vM76b, vM77]. powders [DG75]. power [Jam70, KM79a]. powers [MR73]. pp [Gra78]. Practical [Rob71, Hoa70, LM71, Smi77]. Practice [Smi72]. pre [Bet75]. pre-equilibrium [Bet75]. predict [RS76, Rob79b, Smi78]. Preface [Ano73m, Ano78q, Bis76, Bur79, Mac72]. preprocessor [HRR75]. presence [PW72]. preserving [BS74]. Press [AF72b, Bee73, Bra72c, Bur72a, McN71, Nor73b, Pat73, Smi72, War73]. pressure [RCL75]. Preuss [Yur78]. Price [Gra78]. prime [DW72]. principle [BB79, Wil70]. principles [Gli78, Sut72]. Printer [Moo71b, Cra72, Moo71a]. Printer-plotter [Moo71b, Moo71a]. probabilities [BF71, vdB74]. probability [Mal70]. probes [Ugn78]. problem [BH72, Cam75b, CB77, CB78, De 77, FH74, MAK^{+}72, ZKU79]. problems [BTS73, Cam70, Fle72, Gru75, HKK73, KS73, KS75c, LW75, MS72, Ned72, vdSVSH73, Sch72, Sta72, TYCT73, Zhi72]. procedure [DESL79, Gen72, GMR75b, LM71]. procedures [TYCT73]. Proceedings [Ano79f, Cam72, Zac74, AF72b, Chr75a, Gos78]. processes [BBC^{+}74, BBL^{+}78, LVPG78, RM76a, RM76b, RM76c, vdB74, Moo72]. Processing [AF72b, CMW74, Sch78]. processors [vdSVSH73]. product [CW79a]. production [MT76, PPR72]. products [Chi72, DW72, Klo75b, Sax77a]. profiles [DG75, OBLR73, SLH78]. Program [Ano72o, Ano72p, Ano73n, Ano73o, Ano74p, Ano74q, Ano75s, Ano75r, Ano76p, Ano76q, Ano78r, Ano78s, Ano78t, Ano79p, Ano79q, Ano79r, Bar76, Bet75, Böc75, BTS73, BTS74, Bur70b, Hof71, Hof74, KM74, MT79, NA72, ST73, Ste76, Wil70, AV78, ABH^{+}72, And70, Ano70b, Ano70i, Ano71f, Ano72a, Ano74d, Ano74e, Ano75d, Ano76g, Ano76h, ABG^{+}75, AK76, BRT78, BSP71, Bat74, BGMP79, BBC^{+}74, BBL^{+}78, BD72, BP77, BEM79, Bur70c, But78, Cam75b, CSP^{+}79, Cha73, Cha74, CW79a, CCR76, CCR79, Chi73a, Chi73b, CH71a, CC79, CLS^{+}70, CSW78, DHN79, Des75, Deu79a, Deu79c, DH71, DS75, Duf77, EM74, EM76, FT71, Fem71, FK73, Fer75, FFH76, Fro70, Fro72, Fro78, Gia76, Gio79, Goo74, Gra73a, Gra73b, Gra76b, Han78a]. program [Han78b, Han78c, Han78e, Han78f, Han78d, Hib70, Hib71, Hib74b, Hib74a, Hib75, HRB72, IL74, Jab73, Jam70, Kac77, KF71, KPH78, KSSA78, KE72, KE74, KS75b, Kla71, Klo75a, KS79, KS73, KS75c, Kon78a, KPPR76, KPP79, LCW71, LS79, MRCL76, MP74, MRW76, Mar76, MW71a, MW71b, MS78, Mor72, MT76, Mor78a, O’C74a, O’C74b, PB72a, PR75, PV74, PS78, Pre79b, PGB78, Rob70, Rob73, RCL75, RS76, Rob79a, Rob79b, Rob79c, RBHW76, RSL^{+}78, Sar78, Sax77a, Sax78, Sei73, SR73, SDH74, SS74, SS76, Smi70b, Ste79, Str74, Str79, Tro73, VD74, Vio72, VDM78, WB74, Yat71, Zlo79, Zoh72, dJ78, vM76a, vM76b, vM77, vM78b, vM78a, vM79a, vM79b, MW70, MW71c, WP73]. programme [Ano76o]. programmes [PPR72]. programming [AAB75, Hoa70, Nad76, Pel78, Rob71, Rus73, vdB74]. Programs [FC71, Köl72a, Sch74, AD73, BPZ73, BP73, CR74, Fro73, Gra71, Haw73, PB73, RR74, RV78, SMD71, Tho72, WC78]. Progress [BH72, Sut72]. project [GCCB73]. projection [ABH^{+}72]. projections [CA70, Pre79b]. prompt [War78]. propagation [DH75, LWSS79]. properties [FS73, Joh78, JVS72, Tho79]. property [CMS77]. proton [Fun79a, Fun79b, LL72, MT79]. protons [BBH^{+}77, Mal70]. protostars [Tsc76]. pseudorandom [GR78]. PULSAMP [RS76]. pulsar [Pet76b]. pulse [AK76, DH75, Sou73, vM79b]. pulse-height [AK76, vM79b]. pulses [RS76]. punching [Fem71]. pure [BF71]. purpose [DS75, vdSVSH73, VD74, vM75a, vM76a, vM79a].

quadratically [DESL79]. quadrature [Gen72, Rei79, Tel73]. quadruply [WS79]. quadruply-excited [WS79]. quadrupole [DG75, God78, LJLC72, SS71, vM77]. quadrupole-distorted [vM77]. quantal [RR76]. Quantum [RSA74, BTS73, FH74, LL79, Ner78, Rem72]. Quasi [DeV76, Tal76]. Quasi-bound [DeV76]. quasi-steady [Tal76]. Quasiclassical [CL79]. quasilinear [Dem72]. quasiparticle [HH75].

R [Bee73, Chr75a, Mor78b, Nor73b, Smi77, Sut72, War73]. Racah [AD73, BK74]. Radial [SF78, FS75c, Jam70, Lab79c, Mar73, Pie73, SS71, Ste79]. Radiation [FM79, BS75b, CE75, Hub71, Jon74, KK75, Lan70, SK79, Tor73, TW79]. Radiative [Rem72, EJN74, EG74, HKK73, IL74, SF78, WBG79]. radical [Bel72]. radii [Fun79a, Fun79b]. Radio [BS77, Sch78]. radioactive [Aur77]. radioastronomical [BS75b, Sal75]. Raman [MW70, MW71a, MW71b, MW71c]. RAMSES [DH75]. random [KCA76, Sch74, WDM77]. randomly [WD77]. range [BP77, FKU^{+}76, Hub70d, LR72, NK78, PV74, Ste79, TL74, TUWA79]. ranges [MRW76]. Rapid [Mue71, BKE79, GMR75b]. Rapidity [Jad75]. rate [Rob79c]. rates [TSD76]. rational [GMR75b]. ratios [MRCL76]. ray [AWG71, BCV^{+}72, Deu79a, Deu79b, Deu79c, Deu79d, FK73, FFH76, FC71, Hub73, Wil72, vM76b, Fer75]. Rayleigh [SS74]. rays [RSL^{+}78, SS74, SS76]. RCWFF [Bar76]. RCWFN [Bar76]. REACS [Rob79a]. reactance [BTS73, Hen75, Sar70, Sar72, Sar78]. reaction [CL79, Hod76, vdM72]. reactions [Ass75, Bet75, But78, FKU^{+}76, HIM73, Hei72, MT76, OVQT^{+}76, Rob79a, Rob79c, SM78a, Sei73, SMD71, SR73, SDH74, TCR71, TL74, TUWA79, WMC71]. Reactive [Con79, Dic79, LWSS79]. reactor [CHW^{+}78, CHR75b, KKN^{+}72]. Real [Bar76, AM78, BFSG74, Cam79, CB77, Gli78, Sch78, Taf75a]. real-symmetric [CB77]. real-time [Sch78, Taf75a]. real-time-System [Gli78]. reciprocal [HM71a]. Recognition [BEK^{+}72, Zan72, ZS77]. recoil [PV74]. recombination [BS77, Mag78]. recoupling [Bur70c, Gra73b]. Recursive [PB72b, SG76, Ort72]. REDUCE [Cal72, CB78]. Reduced [Bra78b, Gla78, Klo75b, BC72, BF71, Cha74, Rob73, Sax77a, Sax78, Bra78a, FS75b]. Reduction [Ugn78]. reference [Smi77]. refinement [LM71]. reflection [FK73, Fer75, KN73, PI78]. reflections [FFH76]. Regge [HIM73]. region [vdM72]. regions [BS77, Nor73a]. registered [OVQT^{+}76]. regularized [GVV79]. reinforcement [KM79a]. reinforcement-learning [KM79a]. relations [Han78d]. relative [BPW71, BPW72b]. Relativistic [Bra72b, Bra72a, GH78, Gra79, LCW71, Des75, EJN74, Gra73a, Gra76b, Lab79a, Mor78a, SS76, WB74, Yat71]. relativity [Pay76]. relaxation [CB77, RCL75]. reliable [Han78c, Han78e]. Remarks [Köl72b]. Renyi [Cha76]. representation [Cho70, Hub70b, KSSA78]. representations [Bos76b, MPS75, Net75, Wor73]. repulsion [BDD78]. request [Ano72l]. required [ST73]. research [Aig72, Cas72]. resistivity [Isl73]. resolution [MMK73, Sut72, VG74]. resonance [DG75, SDH74]. resonant [KK75]. resource [Gli78]. respect [JP78]. response [CHR75b, LLC76]. restart [HR74b]. results [ABMT78, BJ76]. retrieval [DSK72, DSK73, HM71b]. Review [AF72b, Ano79f, Bee73, Bra72c, Bur72a, Cam72, Chr75a, Del77, Gos78, Gra78, Hib79, Hoa70, Lai74, May74, McN71, Moo72, Mor78b, Nor73b, Pat73, Smi72, Smi77, Sut72, War73, Zac74, Fro73]. revised [Gra76b]. Richtmeyer [Chr75a]. rigid [KN79b]. Ring [TS75a, Wil78]. RKR [Tel73]. Role [Nad76]. roots [Bos76a]. rotation [BC72]. rotational [BB74, DT79, Hir73, HH75]. rotationally [BKE79]. Rotenberg [Bra72c, McN71]. routine [Gru79, HR74a, Isl73, McC75, Moo71a, Moo71b, Ste71, WS74, Zlo78, vM75a, vM75b, vM78a]. routines [BPW72c, Gar74]. Rovibrational [Hen75]. run [BTS74, DTE74, NA72]. Rurp [Rin79].

S [Bra72c, McN71]. safety [CHW^{+}78]. sampling [CL79, KVW70]. Sanderson [Hoa70]. Saxon [Cug73, DHD74, Hir73, HH75]. Scalar [Klu72b, Chi72, KS73, vM78b]. scale [AAB75, vM75c]. scaling [vM79b]. scattered [KPH78]. scattering [AT75, BJ76, BTS73, Bra72c, BBL79, CSP^{+}79, Cha73, Chi73a, Com78, CLS^{+}70, Cop74, Cop75a, Cop75b, DT79, Deu79a, Deu79b, Deu79c, Deu79d, EM74, EM76, Fun79a, Fun79b, Gro72, Han78d, KF71, MS78, Moo71c, Nes79, Nut73, Rei73, Sin70, Sko74, SS74, SS76, Smi70a, Smi70b, ST73, SR78, Ste79, TTW79a, TYCT73]. SCF [DESL79, KPH78, WB74]. scheduling [BP73]. scheme [PB73, Wil70]. schemes [FC71]. Schmidt [O’C74a]. school [Nad75]. SCHOONSCHIP [Str74, Ner78]. SCHOONSHIP [Str79]. Schrödinger [Has76, BG72, Bel72, HH79, HLS75, Kil79, LR71, Mar73, RA78, Wei72]. Schwinger [Moi73]. science [AAB75]. Sciences [Mor78b]. scientific [Aig72]. scientists [Del77]. screening [DH71]. Sea [Tom73]. Search [O’C74b, Smi70b]. searches [GMP76]. Second [Cam72, BCE79, Cha73, Dem72, GMP76, Hol79, Lew70b, Sil78b]. second-order [Cha73, Lew70b, Sil78b]. section [Bet75, SR73, Yos78]. sections [All72, BKE79, BTS73, BTS74, Com78, CLS^{+}70, Hei72, Hen75, LJLC72, MMM74, MT79, PV74, PS77, Sar72, Sar78, SMD71, vdB74]. Selection [HM71a]. Self [CCR76, Des70, LCW71, WD77]. Self-consistent [CCR76, LCW71]. self-diffusion [WD77]. selfgravitating [TW79]. semiconductors [NW79, RM76a, RM76b, RM76c, Ryc78]. sense [Bok78]. separation [MR73]. September [Ano79f]. sequence [HH79, Kno70]. Sequential [KCA76]. series [Del79b, Jam70, Ort72]. service [Ano74d, Ano74e, Ano76h]. set [BPW72c, Han78a, SV79]. sets [AN79, AZS^{+}79, Moo79a, Rei73]. SEURBNUK [CHW^{+}78]. SEURBNUK-2 [CHW^{+}78]. several [vM79a]. shapes [vM75a]. sharing [Gli78]. shasta [WBG79]. shell [AM74, BDD78, Chi73b, GA74, Hub70a, Klu72a, KL72, LW75, MNJL71, Sin70, KN79a]. shells [DHN79, DH71]. shift [NM73]. shifts [KF71, MS78, Ste79]. shock [RD79]. short [BP77, Ste79]. short-range [Ste79]. Shortley [Tob79]. showers [Goo74, Sta72]. SI [Cam75a]. Siegmund [Hib79]. signal [vM79b]. significant [O’C74b]. Simple [Sta72, Bos76a, Bos76b, GMR75b, vM76b]. simulate [vM77]. simulated [WDM77]. simulating [Rou72]. Simulation [DEW78, LVPG78, LNV78, Nor73b, OBLR73, BB74, Ber77, BMS72, Bun76, Chr74a, Chr74b, CKM^{+}73, DBB71, Daw72, Goo74, Gro72, Kas75, KM79a, LL72, LB78, LNS76, RM76a, RM76b, RM76c, Ryc78, Sch73, ST72, Sta72, SV79, Tom73, Tor73, VD74, WD77]. simulations [JB77, MGTD73, OC78]. SINAC [MP74]. sine [TS75b]. Single [Lew70b, Hub70a, MT76, SDH74]. Single-particle [Lew70b]. single-particle-inclusive [MT76]. SIPSOL [Jes79]. skin [TT78]. SLAB [Cop75a]. Slater [AT75, Des70, DF76, FP73, Gol78, Lab79d, ST73, WB74, Yur78]. Slater-Transform [Yur78]. Slater-type [AT75, DF76, ST73]. Slit [Deu79a, Deu79b, Deu79c, Deu79d]. small [Aur77, DTE74, Deu79a, Deu79b, Deu79c, Deu79d, ETH73]. smearing [Deu79a, Deu79b, Deu79c, Deu79d]. Smith [Moo72]. smooth [CL79]. Society [Bis75]. sodium [ÖW70]. Software [BPZ73, Kar78, For78, HR74a, Taf75b]. Solid [Bee73, JB77, KKN^{+}72, SK79]. solid-state [KKN^{+}72]. solids [HRB72, JVS72]. soliton [Bog77]. Solitons [Zab73]. Solution [Hug71, LW75, Par73, Bel72, Bor76, CE75, Cha75, Chi73a, Col77, CT79, CSW78, De 77, EJ77, Fle72, HH79, HD78, HKK73, Jes79, Jon72, KS75a, LNV78, LR71, Mar73, MAK^{+}72, RA78, TA72, Wei72]. solutions [Has76, Sch72, SK79]. solve [TSD76]. solver [CH71a]. solving [Cha73, CB77, EG74, Gru75, KS79, PPR72, Sta72, WBG79]. Some [Bog77, Kil79, SSZ73, TA72, Sch72, Zhi72]. Sons [Del77, Gra78]. sound [Mor72]. source [Hub71, Pie73, Sal75]. sources [BS75b]. space [DE72, ET76, Jab73, Jad75, KK71, Mor78a, Net75, PB72b]. space-charge [DE72]. spaced [AAMB79, CM79]. spark [BCV^{+}72]. speakeasy [Coh71]. special [De 77, LS79, vdSVSH73]. specialist [Ano74j]. species [Rob79b]. spectra [AK76, AWG71, BB74, Bet75, CT71, Gro72, KE72, KE74, MGTD73, VDM78, War78, WS74, ZKU79, Zlo79, Sut72]. Spectral [Lan70, SLH78, vM75b]. Spectro [Zlo78]. Spectro-oriented [Zlo78]. spectrometer [LLC76]. spectrometry [Sou73]. spectroscopic [FS75b, HEL71, Lab79a, Sax77b, vM75a]. spectroscopy [FC74]. spectrum [BGT76, Sal75, SR78, VG74, vM75c]. speed [SC71]. spheric [Chr74a].

Spherical [AM78, Bra73, Del79b, HKK73, MW71b, MW71c]. spherically [Joh73]. spin [CB78, Ell75, Hei72, Hub70a, KF71, Klo75a, MGTD73, SM78a, Sax77b, SDH74, Ste76, DG75]. spin-Hamiltonian [Ste76]. spin-orbit [Ell75, Klo75a, Sax77b]. spin-up [CB78]. spline [AAMB79, Cha75, CM79]. splines [Bok78, OC78, PS78]. Springer [Cam72, Chr75a, Gos78, Mor78b, Sut72]. Springer-Verlag [Cam72, Chr75a, Gos78, Mor78b, Sut72]. Spying [Taf75a]. squares [Bok78, Han78f, LM71, O’C74a, Pac72, RFSG72b, WS74, vM75b]. SSOR [Cam75a]. SSOR-SI [Cam75a]. stability [ABG^{+}75, CMS77, Fro72, GJ76]. standard [CR74, Sal76]. star [BM76, HEL71, HE71, HE73]. stars [Gra76a, Mik74]. State [Bee73, And70, Cug73, DeV76, FCW78, KKN^{+}72, KH78, LW75, LR71, Mor73, Sch73, Tal76]. statements [Sal76]. states [Bra72b, Bra72a, BF71, CCR79, DHD74, Fro73, Hir73, HH75, Lab79a, SM78a, Tal76, WS79]. Static [Rin79, FT71, Gia76, Han78b]. stationary [BB79]. Statistical [GR78, Ald72, CE75, CIT73, Hib79]. steady [FCW78, Tal76]. stellar [Bou72]. stereographic [Pre79b]. stimulated [LVPG78]. Stirling [BPW72c]. STLPLT [CA70]. storage [Nas74, Tob79, HM71b]. straggling [CJ71, JPB78]. straight [BGT76, GPS75, Pet76a, PS77]. strategy [TS75b]. streamer [EGMS72]. strengths [Cla78, Duf77, FS75c, God78, Hib75, Sar70]. strong [KK77]. strong-field [KK77]. strongly [vM75b]. structure [BB74, Cro72, Fro73, GH76, Gra73a, Gra76b, Gra79, Han78b, Han78d, Hib70, Hib71, Hib74b, Hib74a, Hof71, HM71a, HM71b, HM71c, HRB72, KN79b, KN79a, Mik74, Sar72, Sar78]. structure-matching [Hof71]. structured [Nad76]. Structures [Smi72, EJN74]. studies [CHW^{+}78, Sch73, Tor73]. study [BP77, Cha73, CL79, DT79, FS73, RM76a, RM76b, RM76c, Sou73, Tom73]. SU3 [MNJL71]. sub [Smi78]. sub-millimeter [Smi78]. subject [HE73]. SUBMMW [Smi78]. subprograms [Jes79]. Subroutine [Sta76, Bok78, GMR75b, Mor73, Smi70a]. subroutines [SV79]. subscription [Ano72a]. substitution [Lew70b]. Subtracted [MP74]. suite [Jes79]. suited [Aue78]. summation [CT79]. summations [Klo75b, Rob73]. Summer [Nad75]. sums [Lew70b]. super [Nil72]. super-heavy [Nil72]. surface [Ugn78]. surfaces [Nil72]. survey [Fos72]. susceptibility [vM78a]. switching [Ryc78]. symbol [FP73]. Symbolic [BKK76, GVV79, BH79, KS75a, Str74, KPPR76]. symbols [Cra72, Sha70]. Symmetric [Hub70c, Hub70d, Cam75b, CB77, DEW78, Joh73, MW70, MW71a, SM78b, SF78, Ugn78]. Symmetrization [RS75]. Symmetrized [HRB72, HM71c]. Symmetry [De 75b, HM71a, HM71b, HM71c, BDD78, PI78, Ste76]. Symposium [AF72b, Zac74]. synchrotron [Lan70]. System [Gli78, ADF^{+}72, AZS^{+}79, BP73, BEK^{+}72, CMW74, Coh71, EK72, EK73, JR75, KKN^{+}72, MR77, PB73, Pen78, Rob74]. Systems [AAB75, Bou72, CIT73, CLS^{+}70, CHR75b, CH74, Kar78, KM79a, KH78, KS79, Moo71c, ST72, Smi77, Sou73, TA72, Zac72].

T [Bee73, CL79, Smi72, Blo72]. T-junction [Blo72]. tableaux [SM78b]. tail [AL79]. tailor [RD79]. tailor-made [RD79]. take [Rob79a]. Talmi [Zoh72]. tapes [Gol70]. target [CW79b, Hub71, MS78]. targets [KF71]. teaching [BS75a]. technique [Pac72, Tro73]. Techniques [EJN74, Nad75, Rei79, Rob71, Vel72, Pat73]. temperature [MR77]. Tensor [FS75b, Cha74, Gla78, Klo75b, Klu72b, LB76, LR71, LR72, Pay72, Rob73, Sax77a]. terminal [MMK73]. terms [BB77, Lab79b, RD79]. Test [Gra71, RCL75]. tests [Cha76, GR78]. tetragonal [Ste76]. text [Day77]. THALIA [ABG^{+}75]. their [War73]. theoretic [EJ77]. theoretical [ABMT78, Bur72a, EK72, HM71b, Smi78, WW74, WW72]. Theory [Smi72, BCE79, CE75, CSW78, Daw72, Joa73, Kil79, Lew70b, MS72, MS78, PW73, ST73, WP73]. thermal [BS75b, Cop74, Cop75a, Cop75b, Sal75]. thermally [LVPG78]. thermodynamic [FS73]. thick [KK75]. thin [Hub70c, Hub70d, Sko74]. Third [Ano79f, JS72, Sil78b, Wil78]. third-order [Sil78b, Wil78]. three [Ass75, BDD78, Bra72a, MR73, RR74, Aue78, RM76c]. three-body [Ass75, Bra72a]. three-dimensional [RR74, RM76c]. three-open-shell [BDD78]. threshold [CL79, Ryc78]. tidal [HE73]. time [Gli78, Has76, Mag78, MPW73, Rob79b, Sch78, Taf75a, Wei72]. time-dependent [Has76, Mag78, MPW73, Wei72]. Timer [HR74a]. timing [HR74a]. TLASER [DST75]. Tokamak [MBW76]. Topics [RC72]. tops [MW70, MW71a, MW71b, MW71c]. toroidal [Mik74, ST72]. total [SMD71, SR73, WB74]. town [Tom73]. track [BSP71]. trajectories [BS75a, HIM73, PS76]. Trajectory [vdM72]. Tranal [KPPR76]. Transfer [KK75, But78, EG74, FKU^{+}76, HKK73, Mal70, PV74]. Transform [CH71b, Mar76, Yur78]. transformation [SZ79, SG72]. transformations [O’C74a]. transforms [Col77, CT79, EJ77, TS75b]. transition [BTS73, BF71, FS73, GPS75, MR77, PS76, PS77, Wil70]. translation [KPPR76]. translations [SMD71]. transmission [Deu79c, Deu79d]. Transport [OS71, Tho79, Cha75, KK75, MBW76, NA72, Sch72, TSD76, Zab73]. transportable [For78]. trapezoidal [CT79]. treatment [GPS75, PS76, PS77]. Trip [Mag78]. triply [MW71c]. tube [RD79]. turbulent [ATV76]. twin [SDH74]. Two [LL72, Taf75b, Blo72, Bra72b, Cho70, Chr74b, CH71a, CW79b, CKM^{+}73, DH75, EG74, FKU^{+}76, HIM73, Jam70, LR71, LR72, MPW73, WBG79, ZKU79, Zlo79, dJ78, RM76b]. two-body [Bra72b, HIM73]. Two-dimensional [LL72, Blo72, Cho70, CH71a, CW79b, DH75, EG74, MPW73, WBG79, Zlo79, RM76b]. two-layered [Chr74b]. Two-level [Taf75b]. two-nucleon [FKU^{+}76, LR71, LR72]. type [AT75, Che73, DF76, DH75, Mat72, SMD71, SR73, ST73]. typewritter [MMK73].

ultrasonic [LNV78]. unequally [AAMB79, CM79]. unfolding [Rou72]. unitary [ADO78, TS78]. University [Pat73, BS75a]. unmagnetized [TT78]. unsteady [Dem78, RD79]. UPEAK [Zlo78]. Use [AN79, GMP76, Nut73, Pay76, Rob72, Web73]. used [CSP^{+}79]. user [AD73]. Users [EK73, KKN^{+}72]. using [AT75, All72, ABG^{+}75, BBC^{+}74, BBL^{+}78, BCE79, CB78, Cra72, EJ77, FP73, Hof71, KN79b, MS78, Nas74, NW79, O’C74a, Pre79a, Rei73, RM76a, RM76b, RM76c, ST73]. utility [CR74, HRL75]. utilizing [Dem78].

V103 [HW77]. vacancies [WD77, WDM77]. vacuum [CS78]. validity [JPB78]. value [CR74, Pas72]. values [Hub70a, RFSG72a, RFSG72b]. variable [HEL71, HE73]. variables [Mor78a, O’C74b, dJ78, vM79a]. variation [Rob79b]. variational [AT75, DHD74, Moi73, Nes73, Nes79]. Vavilov [JPB78, Sch74]. Vector [DW72, Nus70, CN73, Klu72b, KS75c]. vectors [HM71a]. Venice [Tom73]. Verlag [Cam72, Chr75a, Gos78, Mor78b, Sut72]. versatile [KE72, Sar78]. version [BBL^{+}78, Chi73a, Chi73b, DTE74, FS75b, Fun79a, Fun79b, Gia76, Gra76b, Hib71, KE74, LS79, Moo71a, Sax77a, Sax77b, Sax78, VDM78, War78, WW74]. via [WD77]. Vibrational [Mag75, TSD76]. vibrationally [BKE79]. vibrations [WW72]. virial [JS72, Joh73, TS75a]. viscosities [Mor72]. viscous [Blo72]. visual [AMB75]. visually [ABMT78]. Vol [McN71, Mor78b, Sut72, Bra72c]. Volume [Ano70h, Ano70i, Ano71d, Ano71e, Ano71f, Ano72e, Ano72f, Ano73c, Ano73d, Ano73e, Ano73o, Ano74c, Ano74g, Ano74q, Ano75b, Ano75a, Ano75f, Ano75e, Ano75s, Ano75r, Ano76b, Ano76c, Ano76f, Ano76e, Ano76p, Ano76q, Ano78a, Ano78b, Ano78c, Ano78f, Ano78g, Ano78h, Ano78r, Ano78s, Ano79c, Ano79d, Ano79e, Ano79g, Ano79h, Ano79i, Ano79p, Ano79q, Ano79r, KK71]. volumes [Ano78t]. Vooren [Gos78].

W [Bee73]. walls [Hub70c, Hub70d]. water [CH74]. Wave [RFSG72b, She76, And70, ATV76, BFSG74, BKK76, CHMM73, Cug73, Has76, Hib75, Jam70, KPH78, Sal78, She74, Smi78, Ste79, TCR71]. wave-functions [Sal78]. wave-particle [ATV76]. Wavefunction [Bar76, ABH^{+}72, Kla71]. Wavefunctions [Köl72b, DeV76, Kno70]. weakly [ATV76]. Weight [ADO78]. Weizmann [GA74]. well [DHD74]. Weyl [BH72, BK72]. which [Hei72, Rob79a, dJ78]. Wigner [AD73, CMS77]. Wiley [Del77, Gra78, Lai74, May74, Moo72]. Wilson [CSC74]. within [Bet75, Rob79b]. without [DH71, HD78]. Woods [Cug73, DHD74, Hir73, HH75]. work [KM74]. Worker [BEM79]. writer [Bar77]. WWR [KKN^{+}72]. WWR-M [KKN^{+}72].

X [Bra72c, CL79, KPH78, CL79, Deu79a, Deu79b, Deu79c, Deu79d, FK73, Fer75, FFH76, vM76b]. X-ray [Fer75, Deu79a, Deu79b, Deu79c, Deu79d, FK73, FFH76, vM76b]. XF [CL79]. xi [Gra78].

York [AF72b].

Zanbergen [Gos78]. Zeeman [DG75]. zero [dJ78]. zeros [CSP^{+}79]. zincblende [KN79b, KN79a].

References

Ayoub:1975:SAL

[AAB75] E. E. Ayoub, R. J. Ascuitto, and D. A. Bromley. Systems aspects of large scale nuclear science programming. Computer Physics Communications, 10(4):203–222, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900661.

Allard:1972:HFN

[AAH72] J. F. Allard, A. Abzouzi, and B. Houssais. Hartree–Fock nuclear calculations with Gaussian potentials. Computer Physics Communications, 3(1):22–30, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900501.

Aldeen:1976:CEE

[AAJ76] I. H. Aldeen, A. C. Allison, and M. J. Jamieson. The calculation of eigenvalues and eigenfunctions in an asymptotically Coulomb potential. Computer Physics Communications, 12(3):261–265, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900850.

Anderson:1979:CSI

[AAMB79] J. Anderson, R. W. B. Ardill, K. J. M. Moriarty, and R. C. Beckwith. A cubic spline interpolation of unequally spaced data points. Computer Physics Communications, 16(2):199–206, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900882.

Appert:1975:TOD

[ABG^{+}75] K. Appert, D. Berger, R. Gruber, F. Troyon, and K. V. Roberts. THALIA — A one-dimensional magnetohydrodynamic stability program using the method of finite elements. Computer Physics Communications, 10(1):11–29, July 1975. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900879.

Allard:1972:NHF

[ABH^{+}72] J. Allard, N. Boumahrat, B. Houssais, M. Hadj Hassan, and M. Lambert. A nuclear Hartree–Fock intrinsic wavefunction projection program. Computer Physics Communications, 4(2):239– 248, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557290015X.

Anderson:1978:PPV

[ABMT78] J. Anderson, R. C. Beckwith, K. J. M. Moriarty, and J. H. Tabor. A plotting package for visually comparing theoretical and experimental results. Computer Physics Communications, 15(5):437– 441, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900723.

Akiyama:1973:UGF

[AD73] Yoshimi Akiyama and J. P. Draayer. A user’s guide to Fortran programs for Wigner and Racah coefficients of SU_{3}. Computer Physics Communications, 5(6):405–406, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900775.

Akolk:1972:DLS

[ADF^{+}72] F. Akolk, H. Dilcher, H. Frese, G. Hochweller, P. Kuhlmann, and E. Raubold. Desy on-line system. Computer Physics Communications, 4(3):275–278, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900872.

Amar:1978:WMU

[ADO78] V. Amar, U. Dozzio, and C. Oleari. Weight multiplicity for unitary groups. Computer Physics Communications, 14(5–6):413–421, July/August 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557890005X.

Abecasis:1972:ELC

[AF72a] S. M. Abecasis and F. R. Femenia. Energy level calculations with AROVMI model. Computer Physics Communications, 4(2):262– 267, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900185.

Allison:1972:BRB

[AF72b] D. C. S. Allison and D. Q. M. Fay. Book review: Proceedings of the Symposium on Computer Processing in Communications, New York 1969: ed. J. Fox, Polytechnic Press £9.50. Computer Physics Communications, 3(3):273, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900744.

Abecasis:1974:ELC

[AF74] S. M. Abecasis and F. R. Femenia. Energy-level calculations with the extended AROVMI model. Computer Physics Communications, 7(3):145–150, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900046.

Abecasis:1971:ELC

[AFH71] S. M. Abecasis, F. R. Femenia, and E. S. Hernandez. Energy level calculations in Davydov model. Computer Physics Communications, 2(1):33–39, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900129.

Ahlrichs:1979:MBP

[Ahl79] Reinhart Ahlrichs. Many body perturbation calculations and coupled electron pair models. Computer Physics Communications, 17(1–2):31–45, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900675.

Aigrain:1972:CSR

[Aig72] P. Aigrain. Computing and scientific research in Europe. Computer Physics Communications, 3(S1):166–173, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572901257.

Assimakopoulos:1976:AIP

[AK76] P. A. Assimakopoulos and S. Kossionides. Anna: An interactive program for analysis of one-dimensional pulse-height spectra. Computer Physics Communications, 11(1):37–48, January/ February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900382.

Aquilanti:1979:CEP

[AL79] V. Aquilanti and A. Laganà. On the computation of eigenenergies for potentials with a Coulomb tail. Computer Physics Communications, 17(1–2):113–116, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900742.

Alder:1972:NES

[Ald72] B. J. Alder. Numerical experiments in statistical mechanics. Computer Physics Communications, 3(S1):86–91, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572901178.

Allison:1972:CAE

[All72] Arthur C. Allison. The calculation of absorption and elastic cross sections using the optical potential. Computer Physics Communications, 3(3):173–179, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900665.

Allison:1974:FPC

[AM74] D. C. S. Allison and J. E. McNulty. Fractional parentage coefficients for equivalent f shell electrons. Computer Physics Communications, 8(3):246–256, October 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574901003.

Ardill:1977:BFC

[AM77] R. W. B. Ardill and K. J. M. Moriarty. The Bessel functions J_{0} and J_{1} of complex argument. Computer Physics Communications, 13 (1):17–24, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900236.

Ardill:1978:SBF

[AM78] R. W. B. Ardill and K. J. M. Moriarty. Spherical Bessel functions j_{n} and y_{n} of integer order and real argument. Computer Physics Communications, 14(3–4):261–265, May/June 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046557890019X.

Ardill:1979:ABF

[AM79] R. W. B. Ardill and K. J. M. Moriarty. Accurate Bessel functions J_{n}(z), Y_{n}(z), H_{n}^{(1)}(z) and H_{n}^{(2)}(z) of integer order and complex argument. Computer Physics Communications, 17(3):321–336, June 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Anderson:1975:PPV

[AMB75] J. Anderson, K. J. M. Moriarty, and R. C. Beckwith. A plotting package for visual comparison of points and curves. Computer Physics Communications, 9(2):85–91, February 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900259.

Ahmad:1979:UDB

[AN79] S. Ahmad and D. J. Newman. Use of discrete basis sets in configuration interaction calculation. Computer Physics Communications, 18(3):331–337, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900031.

Andre:1970:PGP

[And70] Jean-Marie Andre. Polymol: a general program for the calculation of ground-state wave functions of polymers. Computer Physics Communications, 1(6):391–414, December 1970. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465570900135.

Anonymous:1970:C

[Ano70a] Anonymous. Contents. Computer Physics Communications, 1(6): 478, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900287.

Anonymous:1970:CPC

[Ano70b] Anonymous. The C.P.C. program library. Computer Physics Communications, 1(6):473–476, December 1970. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900263.

Anonymous:1970:Ea

[Ano70c] Anonymous. Erratum. Computer Physics Communications, 1(3): 224, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557090010X.

Anonymous:1970:Eb

[Ano70d] Anonymous. Erratum. Computer Physics Communications, 1(4): 291, April 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900457.

Anonymous:1970:Ec

[Ano70e] Anonymous. Erratum. Computer Physics Communications, 1(6):470–472, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900251.

Anonymous:1970:ENa

[Ano70f] Anonymous. Erratum notice. Computer Physics Communications, 1(3):223, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900093.

Anonymous:1970:ENb

[Ano70g] Anonymous. Erratum notices. Computer Physics Communications, 1(6):469, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557090024X.

Anonymous:1970:VAI

[Ano70h] Anonymous. Volume 1 author index. Computer Physics Communications, 1(6):481–482, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900305.

Anonymous:1970:VPI

[Ano70i] Anonymous. Volume 1 program index. Computer Physics Communications, 1(6):479–480, December 1970. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900299.

Anonymous:1971:EB

[Ano71a] Anonymous. Editorial board. Computer Physics Communications, 2(1):??, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900075.

Anonymous:1971:ENa

[Ano71b] Anonymous. Erratum notices. Computer Physics Communications, 2(3):173–174, April 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557190049X.

Anonymous:1971:ENb

[Ano71c] Anonymous. Erratum notices. Computer Physics Communications, 2(7):471–472, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900415.

Anonymous:1971:VAI

[Ano71d] Anonymous. Volume 2 author index. Computer Physics Communications, 2(7):476–477, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900440.

Anonymous:1971:VC

[Ano71e] Anonymous. Volume 2 contents. Computer Physics Communications, 2(7):473, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900427.

Anonymous:1971:VPI

[Ano71f] Anonymous. Volume 2 program index. Computer Physics Communications, 2(7):474–475, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900439.

Anonymous:1972:ASP

[Ano72a] Anonymous. Application for a subscription to the program library. Computer Physics Communications, 3(2):??, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557290063X.

Anonymous:1972:AIa

[Ano72b] Anonymous. Author index. Computer Physics Communications, 3(S1):??, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901270.

Anonymous:1972:AIb

[Ano72c] Anonymous. Author index. Computer Physics Communications, 3(4):354–355, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900860.

Anonymous:1972:AIc

[Ano72d] Anonymous. Author index. Computer Physics Communications, 4(3):387–389, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901051.

Anonymous:1972:CVa

[Ano72e] Anonymous. Contents of volume 3. Computer Physics Communications, 3(4):351, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900847.

Anonymous:1972:CVb

[Ano72f] Anonymous. Contents of volume 4. Computer Physics Communications, 4(3):385, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901038.

Anonymous:1972:EBa

[Ano72g] Anonymous. Editorial board. Computer Physics Communications, 3(1):??, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557290046X.

Anonymous:1972:EBb

[Ano72h] Anonymous. Editorial Board. Computer Physics Communications, 3(S1):??, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901063.

Anonymous:1972:EBc

[Ano72i] Anonymous. Editorial Board. Computer Physics Communications, 4(1):??, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900203.

Anonymous:1972:E

[Ano72j] Anonymous. Erratum. Computer Physics Communications, 4(3):382–383, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901014.

Anonymous:1972:EN

[Ano72k] Anonymous. Erratum notices. Computer Physics Communications, 3(3):275–276, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900768.

Anonymous:1972:IRO

[Ano72l] Anonymous. Individual request order form. Computer Physics Communications, 3(2):??, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900641.

Anonymous:1972:LP

[Ano72m] Anonymous. List of participants. Computer Physics Communications, 3(S1):ix–xvi, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901099.

Anonymous:1972:OC

[Ano72n] Anonymous. Organizing committee. Computer Physics Communications, 3(S1):viii, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901087.

Anonymous:1972:PIa

[Ano72o] Anonymous. Program index. Computer Physics Communications, 3(4):352–353, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900859.

Anonymous:1972:PIb

[Ano72p] Anonymous. Program index. Computer Physics Communications, 4(3):386, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557290104X.

Anonymous:1973:A

[Ano73a] Anonymous. Announcement. Computer Physics Communications, 6(5):243, November 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900969.

Anonymous:1973:AI

[Ano73b] Anonymous. Author index. Computer Physics Communications, 6(6):380–381, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900489.

Anonymous:1973:AIV

[Ano73c] Anonymous. Author index volume 5. Computer Physics Communications, 5(6):481–483, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900842.

Anonymous:1973:CVa

[Ano73d] Anonymous. Contents volume 5. Computer Physics Communications, 5(6):479, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900830.

Anonymous:1973:CVb

[Ano73e] Anonymous. Contents volume 6. Computer Physics Communications, 6(6):377–378, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900465.

Anonymous:1973:EBa

[Ano73f] Anonymous. Editorial Board. Computer Physics Communications, 5(1):??, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900015.

Anonymous:1973:EBb

[Ano73g] Anonymous. Editorial Board. Computer Physics Communications, 6(1):??, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900155.

Anonymous:1973:Ea

[Ano73h] Anonymous. Erratum. Computer Physics Communications, 5(4): 308, April 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557390074X.

Anonymous:1973:Eb

[Ano73i] Anonymous. Erratum. Computer Physics Communications, 6(2): 98, August 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900581.

Anonymous:1973:ENa

[Ano73j] Anonymous. Erratum notice. Computer Physics Communications, 5(4):304–307, April 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900738.

Anonymous:1973:ENc

[Ano73k] Anonymous. Erratum notice. Computer Physics Communications, 6(1):59–61, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900222.

Anonymous:1973:ENb

[Ano73l] Anonymous. Erratum notices. Computer Physics Communications, 5(5):395–396, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900660.

Anonymous:1973:P

[Ano73m] Anonymous. Preface. Computer Physics Communications, 6(6): 245, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900337.

Anonymous:1973:PI

[Ano73n] Anonymous. Program index. Computer Physics Communications, 6(6):379, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900477.

Anonymous:1973:PIV

[Ano73o] Anonymous. Program index volume 1–5. Computer Physics Communications, 5(6):484–490, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900854.

Anonymous:1974:A

[Ano74a] Anonymous. Announcement. Computer Physics Communications, 8(1):??, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900770.

Anonymous:1974:AI

[Ano74b] Anonymous. Author index. Computer Physics Communications, 7(7):436–439, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900757.

Anonymous:1974:AIV

[Ano74c] Anonymous. Author index volume 8. Computer Physics Communications, 8(5):409–411, December 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900174.

Anonymous:1974:CPCa

[Ano74d] Anonymous. Computer Physics Communications program library : New service charges—January 1975. Computer Physics Communications, 8(2):??, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900654.

Anonymous:1974:CPCb

[Ano74e] Anonymous. Computer Physics Communications program library new service charges—January 1975. Computer Physics Communications, 8(4):??, November 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900241.

Anonymous:1974:C

[Ano74f] Anonymous. Contents. Computer Physics Communications, 7(7): 429–431, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900733.

Anonymous:1974:CV

[Ano74g] Anonymous. Contents volume 8. Computer Physics Communications, 8(5):404–405, December 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900150.

Anonymous:1974:EBa

[Ano74h] Anonymous. Editorial Board. Computer Physics Communications, 7(1):??, January 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900526.

Anonymous:1974:EBb

[Ano74i] Anonymous. Editorial Board. Computer Physics Communications, 8(1):??, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900769.

Anonymous:1974:ESF

[Ano74j] Anonymous. Editors in specialist fields. Computer Physics Communications, 8(1):xi–xii, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900800.

Anonymous:1974:E

[Ano74k] Anonymous. Erratum. Computer Physics Communications, 7(3): 177, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900101.

Anonymous:1974:ENa

[Ano74l] Anonymous. Erratum notice. Computer Physics Communications, 7(4):236, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900940.

Anonymous:1974:ENb

[Ano74m] Anonymous. Erratum notice. Computer Physics Communications, 8(2):141, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900642.

Anonymous:1974:ENc

[Ano74n] Anonymous. Erratum notice. Computer Physics Communications, 8(4):333–336, November 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557490023X.

Anonymous:1974:IA

[Ano74o] Anonymous. Instructions to authors. Computer Physics Communications, 8(1):v–x, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900794.

Anonymous:1974:PI

[Ano74p] Anonymous. Program index. Computer Physics Communications, 7(7):432–435, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900745.

Anonymous:1974:PIV

[Ano74q] Anonymous. Program index volume 8. Computer Physics Communications, 8(5):406–408, December 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900162.

Anonymous:1975:AIVb

[Ano75a] Anonymous. Author index of volume 10. Computer Physics Communications, 10(6):444–446, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900508.

Anonymous:1975:AIVa

[Ano75b] Anonymous. Author index of volume 9. Computer Physics Communications, 9(6):413–415, June 1975. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900223.

Anonymous:1975:CEC

[Ano75c] Anonymous. Centre Européen de Calcul Atomique et Moléculaire — activities in 1976. Computer Physics Communications, 10(5): 341, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575901010.

Anonymous:1975:CPC

[Ano75d] Anonymous. Computer Physics Communications program library. Computer Physics Communications, 10(3):203, September 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900077.

Anonymous:1975:CVb

[Ano75e] Anonymous. Contents of volume 10. Computer Physics Communications, 10(6):439–440, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557590048X.

Anonymous:1975:CVa

[Ano75f] Anonymous. Contents of volume 8. Computer Physics Communications, 9(6):408–409, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557590020X.

Anonymous:1975:EBa

[Ano75g] Anonymous. Editorial Board. Computer Physics Communications, 9(1):??, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557590051X.

Anonymous:1975:EBb

[Ano75h] Anonymous. Editorial Board. Computer Physics Communications, 10(1):??, July 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900855.

Anonymous:1975:E

[Ano75i] Anonymous. Erratum. Computer Physics Communications, 10(6): 438, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900478.

Anonymous:1975:ENa

[Ano75j] Anonymous. Erratum notice. Computer Physics Communications, 9(2):129, February 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900296.

Anonymous:1975:ENb

[Ano75k] Anonymous. Erratum notice. Computer Physics Communications, 9(4):268–269, April 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900132.

Anonymous:1975:ENc

[Ano75l] Anonymous. Erratum notice. Computer Physics Communications, 10(1):70, July 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900922.

Anonymous:1975:ENd

[Ano75m] Anonymous. Erratum notice. Computer Physics Communications, 10(1):71, July 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900934.

Anonymous:1975:ENe

[Ano75n] Anonymous. Erratum notice. Computer Physics Communications, 10(4):251–256, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900703.

Anonymous:1975:ENf

[Ano75o] Anonymous. Erratum notice. Computer Physics Communications, 10(4):257, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900715.

Anonymous:1975:ENg

[Ano75p] Anonymous. Erratum notice. Computer Physics Communications, 10(6):434–435, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900454.

Anonymous:1975:ENh

[Ano75q] Anonymous. Erratum notice. Computer Physics Communications, 10(6):436–437, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900466.

Anonymous:1975:PIVb

[Ano75r] Anonymous. Program index of volume 10. Computer Physics Communications, 10(6):441–443, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900491.

Anonymous:1975:PIVa

[Ano75s] Anonymous. Program index of volume 9. Computer Physics Communications, 9(6):410–412, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900211.

Anonymous:1976:A

[Ano76a] Anonymous. Announcement. Computer Physics Communications, 12(2):259, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900734.

Anonymous:1976:AIVa

[Ano76b] Anonymous. Author index of volume 11. Computer Physics Communications, 11(3):409–411, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046557690028X.

Anonymous:1976:AIVb

[Ano76c] Anonymous. Author index of volume 12. Computer Physics Communications, 12(3):345–347, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900953.

Anonymous:1976:CPC

[Ano76d] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 11(1):??, January/ February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900333.

Anonymous:1976:CVb

[Ano76e] Anonymous. Content of volume 12. Computer Physics Communications, 12(3):341–342, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557690093X.

Anonymous:1976:CVa

[Ano76f] Anonymous. Contents of volume 11. Computer Physics Communications, 11(3):412–413, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900291.

Anonymous:1976:CPLa

[Ano76g] Anonymous. CPC program library. Computer Physics Communications, 11(1):??, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900321.

Anonymous:1976:CPLb

[Ano76h] Anonymous. CPC program library — new service charges, January 1977. Computer Physics Communications, 12(1):??, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900163.

Anonymous:1976:EBa

[Ano76i] Anonymous. Editorial Board. Computer Physics Communications, 11(1):??, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557690031X.

Anonymous:1976:EBb

[Ano76j] Anonymous. Editorial Board. Computer Physics Communications, 12(1):??, September/October 1976. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900011.

Anonymous:1976:ENa

[Ano76k] Anonymous. Erratum notice. Computer Physics Communications, 11(3):407, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900278.

Anonymous:1976:ENb

[Ano76l] Anonymous. Erratum notice. Computer Physics Communications, 12(3):339–340, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900928.

Anonymous:1976:IA

[Ano76m] Anonymous. Instructions to authors. Computer Physics Communications, 11(1):??, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900345.

Anonymous:1976:LP

[Ano76n] Anonymous. List of participants. Computer Physics Communications, 12(1):ix–x, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900047.

Anonymous:1976:OPC

[Ano76o] Anonymous. Organizing and programme committees. Computer Physics Communications, 12(1):viii, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900035.

Anonymous:1976:PIVa

[Ano76p] Anonymous. Program index of volume 11. Computer Physics Communications, 11(3):414–416, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465576900308.

Anonymous:1976:PIVb

[Ano76q] Anonymous. Program index of volume 12. Computer Physics Communications, 12(3):343–344, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900941.

Anonymous:1977:CPC

[Ano77a] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 13(1):v–vii, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465577900182.

Anonymous:1977:EB

[Ano77b] Anonymous. Editorial Board. Computer Physics Communications, 13(1):??, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900170.

Anonymous:1977:ENa

[Ano77c] Anonymous. Erratum notice. Computer Physics Communications, 13(1):71, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900297.

Anonymous:1977:ENb

[Ano77d] Anonymous. Erratum notice. Computer Physics Communications, 13(1):72, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900303.

Anonymous:1977:ENc

[Ano77e] Anonymous. Erratum notice. Computer Physics Communications, 13(2):137–140, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900388.

Anonymous:1977:ENd

[Ano77f] Anonymous. Erratum notice. Computer Physics Communications, 13(2):141–145, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557790039X.

Anonymous:1977:ENf

[Ano77g] Anonymous. Erratum notice. Computer Physics Communications, 13(4):295, November/December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557790008X.

Anonymous:1977:ENe

[Ano77h] Anonymous. Erratum notices. Computer Physics Communications, 13(3):225–230, September/October 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465577900169.

Anonymous:1977:IA

[Ano77i] Anonymous. Instructions to authors. Computer Physics Communications, 13(1):ix–xiv, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900194.

Anonymous:1978:AIVa

[Ano78a] Anonymous. Author index to volume 12. Computer Physics Communications, 13(5–6):435–439, January/February 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900437.

Anonymous:1978:AIVb

[Ano78b] Anonymous. Author index to volume 14. Computer Physics Communications, 14(5–6):451–453, July/August 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900097.

Anonymous:1978:AIVc

[Ano78c] Anonymous. Author index to volume 15. Computer Physics Communications, 15(5):447–450, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900759.

Anonymous:1978:CPCa

[Ano78d] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 15(1–2):v–vii, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900784.

Anonymous:1978:CPCb

[Ano78e] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 16(1):v–vii, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578901017.

Anonymous:1978:CVa

[Ano78f] Anonymous. Contents to volume 12. Computer Physics Communications, 13(5–6):440–442, January/February 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900449.

Anonymous:1978:CVb

[Ano78g] Anonymous. Contents to volume 14. Computer Physics Communications, 14(5–6):449–450, July/August 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900085.

Anonymous:1978:CVc

[Ano78h] Anonymous. Contents to volume 15. Computer Physics Communications, 15(5):445–446, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900747.

Anonymous:1978:EBa

[Ano78i] Anonymous. Editorial Board. Computer Physics Communications, 14(1–2):v–vii, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900462.

Anonymous:1978:EBb

[Ano78j] Anonymous. Editorial Board. Computer Physics Communications, 15(1–2):??, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900772.

Anonymous:1978:EBc

[Ano78k] Anonymous. Editorial Board. Computer Physics Communications, 16(1):??, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578901005.

Anonymous:1978:ENa

[Ano78l] Anonymous. Erratum notice. Computer Physics Communications, 13(5–6):429, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900401.

Anonymous:1978:ENb

[Ano78m] Anonymous. Erratum notice. Computer Physics Communications, 14(3–4):311, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900243.

Anonymous:1978:ENc

[Ano78n] Anonymous. Erratum notice. Computer Physics Communications, 14(3–4):312, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900255.

Anonymous:1978:ENd

[Ano78o] Anonymous. Erratum notice. Computer Physics Communications, 15(5):443, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900735.

Anonymous:1978:ICF

[Ano78p] Anonymous. International conference on finite elements in nonlinear mechanics. Computer Physics Communications, 13(5–6): 431–433, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900425.

Anonymous:1978:P

[Ano78q] Anonymous. Preface. Computer Physics Communications, 15(5): 303, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900589.

Anonymous:1978:PIVa

[Ano78r] Anonymous. Program index to volume 12. Computer Physics Communications, 13(5–6):443–445, January/February 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900450.

Anonymous:1978:PIVb

[Ano78s] Anonymous. Program index to volume 14. Computer Physics Communications, 14(5–6):454–457, July/August 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900103.

Anonymous:1978:PIVc

[Ano78t] Anonymous. Program index to volumes 11-15. Computer Physics Communications, 15(5):451–465, November 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900760.

Anonymous:1979:Aa

[Ano79a] Anonymous. Announcement. Computer Physics Communications, 16(3):401, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557990047X.

Anonymous:1979:Ab

[Ano79b] Anonymous. Announcement. Computer Physics Communications, 18(3):443, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900146.

Anonymous:1979:AIVa

[Ano79c] Anonymous. Author index to volume 16. Computer Physics Communications, 16(3):405–407, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900493.

Anonymous:1979:AIVb

[Ano79d] Anonymous. Author index to volume 17. Computer Physics Communications, 17(4):433–436, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579901097.

Anonymous:1979:AIVc

[Ano79e] Anonymous. Author index to volume 18. Computer Physics Communications, 18(3):447–450, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557990016X.

Anonymous:1979:BRB

[Ano79f] Anonymous. Book review: Computational Atomic and Molecular Physics: Proceedings of the Third EPS Computational Physics Conference, Nottingham, England, 12–15 September 1978. Computer Physics Communications, 17(1–2):ix, April/May 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900626.

Anonymous:1979:CVa

[Ano79g] Anonymous. Contents to volume 16. Computer Physics Communications, 16(3):403–404, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900481.

Anonymous:1979:CVb

[Ano79h] Anonymous. Contents to volume 17. Computer Physics Communications, 17(4):431–432, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901085.

Anonymous:1979:CVc

[Ano79i] Anonymous. Contents to volume 18. Computer Physics Communications, 18(3):445–446, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900158.

Anonymous:1979:EBa

[Ano79j] Anonymous. Editorial Board. Computer Physics Communications, 17(1–2):??, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900614.

Anonymous:1979:EBb

[Ano79k] Anonymous. Editorial Board. Computer Physics Communications, 18(1):v–vii, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900183.

Anonymous:1979:ENa

[Ano79l] Anonymous. Erratum notice. Computer Physics Communications, 17(4):423–424, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557990105X.

Anonymous:1979:ENc

[Ano79m] Anonymous. Erratum notice. Computer Physics Communications, 17(4):427–430, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901073.

Anonymous:1979:ENd

[Ano79n] Anonymous. Erratum notice. Computer Physics Communications, 18(3):441, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900134.

Anonymous:1979:ENb

[Ano79o] Anonymous. Erratum notices. Computer Physics Communications, 17(4):425, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901061.

Anonymous:1979:PIVa

[Ano79p] Anonymous. Program index to volume 16. Computer Physics Communications, 16(3):408–410, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557990050X.

Anonymous:1979:PIVb

[Ano79q] Anonymous. Program index to volume 17. Computer Physics Communications, 17(4):437–439, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579901103.

Anonymous:1979:PIVc

[Ano79r] Anonymous. Program index to volume 18. Computer Physics Communications, 18(3):451–453, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900171.

Assimakopoulos:1975:KTB

[Ass75] P. A. Assimakopoulos. Kinematics of three-body reactions. Computer Physics Communications, 10(6):385–400, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900429.

Abdallah:1975:CEI

[AT75] Joseph Abdallah, Jr. and Donald G. Truhlar. Continuum exchange integrals for algebraic variational calculations of electronatom scattering using Slater-type orbitals as basis functions. Computer Physics Communications, 9(5):327–336, May 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900843.

Appert:1976:FEA

[ATV76] K. Appert, T. M. Tran, and J. Vaclavik. Finite element approximation for the wave-particle interaction in weakly turbulent plasmas. Computer Physics Communications, 12(2):135–144, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557690062X.

Auerbach:1978:TGO

[Aue78] E. H. Auerbach. A-Three: a general optical model code especially suited to heavy-ion calculations. Computer Physics Communications, 15(3–4):165–192, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900917.

Aurela:1977:CSN

[Aur77] A. M. Aurela. Counting a small number of radioactive atoms. Computer Physics Communications, 13(4):281–287, November/ December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900066.

Akylas:1978:MAC

[AV78] V. R. Akylas and P. Vogel. Muonic atom cascade program. Computer Physics Communications, 15(3–4):291–302, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900991.

Atkinson:1971:CAA

[AWG71] G. D. Atkinson, Jr., J. B. Whitworth, and S. J. Gage. Computerassisted analysis of gamma-ray spectra. Computer Physics Communications, 2(1):40–46, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900130.

Armenise:1979:PIS

[AZS^{+}79] N. Armenise, G. Zito, A. Silvestri, E. Lefons, M. T. Pazienza, and F. Tangorra. POL: An interactive system to analyze large data sets. Computer Physics Communications, 16(2):147–157, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900833.

Barnett:1976:RMR

[Bar76] A. R. Barnett. RCWFF — modification of the real Coulomb wavefunction program RCWFN. Computer Physics Communications, 11(1):141–142, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557690045X.

Barnard:1977:FFF

[Bar77] A. J. Barnard. Furi — a FORTRAN function writer. Computer Physics Communications, 13(4):271–280, November/December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465577900054.

Bates:1974:FCH

[Bat74] G. N. Bates. A fixed core Hartree–Fock program for calculating bound and continuum orbitals. Computer Physics Communications, 8(3):220–235, October 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900988.

Basso:1970:AGP

[BB70] P. Basso and C. Bourrely. An algorithm generating permutations with identical objects. Computer Physics Communications, 1(6):415–419, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900147.

Baas:1974:AID

[BB74] R. Ch. Baas and C. I. M. Beenakker. Analysis of the intensity distribution in the rotational structure of the electronic spectra of diatomic molecules by computer simulation. Computer Physics Communications, 8(3):236–245, October 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557490099X.

Bose:1977:CCA

[BB77] A. K. Bose and A. Bose. Combinatorial computation of atomic terms for equivalent electrons. Computer Physics Communications, 13(4):241–246, November/December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900029.

Bhattacharya:1979:NPS

[BB79] R. N. Bhattacharya and Iva Basu. Note on the principle of stationary phase. Computer Physics Communications, 16(2): 167–173, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900857.

Berrington:1974:GPC

[BBC^{+}74] K. A. Berrington, P. G. Burke, J. J. Chang, A. T. Chivers, W. D. Robb, and K. T. Taylor. A general program to calculate atomic continuum processes using the R-matrix method. Computer Physics Communications, 8(3):149–198, October 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900964.

Banks:1977:CCP

[BBH^{+}77] D. Banks, K. S. Barnes, P. E. Hughes, I. C. Percival, D. Richards, N. A. Valentine, and J. Mc. B. Wilson. Classical collisions of protons with hydrogen atoms. Computer Physics Communications, 13(4):251–269, November/December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900042.

Berrington:1978:NVG

[BBL^{+}78] K. A. Berrington, P. G. Burke, M. Le Dourneuf, W. D. Robb, K. T. Taylor, and Vo Ky Lan. A new version of the general program to calculate atomic continuum processes using the R-matrix method. Computer Physics Communications, 14(5–6):367–412, July/August 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900048.

Buckley:1979:MCE

[BBL79] B. D. Buckley, P. G. Burke, and Vo Ky Lan. R-matrix calculations for electron-molecule scattering. Computer Physics Communications, 17(1–2):175–179, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900808.

Braithwaite:1972:RRM

[BC72] W. J. Braithwaite and J. G. Cramer. The reduced rotation matrix. Computer Physics Communications, 3(4):318–321, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557290080X.

Berrington:1979:RDE

[BC79] Keith Berrington and Martin Crees. Recent developments in electron collision calculations. Computer Physics Communications, 17(1–2):181–205, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557990081X.

Burton:1979:CHC

[BCE79] B. Burton, T. A. Claxton, and Y. Ellinger. Calculation of hyperfine coupling constants using second order double perturbation theory and the configuration interaction method. Computer Physics Communications, 17(1–2):27–30, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900663.

Bignami:1972:MAA

[BCV^{+}72] G. F. Bignami, G. Cioni, A. Della Ventura, P. Mussio, M. J. L. Turner, and U. Volonte’. A method for the automatic analysis of gamma ray events in astronomical spark chambers. Computer Physics Communications, 4(3):299–314, December 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900914.

Bettels:1972:PGA

[BD72] J. Bettels and P. Dodd. A program for the generation of artificial bubble chamber events. Computer Physics Communications, 3(2): 136–154, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900616.

Bird:1978:IER

[BDD78] B. Bird, C. Daul, and P. Day. Inter-electron repulsion integrals for three-open-shell configurations in cubic symmetry. Computer Physics Communications, 14(3–4):273–285, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900218.

Bardin:1972:CFE

[BDG^{+}72] C. Bardin, Y. Dandeu, L. Gauthier, J. Guillermin, T. Lena, J.- M. Pernet, H. H. Wolter, and T. Tamura. Coulomb functions in entire (η, ρ) plane. Computer Physics Communications, 3(2): 73–87, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900574.

Beeler:1973:BRB

[Bee73] Joe R. Beeler, Jr. Book review: Computational Solid State Physics: edited by F. Herman, N. W. Dalton and T. R. Koehler, Plenum Press, 1972, £12.15. Computer Physics Communications, 6(1): 63–64, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900234.

Brokate:1972:RBC

[BEK^{+}72] K. Brokate, R. Erbe, E. Keppel, D. Kropp, E. Grimm, H. Schneider, and J. Wells. Recognition of bubble chamber events exemplified by the Heidelberg automatic system. Computer Physics Communications, 4(3):315–326, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900926.

Beleznay:1972:NSR

[Bel72] F. Beleznay. Numerical solution of the radical Schrödinger equation. Computer Physics Communications, 3(4):334–338, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900823.

Bolger:1979:WPH

[BEM79] Joseph E. Bolger, Hunter Ellinger, and C. Fred Moore. Worker, a program for histogram manipulation. Computer Physics Communications, 16(3):345–361, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900419.

Bermudez:1977:EFS

[Ber77] Victor M. Bermudez. Ellips — A FORTRAN simulation of a polarization-modulation ellipsometer. Computer Physics Communications, 13(3):207–224, September/October 1977. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465577900157.

Betak:1975:PSC

[Bet75] Emil Beták. Program for spectra and cross-section calculations within the pre-equilibrium model of nuclear reactions. Computer Physics Communications, 9(2):92–101, February 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900260.

Browne:1971:EMR

[BF71] E. Browne and R. F. Femenia. Electromagnetic M1 reduced transition probabilities for pure and mixed Nilsson states in odd a nuclei. Computer Physics Communications, 2(6):331–340, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900257.

Bottcher:1977:EMI

[BFKD77] C. Bottcher, A. L. Ford, and K. Kirby-Docken. Evaluation of molecular integrals involving continuum orbitals. Computer Physics Communications, 13(1):11–15, May 1977. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465577900224.

Barnett:1974:CWF

[BFSG74] A. R. Barnett, D. H. Feng, J. W. Steed, and L. J. B. Goldfarb. Coulomb wave functions for all real η and ρ. Computer Physics Communications, 8(5):377–395, December 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900137.

Beiner:1972:DNI

[BG72] M. Beiner and P. Gara. Direct numerical integration of coupled one-dimensional Schrödinger equations. Computer Physics Communications, 4(1):1–9, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900227.

Batana:1979:AFI

[BGM79] Alicia Batana, Ernesto R. Gonzalez, and Maria C. Monard. Analysis of Faradaic impedance experimental measurements. Computer Physics Communications, 18(1):27–34, September 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900225.

Beatham:1979:MPC

[BGMP79] N. Beatham, I. P. Grant, B. J. McKenzie, and N. C. Pyper. MCBP — A program to calculate angular coefficients of the Breit interaction between electrons in the low energy limit. Computer Physics Communications, 18(2):245–260, October/ November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901164.

Berger:1976:FEA

[BGT76] D. Berger, R. Gruber, and F. Troyon. A finite element approach to the computation of the MHD spectrum of straight noncircular plasma equilibria. Computer Physics Communications, 11(3):313–323, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900187.

Baltes:1972:PWP

[BH72] H. P. Baltes and E. R. Hilf. Progress in Weyl’s problem achieved by computational methods. Computer Physics Communications, 4(2):208–213, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900100.

Brown:1979:ASA

[BH79] W. S. Brown and A. C. Hearn. Applications of symbolic algebraic computation. Computer Physics Communications, 17(1–2):207– 215, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900821.

Biskamp:1975:EPS

[Bis75] D. Biskamp. European Physical Society. Computer Physics Communications, 10(5):259, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900946.

Biskamp:1976:P

[Bis76] D. Biskamp. Preface. Computer Physics Communications, 12 (1):vii, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900023.

Baas:1976:CAE

[BJ76] R. Ch. Baas and R. H. J. Jansen. Computer analysis of experimental results on differential scattering of electrons by gases. Computer Physics Communications, 12(3):267–276, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900862.

Beck:1972:CGW

[BK72] Robert E. Beck and Bernard Kolman. Computer generated Weyl groups. Computer Physics Communications, 3(2):155–158, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900628.

Beck:1974:ROM

[BK74] Robert E. Beck and Bernard Kolman. Racah’s outer multiplicity formula. Computer Physics Communications, 8(2):95–100, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900599.

Balint-Kurti:1979:RCR

[BKE79] G. G. Balint-Kurti and L. Eno. The rapid calculation of rotationally and vibrationally inelastic molecular collision cross sections. Computer Physics Communications, 17(1–2):85–88, April/ May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900717.

Bers:1976:SCN

[BKK76] A. Bers, J. L. Kulp, and C. F. F. Karney. Symbolic computation of nonlinear wave interactions on MACSYMA. Computer Physics Communications, 12(1):81–98, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900126.

Blowers:1972:CTD

[Blo72] R. M. Blowers. The computation of two-dimensional viscous flow through a T-junction. Computer Physics Communications, 4(1): 64–72, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900331.

Brundrit:1976:CFE

[BM76] G. B. Brundrit and M. J. Miketinac. Calculation of the form of an equilibrium poloidal magnetic field contained in a polytropic star. Computer Physics Communications, 11(3):385–395, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900254.

Boujot:1972:NSE

[BMS72] J. P. Boujot and C. Mercier-Soubbaramayer. Numerical simulation of the evolution of a plasma in a confinement device. Computer Physics Communications, 4(1):89–94, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900367.

Bock:1975:PPH

[Böc75] R. K. Böck. Program portability in high-energy physics. Computer Physics Communications, 9(4):221–229, April 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900090.

Bogolubsky:1977:SEI

[Bog77] I. L. Bogolubsky. Some examples of inelastic soliton interaction. Computer Physics Communications, 13(3):149–155, September/October 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900091.

Bok:1978:SAC

[Bok78] J. Bok. A subroutine for approximation by cubic splines in the least squares sense. Computer Physics Communications, 16(1):113–118, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578901145.

Boris:1976:NSC

[Bor76] J. P. Boris. Numerical solution of continuity equations. Computer Physics Communications, 12(1):67–79, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900114.

Bose:1976:AGP

[Bos76a] A. K. Bose. An algorithm for generating the positive roots of simple Lie algebras. Computer Physics Communications, 11(2):159–161, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900473.

Bose:1976:CDI

[Bos76b] A. K. Bose. Computing dimensions of irreducible representations of simple Lie algebras. Computer Physics Communications, 11(1):1–4, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900357.

Bouvier:1972:NEC

[Bou72] P. Bouvier. Numerical experiments concerning isolated and non-isolated stellar systems. Computer Physics Communications, 4(3):345–346, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900963.

Brailsford:1973:OSI

[BP73] David F. Brailsford and John A. Prentice. The OPIT system II. Job control and scheduling for large applications programs. Computer Physics Communications, 5(2):136–146, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900295.

Bradly:1975:CCC

[BP75] D. L. Bradly and R. Perrin. Colour coordinate calculations. Computer Physics Communications, 9(5):305–311, May 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046557590082X.

Bleris:1977:PSS

[BP77] G. L. Bleris and Ch. Polatoglou. A program for the study of short range order of binary alloys. Computer Physics Communications, 13(1):49–56, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900261.

Banks:1971:CRM

[BPW71] D. Banks, I. C. Percival, and J. McB. Wilson. Classical relative motion of 2 particles. Computer Physics Communications, 2(2):114–123, February/March 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557190021X.

Banks:1972:CMP

[BPW72a] D. Banks, I. C. Percival, and J. McB. Wilson. Classical motion of 2 particles (EVA2 edition 01). Computer Physics Communications, 3(3):221–239, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900690.

Banks:1972:CRM

[BPW72b] D. Banks, I. C. Percival, and J. McB. Wilson. Classical relative motion of 2 particles (EVAR edition 02). Computer Physics Communications, 3(3):197–220, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900689.

Banks:1972:SFS

[BPW72c] D. Banks, I. C. Percival, and J. McB. Wilson. Stirling FORDOC 01. A set of documentation conventions for FORTRAN packages and routines. Computer Physics Communications, 3(3):180–196, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900677.

Bock:1973:SCL

[BPZ73] R. K. Böck, E. Pagiola, and J. Zoll. Software concepts for large application programs. Computer Physics Communications, 5(6): 400–403, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900763.

Braithwaite:1972:RKTb

[Bra72a] W. J. Braithwaite. Relativistic kinematics for three-body final states. Computer Physics Communications, 4(2):233–238, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900148.

Braithwaite:1972:RKTa

[Bra72b] W. J. Braithwaite. Relativistic kinematics for two-body final states. Computer Physics Communications, 4(2):227–232, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900136.

Bransden:1972:BRB

[Bra72c] B. H. Bransden. Book review: Methods in Computational Physics, vol. X. Atomic and molecular scattering: eds. B. Alder, S. Fernbach and M. Rotenberg, Academic Press £10.25. Computer Physics Communications, 4(3):384, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901026.

Braithwaite:1973:ALP

[Bra73] W. J. Braithwaite. Associated Legendre polynomials, ordinary and modified spherical harmonics. Computer Physics Communications, 5(5):390–394, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900659.

Braunschweig:1978:IRM

[Bra78a] D. Braunschweig. II. Reduced SU(3) matrix elements. Computer Physics Communications, 15(3–4):259–273, October 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900966.

Braunschweig:1978:RC

[Bra78b] D. Braunschweig. Reduced SU(3) CFP’s. Computer Physics Communications, 14(1–2):109–120, March/April 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900541.

Banuelos:1978:PCA

[BRT78] Alicia Bañuelos and Felix Rodriguez-Trelles. A program for calculating the angular distribution of nonrelativistic Bremsstrahlung intensity. Computer Physics Communications, 15(1–2):125–129, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900863.

Bernussou:1974:CAP

[BS74] J. Bernussou and N. F. Stewart. Computation of area preserving mappings. Computer Physics Communications, 7(2):63–66, February 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557490040X.

Brandt:1975:CDT

[BS75a] Siegmund Brandt and Hermann Schneider. Computer-drawn trajectories of particles in mutual and external fields — an application of computers in university physics teaching. Computer Physics Communications, 9(4):205–220, April 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900089.

Brocklehurst:1975:CLC

[BS75b] M. Brocklehurst and M. Salem. Computation of line and continuum radiation from thermal radioastronomical sources. Computer Physics Communications, 9(4):258–267, April 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900120.

Brocklehurst:1977:RRL

[BS77] M. Brocklehurst and M. Salem. Radio recombination lines from H^{+} regions and cold interstellar clouds: Computation of the b_{n} factors. Computer Physics Communications, 13(1):39–48, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557790025X.

Bastien:1971:TMP

[BSP71] P. L. Bastien, J. N. Snyder, and V. Pless. A track matching program for bubble chamber photographs. Computer Physics Communications, 2(7):394–419, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900336.

Brandt:1973:PCD

[BTS73] Maynard A. Brandt, Donald G. Truhlar, and Richard L. Smith. Program for calculating differential and integral cross sections for quantum mechanical scattering problems from reactance or transition matrices. Computer Physics Communications, 5(6):456– 477, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900829.

Brandt:1974:PAC

[BTS74] Maynard A. Brandt, Donald G. Truhlar, and Richard L. Smith. Program ACRL to calculate differential and integral cross sections adapted to run on IBM computers. Computer Physics Communications, 7(3):172–173, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900071.

Buneman:1976:AEE

[Bun76] O. Buneman. The advance from 2D electrostatic to 3D electromagnetic particle simulation. Computer Physics Communications, 12(1):21–31, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900072.

Burke:1970:E

[Bur70a] P. G. Burke. Editorial. Computer Physics Communications, 1(6): 379, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900111.

Burke:1970:EPA

[Bur70b] P. G. Burke. Editorial: Program annotation. Computer Physics Communications, 1(3):141–150, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900019.

Burke:1970:PCG

[Bur70c] P. G. Burke. A program to calculate a general recoupling coefficient. Computer Physics Communications, 1(4):241–250, April 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900408.

Burke:1972:BRB

[Bur72a] P. G. Burke. Book review: The Padé approximant in theoretical physics: eds. G. A. Baker, Jr. and J. L. Gammel, Academic Press, £8.15. Computer Physics Communications, 3(3):272, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900732.

Burke:1972:E

[Bur72b] P. G. Burke. Editorial. Computer Physics Communications, 4 (1):v, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900215.

Burke:1972:ICA

[Bur72c] P. G. Burke. The impact of computers on atomic physics. Computer Physics Communications, 3(S1):1–13, September 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572901105.

Burke:1973:MMA

[Bur73] P. G. Burke. The R-matrix method in atomic physics. Computer Physics Communications, 6(6):288–302, December 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900386.

Burke:1974:E

[Bur74] P. G. Burke. Editorial. Computer Physics Communications, 8(1): iii–iv, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900782.

Burke:1979:P

[Bur79] P. G. Burke. Preface. Computer Physics Communications, 17(1–2): xi, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900638.

Buttle:1978:DPH

[But78] P. J. A. Buttle. DWBA program for heavy ion transfer reactions. Computer Physics Communications, 14(1–2):133–143, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900565.

Canut-Amoros:1970:SCP

[CA70] Marisa Canut-Amoros. STLPLT — Calcomp plot of crystallographic projections of Laue photographs. Computer Physics Communications, 1(5):293–305, September 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900317.

Calmet:1972:RAC

[Cal72] Jacques Calmet. A REDUCE approach to the calculation of Feynman diagrams. Computer Physics Communications, 4(2):199–204, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900082.

Callaway:1973:POM

[Cal73] J. Callaway. Polarized orbital methods. Computer Physics Communications, 6(6):265–274, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900362.

Campbell:1970:LAP

[Cam70] J. A. Campbell. LISP and its applications to physical problems. Computer Physics Communications, 1(4):251–264, April 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557090041X.

Cameron:1972:BRB

[Cam72] I. G. Cameron. Book review: Lecture Notes in Physics 8, Proceedings of the Second International Conference on Numerical Methods in Fluid Dynamics: ed. M. Holt, Springer-Verlag $7.70. Computer Physics Communications, 3(3):273–274, April/ May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900756.

Campbell:1975:DSS

[Cam75a] J. B. Campbell. Determination of SSOR-SI iteration parameters. Computer Physics Communications, 10(3):194–202, September 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900065.

Campbell:1975:PPD

[Cam75b] J. B. Campbell. A program package for the Dirichlet problem with axially symmetric boundary conditions. Computer Physics Communications, 9(5):283–296, May 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900806.

Campbell:1979:BFR

[Cam79] J. B. Campbell. Bessel functions J_{ν}(x) and Y_{ν}(x) of real order and real argument. Computer Physics Communications, 18(1):133–142, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900304.

Christiansen:1974:MOD

[CAR74] J. P. Christiansen, D. E. T. F. Ashby, and K. V. Roberts. MEDUSA a one-dimensional laser fusion code. Computer Physics Communications, 7(5):271–287, May 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900277.

Casimir:1972:ICI

[Cas72] H. B. G. Casimir. Impact of computers on industrial research and development. Computer Physics Communications, 3(S1):174– 178, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572901269.

Cheung:1977:GCR

[CB77] L. M. Cheung and D. M. Bishop. The group-coordinate relaxation method for solving the generalized eigenvalue problem for large real-symmetric matrices. Computer Physics Communications, 13 (4):247–250, November/December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900030.

Cohen:1978:PCS

[CB78] I. Cohen and F. Bark. Perturbation calculations for the spinup problem using Reduce. Computer Physics Communications, 14(5–6):319–326, July/August 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900024.

Ciulli:1979:GAE

[CC79] M. Ciulli and S. Ciulli. A guide to analytic extrapolations: Part I: a program for optimal extrapolation to interior points. Computer Physics Communications, 18(2):215–244, October/ November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901152.

Chernysheva:1976:SCF

[CCR76] L. V. Chernysheva, N. A. Cherepkov, and V. Radojević. Self-consistent field Hartree–Fock program for atoms. Computer Physics Communications, 11(1):57–73, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900400.

Chernysheva:1979:FCH

[CCR79] L. V. Chernysheva, N. A. Cherepkov, and V. Radojević. Frozen core Hartree–Fock program for atomic discrete and continuous states. Computer Physics Communications, 18(1):87–100, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900262.

Chamayou:1975:ASD

[CE75] J. M. F. Chamayou and M. E. A. El Tom. On the approximate solution of the delay integral equation of the statistical theory of radiation damage. Computer Physics Communications, 9(3):131– 140, March 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575901022.

Christiansen:1971:DTD

[CH71a] J. P. Christiansen and R. W. Hockney. Delsqphi, a two-dimensional Poisson-solver program. Computer Physics Communications, 2(3): 139–156, April 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900464.

Christiansen:1971:FFF

[CH71b] J. P. Christiansen and R. W. Hockney. Four67, a Fast Fourier Transform package. Computer Physics Communications, 2(3):127– 138, April 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900452.

Cuperman:1974:NCM

[CH74] S. Cuperman and A. Harten. A numerical code for multiple “water bag” gravitional systems. Computer Physics Communications, 8(4):307–319, November 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900204.

Crennell:1975:KFI

[CH75] D. J. Crennell and D. E. Hall. Kinematic fitting: An improved method for convergence. Computer Physics Communications, 9(2): 79–84, February 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900247.

Charney:1972:ICM

[Cha72] Jule Charney. Impact of computers on meteorology. Computer Physics Communications, 3(S1):117–126, September 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572901191.

Chandra:1973:GPS

[Cha73] N. Chandra. A general program to study the scattering of particles by solving coupled inhomogeneous second-order differential equations. Computer Physics Communications, 5(6):417– 429, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900799.

Chang:1974:PER

[Cha74] J. J. Chang. A program to evaluate the reduced matrix elements of one-particle tensor operators for the configurations in JJ - coupling. Computer Physics Communications, 7(4):225– 234, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900927.

Chamayou:1975:BSF

[Cha75] J. M. F. Chamayou. Bicubic spline function approximation of the solution of the fast-neutron transport equation. Computer Physics Communications, 10(5):282–291, November 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900983.

Chamayou:1976:AKR

[Cha76] J. M. F. Chamayou. Algorithms for the Kac and Renyi tests. Computer Physics Communications, 12(2):173–178, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900667.

Chen:1973:FTC

[Che73] Joseph C. Y. Chen. Faddeev-type coupled equations. Computer Physics Communications, 6(6):336–346, December 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900428.

Chisholm:1972:GKA

[Chi72] J. S. R. Chisholm. Generalisation of the Kahane algorithm for scalar products of λ-matrices. Computer Physics Communications, 4(2):205–207, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900094.

Chivers:1973:NVPa

[Chi73a] Alfred T. Chivers. A new version of the program to compute the asymptotic solution of coupled equations for electron scattering. Computer Physics Communications, 5(6):416, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900787.

Chivers:1973:NVPb

[Chi73b] Alfred T. Chivers. A new version of the program to compute the fractional parentage coefficients for equivalent d shell electrons. Computer Physics Communications, 6(2):88, August 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900544.

Collins:1973:EPW

[CHMM73] P. A. Collins, B. J. Hartley, R. W. Moore, and K. J. M. Moriarty. An efficient partial-wave analyser for the absorption model. Computer Physics Communications, 5(5):349–364, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900623.

Choudry:1970:IRT

[Cho70] A. Choudry. Isometric representation of two-dimensional matrices. Computer Physics Communications, 1(4):277–280, April 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900433.

Christensen:1974:CSP

[Chr74a] E. R. Christensen. Computer simulation of photons in spheric media for density gauges. Computer Physics Communications, 7(4): 192–199, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900885.

Christensen:1974:MCS

[Chr74b] E. R. Christensen. Monte Carlo simulation of photons in twolayered media for density gauges. Computer Physics Communications, 7(4):185–191, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900873.

Christiansen:1975:BRB

[Chr75a] J. P. Christiansen. Book review: Proceedings of the Fourth International Conference on Numerical Methods in Fluid Dynamics: edited by R. D. Richtmeyer, Springer-Verlag, 1975. $15.20. Computer Physics Communications, 10(4):258, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900727.

Cowler:1975:CDR

[CHR75b] M. S. Cowler, N. E. Hoskin, and A. G. Rowlinson. Calculation of the dynamic response of reactor containment systems to full core explosions. Computer Physics Communications, 9(2):69– 78, February 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900235.

Cameron:1978:CCS

[CHW^{+}78] I. G. Cameron, B. C. Hankin, A. G. P. Warham, A. Benuzzi, and A. Yerkess. The computer code SEURBNUK-2 for fast reactor containment safety studies. Computer Physics Communications, 13(5–6):297–310, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900279.

Chirikov:1973:NES

[CIT73] B. V. Chirikov, F. M. Izrailev, and V. A. Tayursky. Numerical experiments on the statistical behaviour of dynamical systems with a few degrees of freedom. Computer Physics Communications, 5(1): 11–16, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900039.

Clarkson:1971:ELS

[CJ71] R. G. Clarkson and Nelson Jarmie. Energy-loss straggling of heavy charged particles. Computer Physics Communications, 2(7):433– 442, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900361.

Cotterill:1973:MMT

[CKM^{+}73] R. M. J. Cotterill, W. Damgaard Kristensen, J. W. Martin, L. Børsting Pedersen, and K. J. Jensen. Melting of microcrystals in two dimensions: a molecular dynamics simulation. Computer Physics Communications, 5(1):28–31, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900064.

Connor:1979:QSS

[CL79] J. N. L. Connor and A. Laganá. Quasiclassical smooth sampling study of threshold behaviour for the collinear reaction X + F_{2} → XF + F (X = Mu, H, D, T). Computer Physics Communications, 17(1–2):145–148, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900766.

Clark:1978:ICC

[Cla78] Robert E. H. Clark. Intermediate coupling collision strengths from LS coupled R-matrix elements. Computer Physics Communications, 16(1):119–127, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578901157.

Conneely:1970:CPC

[CLS^{+}70] M. J. Conneely, L. Lipsky, K. Smith, P. G. Burke, and R. J. W. Henry. A computer program for the calculation of electron scattering and photoionization cross sections of atomic systems with configuration (np)^{q}. Computer Physics Communications, 1(5):306– 324, September 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900329.

Christie:1979:BSI

[CM79] M. A. Christie and K. J. M. Moriarty. A bicubic spline interpolation of unequally spaced data. Computer Physics Communications, 17(4):357–364, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900985.

Calinon:1977:SPE

[CMS77] R. Calinon, D. Merlini, and R. R. Sari. On the h-stability property and energy of Wigner lattices in the one component classical plasma: a numerical analysis. Computer Physics Communications, 13(1):1–8, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900200.

Clayton:1974:BAB

[CMW74] J. Clayton, R. Matthews, and N. West. BATMAN — an automatic batch processing system. Computer Physics Communications, 8(3):143–148, October 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900952.

Casilio:1973:AFV

[CN73] John M. Casilio and M. E. Noz. Analytic formulation of SU(3) vector coupling coefficients for n particles. Computer Physics Communications, 5(5):365–378, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900635.

Cohen:1971:DSS

[Coh71] Stanley Cohen. The Delphi–speakeasy system. I. Overall description. Computer Physics Communications, 2(1):1–10, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900087.

Collocott:1977:NSK

[Col77] S. J. Collocott. Numerical solution of Kramers–Kronig transforms by a Fourier method. Computer Physics Communications, 13(3): 203–206, September/October 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900145.

Comfort:1978:FES

[Com78] J. R. Comfort. Fitting elastic-scattering cross sections with coupled channels. Computer Physics Communications, 16(1):35–42, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578901066.

Connor:1979:RMC

[Con79] J. N. L. Connor. Reactive molecular collision calculations. Computer Physics Communications, 17(1–2):117–143, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900754.

Copley:1974:MCC

[Cop74] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments. Computer Physics Communications, 7(6):289–317, June 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900289.

Copley:1975:MCCa

[Cop75a] J. R. D. Copley. Monte Carlo calculation of multiple scattering effects in thermal neutron scattering experiments: Modification to SLAB geometry. Computer Physics Communications, 9(1):59– 63, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900570.

Copley:1975:MCCb

[Cop75b] J. R. D. Copley. Monte Carlo calculations of multiple scattering effects in thermal neutron scattering experiments: Modification to horizontal cylinder geometry. Computer Physics Communications, 9(1):64–68, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900582.

Christiansen:1974:OSC

[CR74] J. P. Christiansen and K. V. Roberts. OLYMPUS a standard control and utility package for initial-value FORTRAN programs. Computer Physics Communications, 7(5):245–270, May 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900265.

Cramer:1972:GPO

[Cra72] John G. Cramer. Graphical plotting output on a line printer using high-density plotting symbols. Computer Physics Communications, 3(4):314–317, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900793.

Christiansen:1978:AOD

[CRL78] J. P. Christiansen, K. V. Roberts, and J. W. Long. Athene 1: a one-dimensional equilibrium–diffusion code. Computer Physics Communications, 14(5–6):423–445, July/August 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900061.

Crowther:1972:ABS

[Cro72] R. A. Crowther. Analysis of biological structure at the atomic and molecular level. Computer Physics Communications, 3(S1):137– 138, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557290121X.

Chlouber:1978:ACF

[CS78] Clyde Chlouber and Mark A. Samuel. Accurate computation of fourth-order vacuum polarization. Computer Physics Communications, 15(3–4):153–159, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900899.

Choquard:1974:NDC

[CSC74] Ph. Choquard, R. R. Sari, and R. Calinon. Numerical determination of the critical point in the 3-D Ising model in Wilson’s formalism. Computer Physics Communications, 7(2):67– 71, February 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900411.

Caprini:1979:GAE

[CSP^{+}79] I. Caprini, M. Săraru, C. Pomponiu, M. Ciulli, S. Ciulli, and I. S. Stefănescu. A guide to analytic extrapolations: Part II: a program to be used in finding analytic correlations of data, and for detecting zeros and poles of the scattering amplitude. Computer Physics Communications, 18(3):305–326, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900018.

Crees:1978:IPS

[CSW78] M. A. Crees, M. J. Seaton, and P. M. H. Wilson. Impact, a program for the solution of the coupled integro-differential equations of electron-atom collision theory. Computer Physics Communications, 15(1–2):23–83, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900826.

Chrisman:1971:CAM

[CT71] B. L. Chrisman and T. A. Tumolillo. Computer analysis of Mössbauer spectra. Computer Physics Communications, 2(6): 322–330, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900245.

Collocott:1979:ANS

[CT79] S. J. Collocott and G. J. Troup. Adaptation: Numerical solution of the Kramers–Kronig transforms by trapezoidal summation as compared to a Fourier method. Computer Physics Communications, 17(4):393–395, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901024.

Cugnon:1973:CSB

[Cug73] J. Cugnon. Computation of S-state binding energy and wave functions in a Saxon–Woods potential. Computer Physics Communications, 6(1):17–23, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900180.

Chang:1979:PGC

[CW79a] B. D. Chang and S. S. M. Wong. A program to generate closed basic diagrams for product operators. Computer Physics Communications, 18(1):35–61, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900237.

Christiansen:1979:CTD

[CW79b] J. P. Christiansen and N. K. Winsor. Castor 2: a twodimensional laser target code. Computer Physics Communications, 17(4):397–412, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901036.

Dawson:1972:CCS

[Daw72] John M. Dawson. Contribution of computer simulation to plasma theory. Computer Physics Communications, 3(S1):79–85, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572901166.

Day:1977:PTE

[Day77] Colin Day. A portable text editor. Computer Physics Communications, 13(2):101–105, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900352.

Dahl:1971:CSE

[DBB71] R. E. Dahl, Jr., J. R. Beeler, Jr., and R. D. Bourquin. Computer simulation of extended defects in metals. Computer Physics Communications, 2(6):301–321, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900233.

DeHosson:1975:NCA

[De 75a] J. Th. M. De Hosson. Normal coordinate analysis of crystals. Computer Physics Communications, 10(2):104–116, August 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900776.

DeHosson:1975:SB

[De 75b] J. Th. M. De Hosson. Symmetry and bandstructure. Computer Physics Communications, 10(1):67–69, July 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900910.

DeMey:1977:NSD

[De 77] G. De Mey. Numerical solution of a drift-diffusion problem with special boundary conditions by integral equations. Computer Physics Communications, 13(2):81–88, July 1977. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465577900339.

Davies:1972:CGG

[DE72] A. J. Davies and C. J. Evans. The computation of the growth of a gaseous discharge in space-charge distored fields. Computer Physics Communications, 3(4):322–333, July 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900811.

Delves:1977:BRB

[Del77] I. M. Delves. Book review: Matrix computation for engineers and scientists: Alan Jennings, John Wiley & Sons, 1977 £10.50/ $22.00. Computer Physics Communications, 13(2):147, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465577900406.

Delic:1979:CEA

[Del79a] G. Delic. Chebyshev expansion of the associated Legendre polynomial P_{L}^{M}(x). Computer Physics Communications, 18(1):63–71, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900249.

Delic:1979:CSS

[Del79b] G. Delic. Chebyshev series for the spherical Bessel function j_{l}(r). Computer Physics Communications, 18(1):73–86, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900250.

Demmig:1972:ESO

[Dem72] F. Demmig. Explicit second order method of characteristics for quasilinear hyperbolic differential equations and applications. Computer Physics Communications, 4(2):186–192, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900069.

Demmig:1978:COD

[Dem78] F. Demmig. The computation of one-dimensional unsteady nonequilibrium flows with a method of characteristics utilizing exponential fitting. Computer Physics Communications, 14(1–2):7–12, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900486.

Desclaux:1970:HFS

[Des70] J. P. Desclaux. Hartree Fock Slater self consistent field calculations. Computer Physics Communications, 1(3):216–222, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900081.

Desclaux:1975:MRD

[Des75] J. P. Desclaux. A multiconfiguration relativistic DIRAC–FOCK program. Computer Physics Communications, 9(1):31–45, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900545.

Douady:1979:QCS

[DESL79] J. Douady, Y. Ellinger, R. Subra, and B. Levy. A quadratically convergent SCF procedure. Computer Physics Communications, 17(1–2):23–25, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900651.

Deutsch:1979:SHSa

[Deu79a] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering I: Intensity correction program. Computer Physics Communications, 17(4):337–343, July/August 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046557990095X.

Deutsch:1979:SHSb

[Deu79b] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering II: Computation of the correction function. Computer Physics Communications, 17(4):345–349, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900961.

Deutsch:1979:SHSc

[Deu79c] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering III: Intensity correction program adaptation to arbitrary slit transmission function. Computer Physics Communications, 18(1):143–147, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900316.

Deutsch:1979:SHSd

[Deu79d] Moshe Deutsch. Slit height smearing correction in small angle X-ray scattering IV: Computation of the correction function for an arbitrary slit transmission function. Computer Physics Communications, 18(1):149–154, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900328.

DeVries:1976:QBS

[DeV76] R. M. DeVries. Quasi-bound state wavefunctions. Computer Physics Communications, 11(2):249–256, March/May 1976. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465576900564.

Davies:1978:SGA

[DEW78] A. J. Davies, C. J. Evans, and P. M. Woodison. Simulation of the growth of axially symmetric discharges between plane parallel electrodes. Computer Physics Communications, 14(3–4):287– 297, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557890022X.

Docken:1976:DOI

[DF76] Kate K. Docken and A. Lewis Ford. Dipole and overlap integrals between Slater-type functions and continuum Coulomb functions. Computer Physics Communications, 11(1):49–55, January/ February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900394.

Darville:1975:ECZ

[DG75] J. Darville and A. Gérard. Exact computation of the Zeeman effect on nuclear quadrupole resonance profiles for powders (SPIN I = 32) determination of the asymmetry parameter. Computer Physics Communications, 9(3):173–181, March 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575901046.

Dragoun:1971:PCI

[DH71] O. Dragoun and G. Heuser. A program to calculate internal conversion coefficients for all atomic shells without screening. Computer Physics Communications, 2(7):427–432, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557190035X.

Dudder:1975:RTD

[DH75] H. D. Düdder and D. B. Henderson. RAMSES, a two-dimensional, pic type, laser pulse propagation code. Computer Physics Communications, 10(3):155–166, September 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557590003X.

Delbrouck-Habaru:1974:BSO

[DHD74] J. M. Delbrouck-Habaru and Daniel M. Dubois. Bound states of one nucleon in a Woods–Saxon well from a variational method. Computer Physics Communications, 8(5):396–403, December 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900149.

Der:1979:PCI

[DHN79] R. Der, D. Hinneburg, and M. Nagel. A program to calculate internal conversion coefficients including higher-order corrections for all atomic shells. Computer Physics Communications, 18(3):401– 410, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900092.

Dickson:1979:NRH

[Dic79] A. S. Dickson. Non-reactive heavy particle collision calculations. Computer Physics Communications, 17(1–2):51–80, April/ May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900699.

deJagher:1978:NPC

[dJ78] P. C. de Jagher. Nullijn, a program to calculate zero curves of a function of two variables of which one may be complex. Computer Physics Communications, 15(5):351–373, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900668.

Dobes:1979:CGH

[Dob79] J. Dobeš. Calculations of generalized harmonic oscillator brackets. Computer Physics Communications, 16(3):373–381, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900432.

Dufour:1975:GPM

[DS75] P. Dufour and J. Schlesinger. A general purpose Monte–Carlo program. Computer Physics Communications, 9(6):360–369, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900168.

Davies:1972:DDI

[DSK72] Alan R. Davies, Kenneth Smith, and K. L. Kwok. Dirac, dynamic information retrieval of atomic codes: 1. Physical design criteria. Computer Physics Communications, 3(4):277–295, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557290077X.

Davies:1973:DDI

[DSK73] Alan R. Davies, Kenneth Smith, and K. L. Kwok. DIRAC dynamic information retrieval of atomic codes: II. Implementation. Computer Physics Communications, 6(4):165–186, October 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900982.

Davies:1975:TCL

[DST75] A. R. Davies, Kenneth Smith, and R. M. Thomson. TLASER — A CO_{2} laser kinetics code. Computer Physics Communications, 10(2):117–132, August 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900788.

Davies:1979:CSE

[DT79] D. W. Davies and S. J. Till. A computational study of the effect of the interaction potential for inelastic atom-molecule rotational scattering. Computer Physics Communications, 17(1–2):81– 84, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900705.

Datta:1974:VNO

[DTE74] S. K. Datta, W. J. Thompson, and D. O. Elliott. A version of a nuclear optical model code for small computers designed to run on a PDP-15. Computer Physics Communications, 7(6): 343, June 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900319.

Dufton:1977:PCC

[Duf77] P. L. Dufton. A program to calculate coronal emission line strengths. Computer Physics Communications, 13(1):25–38, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465577900248.

Dodds:1972:VCC

[DW72] R. McD. Dodds and G. Wiechers. Vector coupling coefficients as products of prime factors. Computer Physics Communications, 4(2):268–274, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900197.

Ellison:1974:NMS

[EG74] D. Ellison and I. P. Grant. A numerical method for solving the radiative transfer equation in two-dimensional X–Y geometry. Computer Physics Communications, 8(4):257–269, November 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900186.

Eggert:1972:AMS

[EGMS72] K. Eggert, W. Gürich, W. Moze, and C. Stolze. Automatic measurement of streamer chamber pictures. Computer Physics Communications, 4(3):294–298, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900902.

Eastwood:1977:SEP

[EJ77] J. W. Eastwood and C. R. Jesshope. The solution of elliptic partial differential equations using number theoretic transforms with application to narrow or limited computer hardware. Computer Physics Communications, 13(4):233–239, November/ December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900017.

Eissner:1974:TCA

[EJN74] W. Eissner, M. Jones, and H. Nussbaumer. Techniques for the calculation of atomic structures and radiative data including relativistic corrections. Computer Physics Communications, 8(4):270– 306, November 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900198.

Eberhard:1972:OSF

[EK72] Philippe H. Eberhard and Werner O. Koellner. The optime system for fitting theoretical expressions. Computer Physics Communications, 3(4):296–313, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900781.

Eberhard:1973:UMO

[EK73] Philippe H. Eberhard and Werner O. Koellner. Users manual for the Optime system. Computer Physics Communications, 5(3):163– 215, March 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573901008.

Ellzey:1975:CCF

[Ell75] M. L. Ellzey, Jr. Complete crystal field calculations including spin-orbit interaction and paramagnetic effects. Computer Physics Communications, 10(3):133–143, September 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900016.

Eisenstein:1974:PCP

[EM74] R. A. Eisenstein and G. A. Miller. Pirk: a computer program to calculate the elastic scattering of pions from nuclei. Computer Physics Communications, 8(2):130–140, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900630.

Eisenstein:1976:DCP

[EM76] R. A. Eisenstein and G. A. Miller. DWPI: a computer program to calculate the inelastic scattering of pions from nuclei. Computer Physics Communications, 11(1):95–112, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900424.

Eisenstein:1976:PMS

[ET76] R. A. Eisenstein and Frank Tabakin. Pipit: a momentum space optical potential code for pions. Computer Physics Communications, 12(2):237–257, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900722.

Eastgate:1973:NOM

[ETH73] R. J. Eastgate, W. J. Thompson, and R. A. Hardekopf. A nuclear optical model code for small computers. Computer Physics Communications, 5(1):69–79, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900131.

Feng:1975:PCC

[FB75] Da Hsuan Feng and A. R. Barnett. PATIWEN — a code for Coulomb-nuclear interference calculations. Computer Physics Communications, 10(6):401–420, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900430.

Foster:1971:PAE

[FC71] B. P. Foster and D. C. Camp. Programs to aid in establishing gamma-ray decay schemes, DCSCH3 and DCSCH4. Computer Physics Communications, 2(5):288–297, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557190004X.

Fleurier:1974:IAI

[FC74] C. Fleurier and J. Chapelle. Inversion of Abel’s integral equation — application to plasma spectroscopy. Computer Physics Communications, 7(4):200–206, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900897.

Fang:1978:CSS

[FCW78] M. T. C. Fang, S. K. Chan, and R. D. Wright. The computation of steady state arcs in nozzle flow. Computer Physics Communications, 13(5–6):363–370, January/February 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900334.

Femenia:1971:CPC

[Fem71] F. R. Femenia. Codnum, a program to change the punching code and to number a deck of cards. Computer Physics Communications, 2(3):168–172, April 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900488.

Ferguson:1975:PCX

[Fer75] I. F. Ferguson. A program for the calculation of X-ray reflection intensities, part 2. Computer Physics Communications, 10 (1):42–55, July 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900892.

Ferguson:1976:PCP

[FFH76] I. F. Ferguson, R. S. Fox, and T. E. Hughes. A program for the calculation of the positions of X-ray reflections. Computer Physics Communications, 12(3):305–321, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900898.

Ford:1974:GEP

[FH74] Brian Ford and George Hall. The generalized eigenvalue problem in quantum chemistry. Computer Physics Communications, 8(5):337– 348, December 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900113.

French:1972:PDH

[FJK^{+}72] B. French, F. James, L. Kowarski, M. Veltman, and A. Zichichi. Panel discussion, high-energy physics. Computer Physics Communications, 3(S1):157–165, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901245.

Ferguson:1973:PCX

[FK73] I. F. Ferguson and J. E. Kirwan. A program for the calculation of X-ray reflection intensities. Computer Physics Communications, 5(5):328–348, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900611.

Feng:1976:EFR

[FKU^{+}76] D. H. Feng, B. T. Kim, T. Udagawa, T. Tamura, and K. S. Low. Exact finite-range microscopic calculations for heavy-ion induced two-nucleon transfer reactions. Computer Physics Communications, 12(3):293–303, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900886.

Fletcher:1972:MSO

[Fle72] R. Fletcher. Methods for the solution of optimization problems. Computer Physics Communications, 3(3):159–172, April/ May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900653.

Fijalkow:1979:RPP

[FM79] E. Fijalkow and G. Mourgues. Radiation potential of a point antenna immersed in drifting cold or hot (hydrodynamical) plasma. Computer Physics Communications, 18(2):297–304, October/November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901218.

Ford:1978:PET

[For78] B. Ford. Parameterisation of the environment for transportable numerical software. Computer Physics Communications, 15(1–2): 1–3, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900796.

Foster:1972:DHC

[Fos72] W. Foster. Diffusion and heat conduction: a brief survey of numerical methods. Computer Physics Communications, 4(2):173– 181, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900045.

FroeseFischer:1973:EES

[FP73] C. Froese Fischer and D. W. B. Prentice. Exact evaluation of Slater integrals using ALTRAN, a symbol manipulation language. Computer Physics Communications, 6(4):157–164, October 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900970.

FroeseFischer:1970:MCH

[Fro70] Charlotte Froese Fischer. A multi-configuration Hartree–Fock program. Computer Physics Communications, 1(3):151–166, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900020.

FroeseFischer:1971:DDC

[Fro71] Charlotte Froese Fischer. The deffered difference correction for the Numerov method. Computer Physics Communications, 2 (2):124–126, February/March 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900221.

FroeseFischer:1972:MCH

[Fro72] Charlotte Froese Fischer. A multi-configuration Hartree–Fock program with improved stability. Computer Physics Communications, 4(1):107–116, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900392.

FroeseFischer:1973:CRP

[Fro73] Charlotte Froese Fischer. A critical review of programs in atomic structure for bound states. Computer Physics Communications, 5(2):147–152, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900301.

FroeseFischer:1978:GMC

[Fro78] Charlotte Froese Fischer. A general multi-configuration Hartree– Fock program. Computer Physics Communications, 14(1– 2):145–153, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900577.

Fortin:1973:NSE

[FS73] X. Fortin and R. Schuttler. Numerical study of the electronic and thermodynamic properties of the metal-non-metal transition in dense media. Computer Physics Communications, 5(1):24– 27, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900052.

Fan:1975:IAI

[FS75a] L. S. Fan and W. Squire. Inversion of Abel’s integral equation by a direct method. Computer Physics Communications, 10(2): 98–103, August 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900764.

FroeseFischer:1975:NVA

[FS75b] C. Froese Fischer and K. M. S. Saxena. A new version of AAKF (Reduced Tensor Matrix Elements) adapted to spectroscopic notation. Computer Physics Communications, 9(6):370– 380, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557590017X.

FroeseFischer:1975:OSN

[FS75c] C. Froese Fischer and K. M. S. Saxena. Oscillator strengths from numerical MCHF radial functions. Computer Physics Communications, 9(6):381–391, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900181.

Fowler:1978:FFI

[FSR78] R. H. Fowler, J. Smith, and J. A. Rome. FIFPC — a Fast Ion Fokker–Planck Code. Computer Physics Communications, 13(5– 6):323–340, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900309.

Faisal:1971:PCS

[FT71] F. H. M. Faisal and A. L. V. Tench. A program for calculating the static interaction potential between an electron and a diatomic molecule. Computer Physics Communications, 2(5):261–271, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900026.

Feng:1975:CHO

[FT75] Da Hsuan Feng and T. Tamura. Calculations of harmonic oscillator brackets. Computer Physics Communications, 10(2):87– 97, August 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900752.

Funsten:1979:NVD

[Fun79a] H. O. Funsten. A new version of DWPI (inelastic pion-nucleus scattering) to incorporate microscopic form factors and differing proton and neutron radii. Computer Physics Communications, 16(3):395– 400, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900468.

Funsten:1979:NVP

[Fun79b] H. O. Funsten. A new version of PIRK (elastic pion-nucleus scattering) to handle differing proton and neutron radii. Computer Physics Communications, 16(3):389–393, March 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900456.

Gross:1974:WSM

[GA74] R. Gross and Y. Accad. Weizmann shell model computational code. Computer Physics Communications, 8(2):101–117, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900605.

Garbow:1974:EPM

[Gar74] Burton S. Garbow. EISPACK — A package of matrix eigensystem routines. Computer Physics Communications, 7(4):179– 184, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900861.

Gault:1973:DEP

[GCCB73] F. D. Gault, B. Chadwick, C. S. Cooper, and B. H. Bransden. The Durham elementary particle data bank, data analysis and model comparison project. Computer Physics Communications, 5(2):98– 103, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900258.

Genz:1972:AMQ

[Gen72] A. Genz. An adaptive multidimensional quadrature procedure. Computer Physics Communications, 4(1):11–15, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900240.

Glass:1976:ACE

[GH76] R. Glass and A. Hibbert. Adaptation of CIV3 to evaluate hyperfine structure. Computer Physics Communications, 11(1): 125–140, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900448.

Glass:1978:REM

[GH78] R. Glass and A. Hibbert. Relativistic effects in many electron atoms. Computer Physics Communications, 16(1):19–34, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578901054.

Gianturco:1976:NVP

[Gia76] F. A. Gianturco. A new version of a program calculating the static interaction potential between an electron and a diatomic molecule. Computer Physics Communications, 11(2):237–247, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900552.

Giorgilli:1979:CPI

[Gio79] Antonio Giorgilli. A computer program for integrals of motion. Computer Physics Communications, 16(3):331–343, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900407.

Grimm:1976:RDC

[GJ76] R. C. Grimm and J. L. Johnson. Recent developments in the computational aspects of MHD stability. Computer Physics Communications, 12(1):45–52, September/October 1976. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465576900096.

Glass:1978:RME

[Gla78] R. Glass. Reduced matrix elements of tensor operators. Computer Physics Communications, 16(1):11–18, December 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578901042.

Gliss:1978:DPR

[Gli78] B. Gliss. Design principles of a resource sharing real-time-system. Computer Physics Communications, 15(5):325–327, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900619.

Gloeckner:1976:UFS

[GMP76] D. H. Gloeckner, M. H. Macfarlane, and Steven C. Pieper. The use of first and second derivatives in optical model parameter searches. Computer Physics Communications, 11(3):299–312, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900175.

Graves-Morris:1975:CCA

[GMR75a] P. R. Graves-Morris and D. E. Roberts. Calculation of Canterbury approximants. Computer Physics Communications, 10(4):234– 244, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900685.

Graves-Morris:1975:SPR

[GMR75b] P. R. Graves-Morris and D. E. Roberts. A subroutine and procedure for the rapid calculation of simple off-diagonal rational approximants. Computer Physics Communications, 9(1):46–50, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900557.

Godefroid:1978:AAC

[God78] M. Godefroid. An adaptation of acrz to calculate electric quadrupole oscillator strengths. Computer Physics Communications, 15(3–4):275–282, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900978.

Goldberg:1970:CBM

[Gol70] J. Goldberg. Conversion of binary magnetic tapes. Computer Physics Communications, 1(6):420–424, December 1970. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465570900159.

Golden:1978:ESI

[Gol78] L. B. Golden. Exact Slater integrals. Computer Physics Communications, 14(3–4):255–260, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900188.

Goorevich:1974:PAS

[Goo74] L. Goorevich. A program for the analytic simulation of extensive air showers. Computer Physics Communications, 7(6):344– 352, June 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900320.

Gosman:1978:BRB

[Gos78] A. D. Gosman. Book review: Proceedings of the fifth international conference on numerical methods in fluid dynamics: edited by A. I. Van de Vooren and P. J. Zanbergen, Springer-Verlag, 1976. $15.20. Computer Physics Communications, 14(3–4):313– 314, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900267.

Gaussorgues:1975:MMT

[GPS75] C. Gaussorgues, R. D. Piacentini, and A. Salin. Multistate molecular treatment of atomic collisions in the impact parameter approximation. I. Integration of coupled equations and calculation of transition amplitudes for the straight line case. Computer Physics Communications, 10(4):223–233, October 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900673.

Garpman:1978:STP

[GR78] S. Garpman and J. Randrup. Statistical tests for pseudorandom number generators. Computer Physics Communications, 15(1–2): 5–13, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900802.

Grant:1971:TP

[Gra71] I. P. Grant. Test programs. Computer Physics Communications, 2(6):383, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900312.

Grant:1972:CFP

[Gra72] I. P. Grant. CFPJJ-fractional parentage coefficients for equivalent electrons in jj-coupling. Computer Physics Communications, 4(3):377–381, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901002.

Grant:1973:GPC

[Gra73a] I. P. Grant. A general program to calculate angular momentum coefficients in relativistic atomic structure. Computer Physics Communications, 5(4):263–282, April 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900696.

Grant:1973:PCG

[Gra73b] I. P. Grant. A program to calculate a general recoupling coefficient. Computer Physics Communications, 5(2):161, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900325.

Graham:1976:CS

[Gra76a] Eric Graham. Convection in stars. Computer Physics Communications, 12(1):121–124, September/October 1976. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465576900151.

Grant:1976:PCA

[Gra76b] I. P. Grant. A program to calculate angular momentum coefficients in relativistic atomic structure — revised version. Computer Physics Communications, 11(3):397–405, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900266.

Grant:1978:BRB

[Gra78] I. P. Grant. Book review: The Hartree–Fock method for atoms. A numerical approach: Charlotte Froese Fischer, John Wiley and Sons, Inc., 1977. pp. xi + 308. Price: £17.20/$29.00. Computer Physics Communications, 14(5–6):447–448, July/August 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900073.

Grant:1979:RAS

[Gra79] I. P. Grant. Relativistic atomic structure calculations. Computer Physics Communications, 17(1–2):149–161, April/May 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900778.

Groves:1972:CSM

[Gro72] J. L. Groves. Computer simulation of Mössbauer scattering spectra. Computer Physics Communications, 3(4):339–350, July 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900835.

Gruber:1975:HBM

[Gru75] R. Gruber. Hymnia — band matrix package for solving eigenvalue problems. Computer Physics Communications, 10(1):30– 41, July 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900880.

Gruschel:1979:FIR

[Gru79] W. J. Gruschel. Freint, an integration routine calculating Fresnel diffraction. Computer Physics Communications, 16(2): 175–179, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900869.

Guest:1974:EEC

[Gue74] P. G. Guest. Errors in eigenvalues calculated by the Numerov– Cooley method. Computer Physics Communications, 8(1):31– 34, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900824.

Gastmans:1979:SED

[GVV79] R. Gastmans, A. Van Proeyen, and P. Verbaeten. Symbolic evaluation of dimensionally regularized Feynman diagrams. Computer Physics Communications, 18(2):201–203, October/ November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901127.

Hansen:1978:CPG

[Han78a] Flemming Yssing Hansen. I. A computer program for generation of a complete set of coordinates and force matrices for normal mode calculations of crystals and molecules. Computer Physics Communications, 14(3–4):193–218, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900152.

Hansen:1978:CPN

[Han78b] Flemming Yssing Hansen. I: a computer program for normalization and instrument correction of neutron diffraction data on non-crystalline materials to obtain the static structure factor. Computer Physics Communications, 15(5):401–415, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900693.

Hansen:1978:ICPa

[Han78c] Flemming Yssing Hansen. II. A computer program for calculation of parameters necessary for the computation of reliable pair distribution functions of non-crystalline materials from limited diffraction data. Computer Physics Communications, 15(5):417– 429, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557890070X.

Hansen:1978:IPC

[Han78d] Flemming Yssing Hansen. II. A program for computing normal modes of molecules, crystal phonon dispersion relations and structure factors for neutron inelastic scattering. Computer Physics Communications, 14(3–4):219–243, May/June 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900164.

Hansen:1978:ICPb

[Han78e] Flemming Yssing Hansen. III. A computer program for calculation of reliable pair distribution functions of non-crystalline materials from limited diffraction data. Computer Physics Communications, 15(5):431–435, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900711.

Hansen:1978:IFC

[Han78f] Flemming Yssing Hansen. III. A force constant adjuster program to obtain least squares fit to observed frequencies of molecules and crystals. Computer Physics Communications, 14(3–4):245– 254, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900176.

Hasse:1976:ODW

[Has76] Rainer W. Hasse. One-dimensional wave packet solutions of timedependent frictional or optical potential schrödinger equations. Computer Physics Communications, 11(3):353–362, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900229.

Hawkes:1973:CCC

[Haw73] P. W. Hawkes. Compilation of a catalogue of computer programs in electron optics. Computer Physics Communications, 5 (6):399, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900751.

Hazony:1972:HCL

[Haz72] Yehonathan Hazony. Hierarchical computing in laboratory experiments: Potential and pitfalls. Computer Physics Communications, 4(3):279–288, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900884.

Hostens:1978:IEM

[HD78] R. Hostens and G. De Mey. An integral equation method for the numerical solution of Laplace’s equation without Green’s function. Computer Physics Communications, 16(1):5–9, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578901030.

Herbison-Evans:1971:DOD

[HE71] D. Herbison-Evans. Derivation of the orbit of a double star from observations made with an intensity interferometer. Computer Physics Communications, 2(2):59–78, February/March 1971. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465571900166.

Herbison-Evans:1973:LCV

[HE73] D. Herbison-Evans. The light curve of a variable star subject to orbital tidal distortion. Computer Physics Communications, 5(5): 315–327, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557390060X.

Heiss:1972:CCS

[Hei72] P. Heiss. Computation of cross sections and polarizations for nuclear reactions, in which only spin 1/2 particles are involved. Computer Physics Communications, 4(3):371–376, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900999.

Herbison-Evans:1971:AVS

[HEL71] D. Herbison-Evans and N. R. Lomb. Analysis of a variable spectroscopic double star. Computer Physics Communications, 2(6): 368–380, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900294.

Henry:1975:RCS

[Hen75] Ronald J. W. Henry. Rovibrational cross sections from reactance matrices calculated in adiabatic nuclei approximation. Computer Physics Communications, 10(6):375–384, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900417.

Hird:1975:DQS

[HH75] B. Hird and K. H. Huang. Deformed quasiparticle states in a Woods–Saxon potential and coupled to rotational states of the core. Computer Physics Communications, 10(5):293–303, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900995.

Hawk:1979:FDS

[HH79] I. L. Hawk and D. L. Hardcastle. Finite-difference solution to the Schrödinger equation for the helium isoelectronic sequence. Computer Physics Communications, 16(2):159–166, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900845.

Hibbert:1970:GPC

[Hib70] A. Hibbert. A general program for calculating angular momentum integrals in atomic structure. Computer Physics Communications, 1(5):359–377, September 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900378.

Hibbert:1971:NVG

[Hib71] A. Hibbert. A new version of a general program to calculate angular momentum integrals in atomic structure. Computer Physics Communications, 2(4):180–190, May/June 1971. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465571900518.

Hibbert:1974:AGPb

[Hib74a] A. Hibbert. Adaptation of a general program to calculate angular momentum integrals in atomic structure: Inclusion of the checking of the configuration data. Computer Physics Communications, 8(4):329–332, November 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900228.

Hibbert:1974:AGPa

[Hib74b] A. Hibbert. Adaptation of a general program to calculate angular momentum integrals in atomic structure: Inclusion of the one-electron part of the Hamiltonian. Computer Physics Communications, 7(6):318–326, June 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900290.

Hibbert:1975:CGP

[Hib75] A. Hibbert. CIV3 — A general program to calculate configuration interaction wave functions and electric-dipole oscillator strengths. Computer Physics Communications, 9(3):141– 172, March 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575901034.

Hibbert:1979:BRB

[Hib79] A. Hibbert. Book review: Statistical and computational methods in data analysis: Siegmund Brandt, North-Holland, 1976. $35.95/ Dfl. 90.00. Computer Physics Communications, 16(2):281, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900948.

Harrison:1973:CER

[HIM73] D. J. Harrison, A. C. Irving, and A. D. Martin. Computation of effective Regge trajectories for high energy two-body reactions. Computer Physics Communications, 5(2):153–160, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900313.

Hird:1973:NOP

[Hir73] B. Hird. Nilsson orbits for a particle in a Woods–Saxon potential with Y_{2}^{0} and Y_{4}^{0} deformations, and coupled to core rotational states. Computer Physics Communications, 6(1):30– 37, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900209.

Hummer:1973:NEF

[HKK73] D. G. Hummer, C. V. Kunasz, and P. B. Kunasz. Numerical evaluation of the formal solution of radiative transfer problems in spherical geometries. Computer Physics Communications, 6(1): 38–57, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900210.

Hirschi:1975:CCC

[HLS75] S. Hirschi, E. Lomon, and N. Spencer. COCHASE, a code for coupled channel Schrödinger equations. Computer Physics Communications, 9(1):11–30, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900533.

Hoffstein:1971:SBSa

[HM71a] V. Hoffstein and O. Moller. Symmetry and band structure I. Selection of reciprocal lattice vectors. Computer Physics Communications, 2(1):11–16, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900099.

Hoffstein:1971:SBSb

[HM71b] V. Hoffstein and O. Moller. Symmetry and band structure II. Storage and retrieval of group theoretical information. Computer Physics Communications, 2(1):17–25, January 1971. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465571900105.

Hoffstein:1971:SBSc

[HM71c] V. Hoffstein and O. Moller. Symmetry and band structure III. Construction of symmetrized Hamiltonian matrix. Computer Physics Communications, 2(1):26–32, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900117.

Hoare:1970:BRB

[Hoa70] C. A. R. Hoare. Book review: Computer languages: a practical guide to the chief programming languages: Peter C. Sanderson, M. A. Newnes-Butterworths £2.15s. Computer Physics Communications, 1(6):477, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900275.

Hodgson:1976:MER

[Hod76] R. J. W. Hodgson. Matrix elements of the reaction matrix in finite nuclei. Computer Physics Communications, 11(1):113–124, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900436.

Hoffstein:1971:PCL

[Hof71] V. Hoffstein. Program for calculating LEED intensities using band structure-matching formalism. Computer Physics Communications, 2(6):341–352, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900269.

Hoffstein:1974:PCL

[Hof74] V. Hoffstein. Program for calculating LEED intensities based on the inelastic collision model: I. Matrix inversion method. Computer Physics Communications, 7(1):50–62, January 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900563.

Holt:1979:CCS

[Hol79] A. R. Holt. The calculation of certain second Born integrals. Computer Physics Communications, 17(1–2):171–173, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900791.

Hughes:1974:TSI

[HR74a] M. H. Hughes and A. P. V. Roberts. Timer — a software instrumentation routine for making timing measurements. Computer Physics Communications, 8(2):118–122, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900617.

Hughes:1974:ORF

[HR74b] M. H. Hughes and K. V. Roberts. Olympus restart facilities. Computer Physics Communications, 8(2):123–129, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900629.

Hoffstein:1972:SPC

[HRB72] V. Hoffstein, D. K. Ray, and M. Belakhovsky. Symmetrized program for calculating energy bands and electronic structure of solids. Computer Physics Communications, 4(3):361–370, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900987.

Hughes:1975:OCU

[HRL75] M. H. Hughes, K. V. Roberts, and G. G. Lister. OLYMPUS control and utility package for the CDC 6500. Computer Physics Communications, 10(3):167–181, September 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900041.

Hughes:1975:OPP

[HRR75] M. H. Hughes, K. V. Roberts, and P. D. Roberts. Olympus and preprocessor package for an IBM 370/165. Computer Physics Communications, 9(1):51–58, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900569.

Hubbard:1970:AVC

[Hub70a] Lincoln B. Hubbard. Allowed values of coupled angular momentum and i-spin for nucleons in a single shell in j-j coupling. Computer Physics Communications, 1(6):453–456, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900202.

Hubbard:1970:CFP

[Hub70b] Lincoln B. Hubbard. Coefficients of fractional parentage in j–j coupling in the isospin representation. Computer Physics Communications, 1(4):225–231, April 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557090038X.

Hubert:1970:SNWa

[Hub70c] A. Hubert. Symmetric Néel walls in thin magnetic films. Computer Physics Communications, 1(5):343–348, September 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900354.

Hubert:1970:SNWb

[Hub70d] A. Hubert. Symmetric Néel walls in thin magnetic films. An adaptation to increase the range of convergence. Computer Physics Communications, 1(6):468, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900238.

Hubbard:1971:GRD

[Hub71] Lincoln B. Hubbard. Gamma-radiation dosimetry for arbitrary source and target geometry. Computer Physics Communications, 2(7):449–454, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900385.

Hubbard:1973:FCG

[Hub73] Lincoln B. Hubbard. First collision gamma-ray dose. Computer Physics Communications, 6(5):240–242, November 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900957.

Hughes:1971:SPE

[Hug71] M. H. Hughes. Solution of Poisson’s equation in cylindrical coordinates. Computer Physics Communications, 2(3):157–167, April 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900476.

Hicks:1977:LNI

[HW77] H. R. Hicks and J. W. Wooten. Linear and nonlinear ideal MHD codes-V103. Computer Physics Communications, 13(2):117– 135, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900376.

Irwin:1974:PER

[IL74] D. J. G. Irwin and A. E. Livingston. A program for the extraction of radiative lifetimes from experimental beam-foil intensity decay data. Computer Physics Communications, 7(2):95– 113, February 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900447.

Islam:1973:RCF

[Isl73] Md. M. Islam. A routine for calculating the form-factor and the electrical resistivity of liquid N.F.E. metals. Computer Physics Communications, 5(4):299–303, April 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900726.

Jabs:1973:AFA

[Jab73] Arthur Jabs. Approximation formula and Algol program of the Lorentz-invariant momentum-space integral for particles of equal masses. Computer Physics Communications, 5(3):217–224, March 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557390101X.

Jadach:1975:RGM

[Jad75] S. Jadach. Rapidity generator for Monte–Carlo calculations of cylindrical phase space. Computer Physics Communications, 9(5): 297–304, May 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900818.

Jamieson:1970:PCR

[Jam70] M. J. Jamieson. A program to calculate the radial parts of interaction matrix elements between two hydrogenic wave functions as power series. Computer Physics Communications, 1(6):437– 439, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900172.

Jackson:1977:MDC

[JB77] D. P. Jackson and J. H. Barrett. Modelling displacement correlations in computer simulations of particle-solid collisions. Computer Physics Communications, 13(3):157–166, September/ October 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900108.

Jesshope:1979:SSS

[Jes79] C. R. Jesshope. SIPSOL — A suite of subprograms for the solution of the linear equations arising from elliptic partial differential equations. Computer Physics Communications, 17(4):383–391, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901012.

Joachain:1973:ETE

[Joa73] Charles J. Joachain. Eikonal theory of electron-atom collisions. Computer Physics Communications, 6(6):358–371, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900441.

Johnson:1973:CFV

[Joh73] C. H. J. Johnson. The computation of fourth virial coefficients for pairwise-additive spherically symmetric interaction potentials. Computer Physics Communications, 6(2):65–75, August 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900520.

Johnson:1978:EEP

[Joh78] C. H. J. Johnson. On the estimation of the equilibrium properties of the kinetic Ising model of ferromagnetism. Computer Physics Communications, 14(3–4):185–192, May/June 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900140.

Jones:1972:NSE

[Jon72] D. J. Jones. The numerical solution of elliptic equations by the method of lines. Computer Physics Communications, 4(2):165– 172, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900033.

Jones:1974:MCA

[Jon74] Michael Jones. A method for calculating the algebra of matrix elements for photoionization and line radiation. Computer Physics Communications, 7(7):353–367, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900666.

Jenkins:1978:CFD

[JP78] H. D. B. Jenkins and K. F. Pratt. Calculation of the first derivatives of Madelung constants with respect to cell lengths. Computer Physics Communications, 13(5–6):341–348, January/ February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900310.

Jarmie:1978:LVV

[JPB78] Nelson Jarmie, Michael S. Pindzola, and H. Bichsel. Limits of validity for the Vavilov energy straggling calculation. Computer Physics Communications, 13(5–6):317–321, January/ February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900292.

James:1975:MSF

[JR75] F. James and M. Roos. Minuit — a system for function minimization and analysis of the parameter errors and correlations. Computer Physics Communications, 10(6):343–367, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900399.

Johnson:1972:CTV

[JS72] C. H. J. Johnson and T. H. Spurling. On the computation of third virial coefficients. Computer Physics Communications, 3 (1):69–72, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900562.

Junginger:1972:CEP

[JVS72] H.-G. Junginger, W. Van Haeringen, and M. F. H. Schlurmans. Computation of electronic properties of solids. Computer Physics Communications, 3(S1):127–136, September 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572901208.

Kaczarowski:1977:CPC

[Kac77] R. Kaczarowski. A computer program for calculation of the Coriolis effect in odd-A nuclei. Computer Physics Communications, 13(1):

63–69, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900285.

Karlov:1978:SGD

[Kar78] A. A. Karlov. Software for graphic display systems. Computer Physics Communications, 15(5):329–334, November 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900620.

Kasman:1975:MCS

[Kas75] J. Kasman. Monte Carlo simulation of the diffractive excitation model. Computer Physics Communications, 9(3):182–192, March 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575901058.

Kolman:1973:FIM

[KB73] Bernard Kolman and Robert E. Beck. Freudenthal’s inner multiplicity formula. Computer Physics Communications, 6(1):24– 29, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900192.

Kuo:1976:SRI

[KCA76] C. T. K. Kuo, T. W. Cadman, and R. J. Arsenault. Sequential random integer generator. Computer Physics Communications, 12(2):163–171, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900655.

Kirkegaard:1972:PVP

[KE72] Peter Kirkegaard and Morten Eldrup. POSITRONFIT: a versatile program for analysing positron lifetime spectra. Computer Physics Communications, 3(3):240–255, April/May 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900707.

Kirkegaard:1974:PEN

[KE74] Peter Kirkegaard and Morten Eldrup. Positronfit extended: a new version of a program for analysing position lifetime spectra. Computer Physics Communications, 7(7):401–409, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900708.

Kelly:1979:MBP

[Kel79] H. P. Kelly. Many body perturbation calculations of photoionization. Computer Physics Communications, 17(1–2):99–111, April/ May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900730.

Kankowsky:1971:PCC

[KF71] R. Kankowsky and D. Fick. A program to calculate complex phase shifts and mixing parameters of elastic scattering of spin 1/2 particles on spin 1/2 targets. Computer Physics Communications, 2(4):223–230, May/June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900567.

Kobe:1978:EGS

[KH78] S. Kobe and A. Hartwig. Exact ground state of finite amorphous Ising systems. Computer Physics Communications, 16(1): 1–4, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578901029.

Killingbeck:1979:SAP

[Kil79] J. Killingbeck. Some applications of perturbation theory to numerical integration methods for the Schrödinger equation. Computer Physics Communications, 18(2):211–214, October/ November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901140.

Kajantie:1971:EVP

[KK71] K. Kajantie and V. Karimäki. The evaluation of the volume of the phase space of N particles. Computer Physics Communications, 2(4):207–213, May/June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900543.

Kunasz:1975:TLR

[KK75] C. V. Kunasz and P. B. Kunasz. Transfer of line radiation in optically thick media allowing for transport of excitation energy: The resonant doublet. Computer Physics Communications, 10(5):304– 340, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575901009.

Konig:1977:EDB

[KK77] E. König and S. Kremer. Energy diagrams based on complete ligand field calculations for arbitrary point groups in strong-field coupling. Computer Physics Communications, 13(2):89–99, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465577900340.

Kadashevich:1972:AMU

[KKN^{+}72] V. I. Kadashevich, I. A. Kondurov, S. N. Nickolaev, Yu. F. Ryabov, and V. I. Vinogradov. Application of multi-users online system to nuclear and solid-state physics at Gatchina’s reactor WWR-M. Computer Physics Communications, 4(3):289– 293, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900896.

Kluge:1972:NFG

[KL72] G. Kluge and K. G. Lenhart. Numerical fits for the geomagnetic shell parameter. Computer Physics Communications, 3 (1):36–41, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900525.

Klapisch:1971:PAW

[Kla71] M. Klapisch. A program for atomic wavefunction computations by the parametric potential method. Computer Physics Communications, 2(5):239–260, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900014.

Klotz:1975:GPC

[Klo75a] W.-D. Klotz. A general program to calculate the matrix of the spinorbit interaction. Computer Physics Communications, 9(2):102– 116, February 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900272.

Klotz:1975:RME

[Klo75b] Wolf-Dieter Klotz. Reduced matrix elements of summations of oneparticle tensor products. Computer Physics Communications, 10 (1):56–66, July 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900909.

Kluge:1972:DCM

[Klu72a] G. Kluge. Direct computation of the magnetic shell parameter. Computer Physics Communications, 3(1):31–35, January/ February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900513.

Kluge:1972:GFM

[Klu72b] G. Kluge. Geomagnetic field models: Scalar and vector potential, induction vector and its gradient tensor computed by a common algorithm. Computer Physics Communications, 4(3):347–360, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900975.

Kerr:1974:PCW

[KM74] I. F. Kerr and C. H. B. Mee. Program for calculating work functions from photoelectric data. Computer Physics Communications, 7(7): 419–427, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900721.

Kitowski:1979:CSH

[KM79a] Jacek Kitowski and Jacek Mościński. Computer simulation of heuristic reinforcement-learning systems for nuclear power plant load changes control. Computer Physics Communications, 18(3):339–352, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900043.

Kruger:1979:CNG

[KM79b] P. B. Kruger and R. M. Mayer. Calculation of the nucleation and growth of defect clusters. Computer Physics Communications, 18(3):385–399, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900080.

Klucker:1973:KKA

[KN73] R. Klucker and U. Nielsen. Kramers–Kronig analysis of reflection data. Computer Physics Communications, 6(4):187–193, October 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900994.

Kunc:1979:LDZb

[KN79a] K. Kunc and O. H. Nielsen. Lattice dynamics of zincblende structure compounds II. Shell model. Computer Physics Communications, 17(4):413–422, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901048.

Kunc:1979:LDZa

[KN79b] K. Kunc and O. Holm Nielsen. Lattice dynamics of zincblende structure compounds using deformation-dipole model and rigid ion model. Computer Physics Communications, 16(2):181–197, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900870.

Knox:1970:CWH

[Kno70] H. O. Knox. Computation of wavefunctions for the helium isoelectronic sequence. Computer Physics Communications, 1(3):167– 180, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900032.

Kolbig:1972:PCL

[Köl72a] K. S. Kölbig. Programs for computing the logarithm of the gamma function, and the digamma function, for complex argument. Computer Physics Communications, 4(2):221–226, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900124.

Kolbig:1972:RCC

[Köl72b] K. S. Kölbig. Remarks on the computation of Coulomb wavefunctions. Computer Physics Communications, 4(2):214–220, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900112.

Kolpakov:1978:CHE

[Kol78] I. F. Kolpakov. CAMAC high energy physics electronics hardware. Computer Physics Communications, 15(5):341–345, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900644.

Konrad:1978:ELP

[Kon78a] A. Konrad. An evaluation of the LACC program. Computer Physics Communications, 14(3–4):177–184, May/June 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900139.

Konrad:1978:LAC

[Kon78b] A. Konrad. A linear accelerator cavity code based on the finite element method. Computer Physics Communications, 13(5– 6):349–362, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900322.

Katsuki:1978:CPS

[KPH78] S. Katsuki, P. Palting, and S. Huzinaga. A compact program of the SCF–Xα scattered wave method. Computer Physics Communications, 14(1–2):13–70, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900498.

Kuo-Petravic:1979:PGI

[KPP79] G. Kuo-Petravic and M. Petravic. A program generator for the incomplete LU decomposition-conjugate gradient (ILUCG) method. Computer Physics Communications, 18(1):13–25, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900213.

Kuo-Petravic:1976:TPT

[KPPR76] L. G. Kuo-Petravic, M. Petravic, and K. V. Roberts. Tranal — a program for the translation of Symbolic Algol into Symbolic Algol II. Computer Physics Communications, 11(1):5–25, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900369.

Konrad:1973:FEP

[KS73] A. Konrad and P. Silvester. A finite element program package for axisymmetric scalar field problems. Computer Physics Communications, 5(6):437–455, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900817.

Kerner:1975:CES

[KS75a] W. Kerner and J. Steuerwald. A contribution to the efficient solution of extensive symbolic computations. Computer Physics Communications, 9(6):337–349, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900144.

Kisak:1975:FEP

[KS75b] E. Kisak and P. Silvester. A finite-element program package for magnetotelluric modelling. Computer Physics Communications, 10(6):421–433, December 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900442.

Konrad:1975:FEP

[KS75c] A. Konrad and P. Silvester. A finite element program package for axisymmetric vector field problems. Computer Physics Communications, 9(3):193–204, March 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557590106X.

Kaiser:1978:CAL

[KS78] F. Kaiser and H. M. Staudenmaier. Computer aided laboratory instruction. Computer Physics Communications, 15(5):335–339, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900632.

Kolbig:1979:PSS

[KS79] K. S. Kölbig and F. Schwarz. A program for solving systems of homogeneous linear inequalities. Computer Physics Communications, 17(4):375–382, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901000.

Katsura:1978:AHP

[KSSA78] Shigetoshi Katsura, Masaru Sampei, Masayuki Suzuki, and Yoshihiko Abe. Application of the HLISP program for the irreducible representation matrix of the permutation groups and the partition function of the dilute linear Heisenberg chain. Computer Physics Communications, 13(5–6):311–315, January/ February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900280.

Kittel:1970:MCG

[KVW70] W. Kittel, L. Van Hove, and W. Wojcik. A Monte Carlo generation method with importance sampling for high energy collisions of hadrons. Computer Physics Communications, 1(6):425– 436, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900160.

Labarthe:1979:GDN

[Lab79a] J. J. Labarthe. I. Generator of determinantal non-relativistic atomic states from spectroscopic notation, computation of matrix elements. Computer Physics Communications, 16(3):285– 299, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900365.

Labarthe:1979:IGA

[Lab79b] J. J. Labarthe. II. Generator of atomic excited terms from angular considerations. Computer Physics Communications, 16(3): 301–309, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900377.

Labarthe:1979:IAA

[Lab79c] J. J. Labarthe. III. Analytic approximations of radial orbitals for multiconfigurational Hartree–Fock computations. Computer Physics Communications, 16(3):311–323, March 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900389.

Labarthe:1979:IAN

[Lab79d] J. J. Labarthe. IV. Approximation of numerical orbitals by Slater functions. Computer Physics Communications, 16(3):325– 329, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900390.

Lackner:1976:CIM

[Lac76] K. Lackner. Computation of ideal MHD equilibria. Computer Physics Communications, 12(1):33–44, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900084.

Laing:1974:BRB

[Lai74] E. W. Laing. Book review: Computational Physics: by David Potter (John Wiley, 1973), £6.95. Computer Physics Communications, 7(3):174, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900083.

Lang:1970:SIA

[Lan70] James Lang. Spectral intensity, angular distribution and polarisation of synchrotron radiation from a monoenergetic electron. Computer Physics Communications, 1(6):440–444, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900184.

Leutz:1976:CAC

[LB76] R. K. Leutz and R. Bauer. Computation of the anisotropic cubic elastic Green’s tensor function and the elastic energy coefficients of point defects in crystals. Computer Physics Communications, 11(3):339–351, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900217.

Lin:1978:NEM

[LB78] S. L. Lin and J. N. Bardsley. The null-event method in computer simulation. Computer Physics Communications, 15(3–4):161–163, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900905.

Liberman:1971:RSC

[LCW71] D. A. Liberman, D. T. Cromer, and J. T. Waber. Relativistic self-consistent field program for atoms and ions. Computer Physics Communications, 2(2):107–113, February/March 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900208.

Lewis:1970:HI

[Lew70a] M. N. Lewis. Hydrogenic R^{k} integrals. Computer Physics Communications, 1(5):325–336, September 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900330.

Lewis:1970:SPS

[Lew70b] M. N. Lewis. Single-particle substitution sums in the secondorder Z-expansion theory of atomic energies. Computer Physics Communications, 1(4):265–276, April 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900421.

Lewis:1979:EME

[Lew79] M. N. Lewis. Z-expansion of matrix elements of one-electron operators for many-electron atoms. Computer Physics Communications, 18(1):109–121, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900286.

Lejeune:1971:CBM

[LJ71] A. Lejeune and J. P. Jeukenne. Computation of Brody–Moshinsky brackets. Computer Physics Communications, 2(4):231–237, May/ June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900579.

Lea:1972:DCS

[LJLC72] Suzanne M. Lea, Vinaya Joshi, and A. B. Lopez-Cepero. Differential cross sections for electric quadrupole Coulomb excitation I. Computer Physics Communications, 3(2):118–119, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900598.

Lapostolle:1972:TDC

[LL72] P. Lapostolle and R. Le Bail. Two-dimensional computer simulation of high intensity proton beams. Computer Physics Communications, 4(3):333–338, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557290094X.

Lubkin:1979:IQM

[LL79] Elihu Lubkin and Thelma Lubkin. An inversion of quantum mechanics. Computer Physics Communications, 16(2):207–219, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900894.

Lotrian:1976:CER

[LLC76] J. Lotrian, M. Leriche, and J. Cariou. Calculation of the energy response of a spectrometer. Computer Physics Communications, 12(2):231–236, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900710.

Lang:1971:PNL

[LM71] J. Lang and R. Müller. A procedure for nonlinear least squares refinement in adverse practical conditions. Computer Physics Communications, 2(2):79–86, February/March 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465571900178.

Long:1976:GOD

[LNS76] J. W. Long, A. A. Newton, and M. C. Sexton. Glowcode: a one-dimensional code for the simulation of plasma afterglows. Computer Physics Communications, 12(2):213–230, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900709.

Linkens:1978:SUD

[LNV78] A. Linkens, J. Niezette, and J. Vanderschueren. Simulation of ultrasonic degradation of macromolecules in solution. Computer Physics Communications, 15(5):375–386, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557890067X.

Lovitch:1971:BSS

[LR71] L. Lovitch and S. Rosati. Bound state solution of the twonucleon Schrödinger equation with tensor forces. Computer Physics Communications, 2(6):353–359, October/November 1971. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465571900270.

Lovitch:1972:TNE

[LR72] L. Lovitch and S. Rosati. The two-nucleon effective-range parameters with tensor forces. Computer Physics Communications, 4(1):138–146, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900422.

Luke:1979:ESV

[LS79] T. M. Luke and H. E. Saraph. Ecsimpact, a special version of program impact for CDC machines with ex-core memory. Computer Physics Communications, 18(2):287–296, October/November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901206.

Linkens:1978:STS

[LVPG78] A. Linkens, J. Vanderschueren, P. Parot, and J. Gasiot. Simulation of thermally stimulated dipolar processes in dielectrics. Computer Physics Communications, 13(5–6):411–419, January/ February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900383.

Lodhi:1975:SBS

[LW75] M. A. K. Lodhi and B. T. Waak. Solution of bound state problems in nuclear shell models with momentum dependent potentials. Computer Physics Communications, 10(3):182–193, September 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900053.

Light:1979:MPM

[LWSS79] John C. Light, Robert B. Walker, Ellen B. Stechel, and Thomas G. Schmalz. R-matrix propagation methods in inelastic and reactive collisions. Computer Physics Communications, 17(1–2):89– 97, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900729.

Macleod:1972:P

[Mac72] G. R. Macleod. Preface. Computer Physics Communications, 3(S1):vii, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901075.

Maguire:1975:VEC

[Mag75] William C. Maguire. Vibrational energies of CO_{2}. Computer Physics Communications, 10(6):368–374, December 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900405.

Magill:1978:TTD

[Mag78] J. Magill. Trip 1, a time-dependent recombination ionisation package. Computer Physics Communications, 16(1):129–137, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578901169.

Mathier:1972:NSE

[MAK^{+}72] E. Mathier, A. Attanasio, J. Keller, P. Nösberger, A. Bauder, and Hs. H. Günthard. Numerical solution of the eigenvalue problem of infinite matrices in molecular dynamics. Computer Physics Communications, 4(1):20–26, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900264.

Malaviya:1970:CCT

[Mal70] V. Malaviya. Computation of charge transfer probability between protons and excited hydrogen atoms. Computer Physics Communications, 1(6):380–390, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465570900123.

Marquez:1973:NEC

[Mar73] L. Marquez. Nearly exact calculation of the solution of the radial Schrödinger equation. Computer Physics Communications, 5(5): 379–389, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900647.

Maruhn:1976:FFF

[Mar76] J. A. Maruhn. Fourgen: a Fast Fourier Transform program generator. Computer Physics Communications, 12(2):147–162, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900643.

Matsuoka:1972:FFG

[Mat72] Osamu Matsuoka. Field and field gradient integrals based on Gaussian type orbitals. Computer Physics Communications, 3(2):130– 135, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900604.

Mayers:1974:BRB

[May74] D. F. Mayers. Book review: Computational methods in ordinary differential equations: J. D. Lambert (John Wiley, 1972), £5.50. Computer Physics Communications, 7(4):235, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900939.

Mercier:1976:CTT

[MBW76] C. Mercier, J. P. Boujot, and F. Werkoff. Computation of Tokamak transport. Computer Physics Communications, 12(1): 109–119, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557690014X.

McCormack:1975:FRC

[McC75] Francis J. McCormack. FORMAC routine for computing the matrix elements of the linearized cross-collision operators of gaseous mixtures. Computer Physics Communications, 9(1):1–10, January 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900521.

McNamara:1971:BRB

[McN71] Brendan McNamara. Book review: Methods in computational physics, Vol. IX — Plasma Physics: eds. B. Alder, S. Fernbach and M. Rotenberg, Academic Press £10.75. Computer Physics Communications, 2(4):238, May/June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900580.

Munick:1973:CSM

[MGTD73] E. Münick, J. L. Groves, T. A. Tumolillo, and P. G. Debrunner. Computer simulations of Mössbauer spectra for an effective spin S = 1/2 Hamiltonian. Computer Physics Communications, 5(3): 225–238, March 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573901021.

Miketinac:1974:CES

[Mik74] M. J. Miketinac. Calculation of the equilibrium structure and oscillations of polytropic stars pervaded by toroidal magnetic fields. Computer Physics Communications, 7(7):410–418, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557490071X.

Manning:1974:DDE

[MM74] Irwin Manning and G. P. Mueller. Depth distribution of energy deposition by ion bombardment. Computer Physics Communications, 7(2):85–94, February 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900435.

Miller:1973:HRG

[MMK73] D. M. Miller, G. E. Miller, and D. P. Kerr. High resolution graphical plotting on a typewritter terminal. Computer Physics Communications, 6(5):195–197, November 1973. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557390091X.

McDowell:1974:EIE

[MMM74] M. R. C. McDowell, L. Morgan, and V. P. Myerscough. Electron impact excitation cross sections. Computer Physics Communications, 7(1):38–49, January 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900551.

Meyer:1971:SFP

[MNJL71] J. Meyer, R. S. Nahabetian, J. Joseph, and J. Lafoucriere. SU3 fractional parentage in the 1p-shell. Computer Physics Communications, 2(7):420–426, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900348.

Moiseiwitsch:1973:HBP

[Moi73] B. L. Moiseiwitsch. Higher Born, Padé and Schwinger variational methods. Computer Physics Communications, 6(6):372– 376, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900453.

Moore:1971:NVP

[Moo71a] C. Fred Moore. A new version of a printer-plotter routine. Computer Physics Communications, 2(7):470, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900403.

Moore:1971:PPR

[Moo71b] C. Fred Moore. Printer-plotter routine. Computer Physics Communications, 2(1):55–57, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900154.

Moores:1971:ASE

[Moo71c] D. L. Moores. Amplitudes for scattering of electrons by hydrogenic and alkali-like atomic systems. Computer Physics Communications, 2(6):360–367, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900282.

Moores:1972:BRB

[Moo72] D. L. Moores. Book review: The Calculation of Atomic Collision Processes: Kenneth Smith, Wiley — Interscience £5.50. Computer Physics Communications, 4(1):147–148, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900434.

Moon:1979:IMD

[Moo79a] Wooil Moon. An interpolation method for data sets with jump discontinuities. Computer Physics Communications, 16(2): 273–280, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900936.

Moon:1979:NEG

[Moo79b] Wooil Moon. Numerical evaluation of geomagnetic dynamo integrals (Elsasser and Adams–Gaunt integrals). Computer Physics Communications, 16(2):267–271, January/February 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900924.

Moraal:1972:PEB

[Mor72] H. Moraal. A program for the extraction of bulk viscosities from sound absorption data. Computer Physics Communications, 3(1):1–8, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900471.

Morgan:1973:EES

[Mor73] K. Morgan. An expansion equation of state subroutine. Computer Physics Communications, 5(1):64–68, January 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046557390012X.

Morton:1976:FDF

[Mor76] K. W. Morton. Finite difference and finite element methods. Computer Physics Communications, 12(1):99–108, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900138.

Morrow:1978:MCI

[Mor78a] Richard A. Morrow. Monte Carlo integration program for the n-particle relativistic phase space integral in invariant variables. Computer Physics Communications, 13(5–6):399–409, January/ February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900371.

Morton:1978:BRB

[Mor78b] K. W. Morton. Book review: Lecture Notes in Physics, Vol. 58: Computing Methods in Applied Sciences: R. Glowinski and J. L. Lions, eds., Springer-Verlag 1976. DM57, $15.20. Computer Physics Communications, 13(5–6):430, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900413.

Maison:1974:SGP

[MP74] Doris Maison and A. Petermann. Subtracted generalized polylogarithms and the SINAC program. Computer Physics Communications, 7(3):121–134, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900022.

McKay:1975:EEL

[MPS75] W. McKay, J. Patera, and R. T. Sharp. Eigenstates and eigenvalues of labelling operators for O(3) bases of U(3) representations. Computer Physics Communications, 10(1):1–10, July 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900867.

Moore:1973:DMT

[MPW73] D. R. Moore, R. S. Peckover, and N. O. Weiss. Difference methods for time-dependent two-dimensional convection. Computer Physics Communications, 6(5):198–220, November 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900921.

Moussa:1973:AIC

[MR73] A. H. Moussa and H. M. A. Radi. Atomic integral containing three odd powers of interelectronic separation coordinates. Computer Physics Communications, 6(2):89–95, August 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900556.

Moscinski:1977:MCC

[MR77] Jacek Mościński and Zbigniew Rycerz. Monte Carlo computations of the order-disorder transition temperature on the lattice model of the binary BCC system. Computer Physics Communications, 13 (1):9–10, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900212.

Macias:1976:PCG

[MRCL76] E. S. Macias, W. D. Ruhter, D. C. Camp, and R. G. Lanier. A program for calculating gamma-gamma directional correlation coefficients and mixing ratios. Computer Physics Communications, 11(1):75–93, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900412.

Manning:1976:APD

[MRW76] Irwin Manning, Mervine Rosen, and J. E. Westmoreland. Adaptation of a program for depth distribution of energy deposition by ion bombardment: Calculation of ion lateral ranges. Computer Physics Communications, 12(3):335–338, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900916.

Makhankov:1972:CIN

[MS72] V. G. Makhankov and B. G. Shchinov. Computer investigation of nonlinear dynamical problems of plasma theory. Computer Physics Communications, 4(3):327–332, December 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900938.

McGreevy:1978:PCE

[MS78] E. McGreevy and A. L. Stewart. A program for calculating elastic scattering phase shifts for an electron colliding with a one-electron target using perturbation theory. Computer Physics Communications, 14(1–2):99–107, March/April 1978. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557890053X.

Moriarty:1976:PCO

[MT76] K. J. M. Moriarty and J. H. Tabor. A program for calculating the observables for single-particle-inclusive production reactions. Computer Physics Communications, 12(3):277–292, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900874.

Moriarty:1979:PAC

[MT79] K. J. M. Moriarty and H. N. Thompson. Program adaptation: To calculate inclusive backward proton cross sections. Computer Physics Communications, 18(1):155–162, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557990033X.

Mueller:1971:REF

[Mue71] G. P. Mueller. Rapid evaluation of four-body cluster contributions. Computer Physics Communications, 2(4):214–222, May/ June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900555.

Masri:1970:PCD

[MW70] F. N. Masri and I. R. Williams. I. Program for calculating degenerate Raman bands of symmetric tops with an adaptation for infrared bands. Computer Physics Communications, 1(5):349–358, September 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900366.

Masri:1971:FPC

[MW71a] F. N. Masri and I. R. Williams. I. A Fortran program for calculating degenerate Raman bands of symmetric tops with an adaptation for infrared bands. Computer Physics Communications, 2(5):298, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900051.

Masri:1971:IFP

[MW71b] F. N. Masri and I. R. Williams. II. A Fortran program for calculating degenerate Raman bands of spherical tops with an adaptation for infrared bands. Computer Physics Communications, 2(5):299, August/September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900063.

Masri:1971:IPC

[MW71c] F. N. Masri and I. R. Williams. II. Program for calculating triply degenerate Raman bands of spherical tops with an adaptation for infrared bands. Computer Physics Communications, 2(2):87–93, February/March 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557190018X.

Neufeld:1972:PAC

[NA72] P. D. Neufeld and R. A. Aziz. Program ACQN to calculate transport collision integrals adapted to run on IBM computers. Computer Physics Communications, 3(3):269–271, April/May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900720.

Nadrchal:1975:SSC

[Nad75] J. Nadrchal. Summer school on computing techniques in physics. Computer Physics Communications, 10(5):260–261, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465575900958.

Nadrchal:1976:RSP

[Nad76] J. Nadrchal. Role of structured programming in physical applications. Computer Physics Communications, 11(2):143–157, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900461.

Nash:1974:HME

[Nas74] J. C. Nash. The Hermitian matrix eigenproblem Hx = eSx using compact array storage. Computer Physics Communications, 8(2):85–94, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900587.

Nedelkov:1972:IPC

[Ned72] I. P. Nedelkov. Improper problems in computational physics. Computer Physics Communications, 4(2):157–164, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900021.

Nerbrant:1978:ASA

[Ner78] P.-O. Nerbrant. An application of Schoonschip for algebraic calculations in quantum chemistry. Computer Physics Communications, 14(5–6):315–318, July/August 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900012.

Nesbet:1973:MVM

[Nes73] R. K. Nesbet. Matrix variational method. Computer Physics Communications, 6(6):275–287, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900374.

Nesbet:1979:MVC

[Nes79] R. K. Nesbet. Matrix variational calculations of electron-atom scattering. Computer Physics Communications, 17(1–2):163–169, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557990078X.

Neto:1975:NCI

[Net75] N. Neto. Numerical calculations of the irreducible representations of space groups. Computer Physics Communications, 9(4):231– 246, April 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900107.

Nilsson:1972:NPE

[Nil72] Sven Gösta Nilsson. Nuclear potential-energy surfaces and the halflives of heavy and super-heavy elements. Computer Physics Communications, 3(S1):92–116, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557290118X.

Nagel:1978:FRC

[NK78] Pierre Nagel and R. D. Koshel. A finite range coupled channel Born approximation code. Computer Physics Communications, 15(3–4):193–226, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900929.

Naccache:1973:PSA

[NM73] P. F. Naccache and M. R. C. McDowell. Phase shift analysis and consistency checks on electron-atom collision data. Computer Physics Communications, 6(2):77–87, August 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900532.

Norcross:1973:NMA

[Nor73a] David W. Norcross. Numerical methods in asymptotic regions. Computer Physics Communications, 6(6):257–264, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900350.

Norgett:1973:BRB

[Nor73b] M. J. Norgett. Book review: Interatomic Potentials and the Simulation of Lattice Defects: edited by P. C. Gehlen, J. R. Beeler and R. J. Jaffee, Plenum Press, 1972, £18.70. Computer Physics Communications, 6(2):97, August 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557390057X.

Nussbaumer:1970:VCC

[Nus70] H. Nussbaumer. Vector coupling coefficients for complex atoms. Computer Physics Communications, 1(3):191–197, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900056.

Nuttall:1973:UAA

[Nut73] J. Nuttall. The use of analyticity in atomic scattering calculations. Computer Physics Communications, 6(6):331–335, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900416.

Nielsen:1979:LDG

[NW79] O. H. Nielsen and W. Weber. Lattice dynamics of group IV semiconductors using an adiabatic bond charge model. Computer Physics Communications, 18(1):101–108, September 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900274.

Olson:1973:SCP

[OBLR73] D. L. Olson, J. L. Blough, T. S. Lakshmanan, and D. A. Rigney. Simulation of chemical profiles during electrotransport. Computer Physics Communications, 5(6):430–436, June 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900805.

OConnell:1974:MLS

[O’C74a] M. J. O’Connell. Multivariate least squares fitting program using modified Gram–Schmidt transformations. Computer Physics Communications, 8(1):56–69, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557490085X.

OConnell:1974:SPS

[O’C74b] M. J. O’Connell. Search program for significant variables. Computer Physics Communications, 8(1):49–55, August 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900848.

OConnell:1976:CLF

[O’C76] M. J. O’Connell. Computing a Laplacian field component from boundary observations only. Computer Physics Communications, 11(2):221–236, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900540.

Okuda:1978:HOM

[OC78] H. Okuda and C. Z. Cheng. Higher order multipoles and splines in plasma simulations. Computer Physics Communications, 14(3–4):169–176, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900127.

Ortiz:1972:RMA

[Ort72] E. L. Ortiz. A recursive method for the approximate expansion of functions in a series of polynomials. Computer Physics Communications, 4(2):151–156, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557290001X.

OHara:1971:TCI

[OS71] H. O’Hara and F. J. Smith. Transport collision integrals for a dilute gas. Computer Physics Communications, 2(1):47–54, January 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900142.

Ortega:1976:INR

[OVQT^{+}76] M. Ortega, A. Vidal-Quadras, M. Tomás, F. Fernández, V. Gandía, and C. Jacquot. Identification of nuclear reactions registered in ionographic detectors. Computer Physics Communications, 11(2):287–297, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900606.

Opik:1970:MPE

[ÖW70] U. Öpik and R. F. Wood. The Madelung potential and electric field intensity for a sodium chloride and for a caesium chloride lattice. Computer Physics Communications, 1(4):281–290, April 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900445.

Pachelski:1972:FTL

[Pac72] Wojciech Pachelski. On the filtering technique of the least squares estimation of parameters. Computer Physics Communications, 4(1):40–46, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557290029X.

Parcell:1973:SDE

[Par73] L. A. Parcell. Solution of differential equations for exchange matrix elements in heavy particle collisions. Computer Physics Communications, 5(4):283–293, April 1973. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900702.

Pasta:1972:HVC

[Pas72] John R. Pasta. The heuristic value of computers in physics. Computer Physics Communications, 3(S1):139–150, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572901221.

Patterson:1973:BRB

[Pat73] T. N. L. Patterson. Book review: Fortran techniques: A. Colin Day, Cambridge University Press, 1972, £2.50. Computer Physics Communications, 5(5):397, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900672.

Payne:1972:CET

[Pay72] A. D. Payne. Computation of the Einstein tensor with FOR-MAC. Computer Physics Communications, 4(1):100–106, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900380.

Payne:1976:CUF

[Pay76] A. D. Payne. Comment on the use of FORMAC73 in general relativity. Computer Physics Communications, 12(2):145–146, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900631.

Parkinson:1972:PNM

[PB72a] J. R. Parkinson and D. T. Birtwistle. A program for normalised Morse functions. Computer Physics Communications, 4(2):257– 261, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900173.

Pirala:1972:RNI

[PB72b] P. Piralä and E. Byckling. Recursive numerical integration of multi-particle phase space with peripheral matrix element. Computer Physics Communications, 4(1):117–127, October 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900409.

Packer:1973:OSF

[PB73] John C. Packer and David F. Brailsford. The OPIT system I. A file-handling scheme for data in large applications programs. Computer Physics Communications, 5(2):123–135, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900283.

Pekeris:1972:G

[Pek72] C. L. Pekeris. Geophysics. Computer Physics Communications, 3(S1):33–49, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901129.

Pelz:1978:PLP

[Pel78] K. Pelz. The programming language ‘PEARL’ and its implementation. Computer Physics Communications, 15(5):317–324, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900607.

Penney:1978:CSA

[Pen78] B. K. Penney. Computer system architecture for laboratory automation. Computer Physics Communications, 15(5):305–316, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900590.

Percival:1973:MCM

[Per73] I. C. Percival. Monte Carlo methods for classical collisions between electrons and atoms. Computer Physics Communications, 6(6):347–357, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557390043X.

Petrasch:1976:NCA

[Pet76a] P. Petrasch. Nonlinear computation of anisotropic elastic fields about straight edge dislocations. Computer Physics Communications, 11(2):279–285, March/May 1976. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557690059X.

Petravic:1976:NMP

[Pet76b] M. Petravić. Numerical modelling of pulsar magnetospheres. Computer Physics Communications, 12(1):9–19, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900060.

Pyper:1978:NPC

[PGB78] N. C. Pyper, I. P. Grant, and N. Beatham. A new program for calculating matrix elements of one-particle operators in jjcoupling. Computer Physics Communications, 15(5):387–400, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900681.

Poenaru:1978:LDM

[PI78] D. N. Poenaru and M. Ivascu. Liquid drop model deformation energies of nuclei with axial symmetry and reflection asymmetry. Computer Physics Communications, 16(1):85–91, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578901121.

Piessens:1973:CRD

[Pie73] R. Piessens. Calculation of the radial distribution of emitters in a cylindrical source. Computer Physics Communications, 5(4):294– 298, April 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900714.

Petravic:1972:APP

[PPR72] M. Petravic, G. Kuo Petravic, and K. V. Roberts. Automatic production of programmes for solving partial differential equations by finite difference methods. Computer Physics Communications, 4(1):82–88, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900355.

Pauli:1975:CPI

[PR75] H. C. Pauli and U. Raff. A computer program for internal conversion coefficients and particle parameters. Computer Physics Communications, 9(6):392–407, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900193.

Preuss:1979:CCO

[Pre79a] E. Preuss. Calculation of crystal orientations using Laue patterns. Computer Physics Communications, 18(2):277–280, October/November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901188.

Preuss:1979:PPL

[Pre79b] E. Preuss. Plot program for Laue patterns and stereographic projections. Computer Physics Communications, 18(2):261–275, October/November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579901176.

Papaconstantopoulos:1974:CCP

[PS74] Dimitrios A. Papaconstantopoulos and Wayne R. Slaughter. Calculation of crystal potentials. Computer Physics Communications, 7(4):207–214, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900903.

Piacentini:1976:MMT

[PS76] R. D. Piacentini and A. Salin. Multistate molecular treatment of atomic collisions in the impact parameter approximation. III — integration of coupled equations and calculation of transition amplitudes for Coulomb trajectories. Computer Physics Communications, 12(2):199–203, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900680.

Piacentini:1977:MMT

[PS77] R. D. Piacentini and A. Salin. Multistate molecular treatment of atomic collisions in the impact parameter approximation. II — calculation of differential cross-sections from the transition amplitudes for the straight line case. Computer Physics Communications, 13(1):57–62, May 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900273.

Pomponiu:1978:FAS

[PS78] C. Pomponiu and M. Sararu. Fourier analysis with splines, a FOR-TRAN program. Computer Physics Communications, 16(1):93– 112, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578901133.

Peterson:1973:EED

[PSBG73] L. R. Peterson, T. Sawada, J. N. Bass, and A. E. S. Green. Electron energy deposition in a gaseous mixture. Computer Physics Communications, 5(4):239–262, April 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900684.

Payne:1974:FPC

[PV74] G. L. Payne and P. L. Von Behren. Fortran program to calculate finite-range no-recoil DWBA transfer cross sections. Computer Physics Communications, 7(1):13–37, January 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046557490054X.

Peckover:1972:CPM

[PW72] R. S. Peckover and N. O. Weiss. Convection in the presence of magnetic fields. Computer Physics Communications, 4(3):339– 344, December 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900951.

Paldus:1973:CGF

[PW73] J. Paldus and H. C. Wong. Computer generation of Feynman diagrams for perturbation theory I. General algorithm. Computer Physics Communications, 6(1):1–7, July 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900167.

Raptis:1978:EFM

[RA78] A. Raptis and A. C. Allison. Exponential-fitting methods for the numerical solution of the Schrödinger equation. Computer Physics Communications, 14(1–2):1–5, March/April 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900474.

Ranft:1972:MCC

[Ran72] J. Ranft. Monte Carlo calculation of hadronic cascades at high energy. Computer Physics Communications, 4(1):59–63, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557290032X.

Rapaport:1974:CGC

[Rap74] D. C. Rapaport. Computer generation of connected graphs. Computer Physics Communications, 8(4):320–328, November 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900216.

Rudnicki-Bujnowski:1975:EFC

[RB75] Georges Rudnicki-Bujnowski. Explicit formulas for Clebsch– Gordan coefficients. Computer Physics Communications, 10(4): 245–250, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0010465575900697.

Rountree:1976:PCG

[RBHW76] S. P. Rountree, T. Burnett, R. J. W. Henry, and C. A. Weatherford. A program to calculate Green’s functions. Computer Physics Communications, 11(1):27–35, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900370.

Roberts:1972:TCF

[RC72] K. V. Roberts and J. P. Christiansen. Topics in computational fluid mechanics. Computer Physics Communications, 3(S1):14– 32, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572901117.

Roberts:1975:ARM

[RCL75] K. V. Roberts, J. P. Christiansen, and J. W. Long. Adiabatic relaxation to 1D MHD pressure equilibrium. EQUIL: a Fortran module and test program. Computer Physics Communications, 10(5):264– 281, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900971.

Ritter:1979:CUN

[RD79] R. P. Ritter and F. Demmig. The computation of unsteady nonequilibrium shock tube flows comprising a tailor-made discretization of boundary layer terms. Computer Physics Communications, 18(2):205–209, October/November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579901139.

Reinhardt:1973:FCE

[Rei73] William P. Reinhardt. Fredholm calculation of electron-atom scattering amplitudes using L^{2} basis sets. Computer Physics Communications, 6(6):303–315, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900398.

Reinhardt:1979:DAM

[Rei79] William P. Reinhardt. L^{2} discretization of atomic and molecular electronic continua: Moment, quadrature and J-matrix techniques. Computer Physics Communications, 17(1–2):1–21, April/ May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557990064X.

Remiddi:1972:RCQ

[Rem72] E. Remiddi. Radiative correction in quantum electrodynamics. Computer Physics Communications, 4(2):193–198, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900070.

Radziemski:1972:CAE

[RFSG72a] Leon J. Radziemski, Jr., Kay J. Fisher, David W. Steinhaus, and Aaron S. Goldman. Calculation of atomic energy level values. Computer Physics Communications, 3(1):9–18, January/ February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900483.

Radziemski:1972:WNC

[RFSG72b] Leon J. Radziemski, Jr., Kay J. Fisher, David W. Steinhaus, and Aaron S. Goldman. Wave number calculation from least-squares level values. Computer Physics Communications, 3(1):19–21, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900495.

Rinker:1979:SDM

[Rin79] G. A. Rinker. Static and dynamic muonic-atom codes — Muon and Rurp. Computer Physics Communications, 16(2):221–242, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900900.

Rycerz:1976:SECa

[RM76a] Zbigniew Rycerz and Jacek Mościński. A study of the electrical conductivity processes in amorphous semiconductors using the computer simulation method I. One-dimensional case. Computer Physics Communications, 11(2):163–168, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900485.

Rycerz:1976:SECb

[RM76b] Zbigniew Rycerz and Jacek Mościński. A study of the electrical conductivity processes in amorphous semiconductors using the computer simulation method II. Two-dimensional case. Computer Physics Communications, 11(2):169–175, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900497.

Rycerz:1976:SECc

[RM76c] Zbigniew Rycerz and Jacek Mościński. A study of the electrical conductivity processes in amorphous semiconductors using the computer simulation method III. Three-dimensional case. Computer Physics Communications, 11(2):177–181, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900503.

Robb:1970:PGN

[Rob70] W. D. Robb. A program to generate numerical orbital functions. Computer Physics Communications, 1(6):457–464, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465570900214.

Roberts:1971:PTC

[Rob71] K. V. Roberts. Practical techniques in computer programming. Computer Physics Communications, 2(7):385–393, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465571900324.

Robb:1972:UMM

[Rob72] W. D. Robb. The use of the R-matrix method in atomic calculations. Computer Physics Communications, 4(1):16–19, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900252.

Robb:1973:PER

[Rob73] W. D. Robb. A program to evaluate the reduced matrix elements of summations of one-particle tensor operators. Computer Physics Communications, 6(3):132–148, September 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900507.

Roberts:1974:IOS

[Rob74] K. V. Roberts. An introduction to the OLYMPUS system. Computer Physics Communications, 7(5):237–243, May 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900253.

Roberts:1979:NMCa

[Rob79a] S. A. Roberts. Numerical modelling of a chemical plasma: I. REACS: a program to generate all reactions which take place in a plasma of given chemical content. Computer Physics Communications, 18(3):353–362, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900055.

Roberts:1979:NMCb

[Rob79b] S. A. Roberts. Numerical modelling of a chemical plasma: II. PLASKEM: a program to predict the variation with time of the number densities of chemical species within a plasma. Computer Physics Communications, 18(3):363–376, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900067.

Roberts:1979:NMCc

[Rob79c] S. A. Roberts. Numerical modelling of a chemical plasma: III. DATSTOR: a program to create a data base containing information on rate coefficients of chemical reactions. Computer Physics Communications, 18(3):377–384, December 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900079.

Routti:1972:IGC

[Rou72] Jorma T. Routti. Interactive graphical computing for simulating and unfolding measured distributions. Computer Physics Communications, 4(1):33–39, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900288.

Ranft:1974:MCP

[RR74] J. Ranft and J. T. Routti. Monte Carlo programs for calculating three-dimensional high-energy (50 MeV–500 GeV) hadron cascades in matter. Computer Physics Communications, 7(6):327– 342, June 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900307.

Rawitscher:1976:EAC

[RR76] G. H. Rawitscher and C. H. Rasmussen. Error analysis of code AROSA for quantal Coulomb excitation calculations. Computer Physics Communications, 11(2):183–198, March/May 1976. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465576900515.

Rundgren:1974:LIC

[RS74] J. Rundgren and A. Salwén. LEED intensity curves by the layerby-layer method and perturbation calculation. Computer Physics Communications, 7(7):369–388, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900678.

Rundgren:1975:SCL

[RS75] J. Rundgren and A. Salwén. Symmetrization and calculation of LEED intensity patterns. Computer Physics Communications, 9(5):312–326, May 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900831.

Roberts:1976:PPP

[RS76] S. A. Roberts and Kenneth Smith. PULSAMP: a program to predict the amplification of nanosecond CO_{2} laser light pulses. Computer Physics Communications, 12(3):323–334, December 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900904.

Rosel:1974:QMC

[RSA74] F. Rösel, J. X. Saladin, and K. Alder. Quantum mechanical coupled channels code for Coulomb excitation. Computer Physics Communications, 8(1):35–48, August 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900836.

Rouse:1978:PCG

[RSL^{+}78] Robert J. Rouse, Jr., Gordon L. Struble, Robert G. Lanier, Lloyd G. Mann, and Edward S. Macias. A program for calculating gamma-gamma directional correlation coefficients and angular distribution coefficients for gamma rays of mixed multipolarities from partially aligned nuclei. Computer Physics Communications, 15(1–2):107–123, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900851.

Russell:1973:ILP

[Rus73] Robert D. Russell. Intermediate-level programming languages for on-line data acquisition and control. Computer Physics Communications, 5(2):89–97, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900246.

Rao:1978:NFP

[RV78] K. Srinivasa Rao and K. Venkatesh. New Fortran programs for angular momentum coefficients. Computer Physics Communications, 15(3–4):227–235, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900930.

Rycerz:1978:CST

[Ryc78] Zbigniew Rycerz. Computer simulation of threshold switching in non-crystalline semiconductors. Computer Physics Communications, 15(1–2):15–21, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900814.

Salem:1975:IAS

[Sal75] M. Salem. ILTHII — analysis of the spectrum of a thermal radioastronomical source. Computer Physics Communications, 9(4): 247–257, April 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900119.

Salem:1976:CFF

[Sal76] M. Salem. Converter of FORTRAN format and data statements to standard form. Computer Physics Communications, 11(2):199–209, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900527.

Salin:1978:CWF

[Sal78] A. Salin. Calculation of wave-functions and collision matrix elements for one-electron diatomic molecules. Computer Physics Communications, 14(1–2):121–132, March/April 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578900553.

Saraph:1970:CSR

[Sar70] H. E. Saraph. Collision strengths from reactance matrices. Computer Physics Communications, 1(4):232–240, April 1970. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465570900391.

Saraph:1972:FSC

[Sar72] H. E. Saraph. Fine structure cross sections from reactance matrices. Computer Physics Communications, 3(3):256–268, April/ May 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572900719.

Saraph:1978:FSC

[Sar78] H. E. Saraph. Fine structure cross sections from reactance matrices — a more versatile development of the program jajom. Computer Physics Communications, 15(3–4):247–258, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900954.

Saxena:1977:ANV

[Sax77a] K. M. S. Saxena. Adaptation of the new version of the reduced tensor matrix elements (AAKP) program: Inclusion of the evaluation of matrix elements of tensor products. Computer Physics Communications, 13(4):289–293, November/December 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465577900078.

Saxena:1977:NVA

[Sax77b] K. M. S. Saxena. A new version of AAKL (the matrix elements of spin-orbit interaction) adapted to spectroscopic notation. Computer Physics Communications, 13(3):193–202, September/October 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900133.

Saxena:1978:ANV

[Sax78] K. M. S. Saxena. Adaptation of the new version of the reduced matrix elements (AAKP) program: Inclusion of the checking of the input data. Computer Physics Communications, 16(1):57– 64, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578901091.

Shipman:1971:HSE

[SC71] L. L. Shipman and R. E. Christoffersen. High speed evaluation of F_{0}(x). Computer Physics Communications, 2(4):201–206, May/June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900531.

Schot:1972:CSS

[Sch72] Joanna W. Schot. Computer solutions of some transport phenomena problems. Computer Physics Communications, 4(1):73– 81, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900343.

Schofield:1973:CSS

[Sch73] P. Schofield. Computer simulation studies of the liquid state. Computer Physics Communications, 5(1):17–23, January 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900040.

Schorr:1974:PLV

[Sch74] B. Schorr. Programs for the Landau and the Vavilov distributions and the corresponding random numbers. Computer Physics Communications, 7(4):215–224, April 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900915.

Schraml:1978:LRT

[Sch78] J. Schraml. On-line and real-time processing in radio astronomy. Computer Physics Communications, 15(5):347–349, November 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900656.

Sheldon:1974:FIP

[SDH74] E. Sheldon, D. R. Donati, and H. R. Hiddleston. “MIA”, a Fortran-IV program for making spin and parity assignments to high-lying single and coherent twin nuclear levels from (alpha, nucleon) angular distributions in on-resonance, compound-nuclear, channelspin-1/2 reactions. Computer Physics Communications, 8(3):199– 219, October 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900976.

Seaton:1973:CC

[Sea73] M. J. Seaton. Close coupling. Computer Physics Communications, 6(6):247–256, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900349.

Seiler:1973:PCF

[Sei73] F. Seiler. A program calculating the formulae for polarization effects in nuclear reactions. Computer Physics Communications, 6(5):229–239, November 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900945.

Sutherland:1976:CCC

[SF76] D. C. Sutherland and W. G. Ferrier. Coulomb coefficients for complex ionic crystals. Computer Physics Communications, 12(2):179– 198, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900679.

Shayler:1978:RRF

[SF78] P. J. Shayler and M. T. C. Fang. Radial radiative flux in cylindrically symmetric arcs. Computer Physics Communications, 16(1):139–146, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578901170.

Sotona:1972:GTB

[SG72] M. Sotona and M. Gmitro. Generalized transformation brackets for the harmonic oscillator functions. Computer Physics Communications, 3(1):53–60, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900549.

Schulten:1976:REC

[SG76] K. Schulten and R. G. Gordon. Recursive evaluation of 3j and 6j coefficients. Computer Physics Communications, 11(2):269–278, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900588.

Shapiro:1970:AS

[Sha70] Joel Shapiro. Arbitrary 3n-j symbols for SU(2). Computer Physics Communications, 1(3):207–215, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557090007X.

Sheorey:1974:CEW

[She74] V. B. Sheorey. Chebyshev expansions for wave functions. Computer Physics Communications, 7(1):1–12, January 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900538.

Sheorey:1976:DCE

[She76] V. B. Sheorey. Double Chebyshev expansions for wave functions. Computer Physics Communications, 12(2):125–134, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900618.

Silver:1978:DMBa

[Sil78a] David M. Silver. Diagrammatic many-body perturbation expansion for atoms and molecules: I. General organization. Computer Physics Communications, 14(1–2):71–79, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900504.

Silver:1978:DMBb

[Sil78b] David M. Silver. Diagrammatic many-body perturbation expansion for atoms and molecules: II. Second-order and third-order ladder energies. Computer Physics Communications, 14(1–2):81–89, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900516.

Sinfailam:1970:ESC

[Sin70] A. L. Sinfailam. Electron scattering by closed shell diatomic molecules. Computer Physics Communications, 1(6):445–452, December 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900196.

Stavroulaki:1979:MCS

[SK79] B. Th. Stavroulaki and S. N. Kaplanis. Monte–Carlo solutions of the solid angle integrals for radiation detectors. Computer Physics Communications, 18(1):7–12, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900201.

Skoog:1974:NEL

[Sko74] R. Skoog. Nuclear energy loss and scattering of ions penetrating thin layers of matter. Computer Physics Communications, 7(7): 392–400, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900691.

Shaub:1978:CSP

[SLH78] W. M. Shaub, S. Lemont, and A. B. Harvey. CARS spectral profiles for homonuclear diatomic molecules. Computer Physics Communications, 16(1):73–83, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557890111X.

Savelli:1978:CCC

[SM78a] C. Savelli and M. Morando. Carla: a code to calculate the population of high spin states through compound nucleus reactions. Computer Physics Communications, 15(3–4):283–290, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557890098X.

Schindler:1978:FTF

[SM78b] Susan Schindler and R. Mirman. Functions on tableaux and frames of the symmetric group. Computer Physics Communications, 15(1–2):147–152, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900887.

Schindler:1978:GCG

[SM78c] Susan Schindler and R. Mirman. Generation of the Clebsch-Gordan coefficients for S_{n}. Computer Physics Communications, 15(1–2):131–145, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900875.

Sheldon:1971:CTD

[SMD71] E. Sheldon, S. Mathur, and D. Donati. Computation of total, differential and double-differential cross sections for compound nuclear reactions of the type (a, b), (a, bσ) and (a, bσ − σ). (III) Fortran translations of the Algol programs ‘Mandy’ and ‘Barbara’. Computer Physics Communications, 2(5):272–287, August/ September 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900038.

Smith:1970:HFN

[Smi70a] William R. Smith. Hauser–Feshbach nuclear scattering subroutine LIANA. Computer Physics Communications, 1(3):181–190, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900044.

Smith:1970:NES

[Smi70b] William R. Smith. Nuclear elastic scattering program with parameter search. Computer Physics Communications, 1(3):198– 206, January 1970. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465570900068.

Smith:1972:BRB

[Smi72] F. J. Smith. Book review: Data Structures: Theory and Practice: A. T. Berztiss, Academic Press £6.75. Computer Physics Communications, 4(1):149, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900446.

Smith:1977:BRB

[Smi77] F. J. Smith. Book review: Database systems: A practical reference: Ian R. Palmer, CACI Inc. London 1975. £22. Computer Physics Communications, 13(1):73, May 1977. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900315.

Smith:1978:STM

[Smi78] Kenneth Smith. SUBMMW: a theoretical model to predict CW sub-millimeter wave laser performance. Computer Physics Communications, 15(1–2):85–96, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900838.

Soucek:1973:ACM

[Sou73] Branko Soućek. Applications of computers and mathematical models to the study of neuronal systems nuclear and neuronal pulse spectrometry. Computer Physics Communications, 5(2):115– 122, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900271.

Sheldon:1973:CTD

[SR73] E. Sheldon and V. C. Rogers. Computation of total and differential cross section for compound nuclear reactions of the type (a, a), (a, a′), (a, b), (a, γ), (a, γ − γ), (a, bγ) and (a, bγ − γ): (IV) FORTRAN program ‘CINDY’. Computer Physics Communications, 6(3):99–131, September 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900490.

Sokolovskij:1978:CPS

[SR78] T. D. Sokolovskij and L. A. Rogoschenko. Computation of phonon spectrum from the cold neutron incoherent inelastic scattering by a polycrystal. Computer Physics Communications, 13(5–6): 381–388, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900358.

Samuel:1971:CEQ

[SS71] M. Samuel and U. Smilansky. Calculation of electric quadrupole radial matrix elements for Coulomb excitation. Computer Physics Communications, 2(7):455–469, December 1971. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465571900397.

Smend:1974:FFP

[SS74] F. Smend and M. Schumacher. Form factor program for Rayleigh scattering of gamma rays by bound electrons. Computer Physics Communications, 7(7):389–391, July 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557490068X.

Smend:1976:NRF

[SS76] F. Smend and M. Schumacher. Non-relativistic form factor program for Compton scattering of gamma rays by bound electrons. Computer Physics Communications, 11(3):363–368, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900230.

Sobolev:1973:SNI

[SSZ73] V. V. Sobolev, V. S. Synakh, and V. E. Zakharov. Some numerical investigations in nonlinear optics. Computer Physics Communications, 5(1):48–50, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557390009X.

Sharp:1972:CSA

[ST72] W. Sharp and J. C. Taylor. Computer simulation of anisotropic magnetoplasma systems in cylindrical and toroidal geometry. Computer Physics Communications, 4(1):95–99, October 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900379.

Smith:1973:PEN

[ST73] Richard L. Smith and Donald G. Truhlar. Program for evaluation of non-exchange type integrals required in electron-atom scattering theory using Slater-type orbitals as basis functions. Computer Physics Communications, 5(1):80–87, January 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900143.

Stanev:1972:SMC

[Sta72] T. Stanev. Simple Monte Carlo simulation for solving methodological problems of detecting the extensive air showers electron component. Computer Physics Communications, 4(1):47– 50, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900306.

Starkand:1976:SCM

[Sta76] Yair Starkand. Subroutine for calculation of matrix Padé approximants. Computer Physics Communications, 11(3):325–330, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900199.

Stedman:1971:LCF

[Ste71] G. E. Stedman. Lanthanide crystal field fitting routine. Computer Physics Communications, 2(4):191–200, May/June 1971. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046557190052X.

Stefanescu:1976:POC

[Ste76] A. Sabba Stefănescu. Program for optimal computation of EPR spin-Hamiltonian parameters for ions in tetragonal symmetry. Computer Physics Communications, 11(2):257–267, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900576.

Stern:1979:IEP

[Ste79] M. S. Stern. An integral equation program to calculate radial wave functions and scattering phase shifts of short-range local interactions. Computer Physics Communications, 17(4):365–374, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900997.

Strubbe:1974:MSC

[Str74] H. Strubbe. Manual for SCHOONSCHIP a CDC 6000/7000 program for symbolic evaluation of algebraic expressions. Computer Physics Communications, 8(1):1–30, August 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900812.

Strubbe:1979:DSP

[Str79] H. Strubbe. Development of the SCHOONSHIP program. Computer Physics Communications, 18(1):1–5, September 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579900195.

Sutcliffe:1972:BRB

[Sut72] L. H. Sutcliffe. Book review: NMR basic principles and progress… Grundlagen und Fortschritte: eds. P. Diehl, E. Fluck and R. Kosfeld, Vol. 6: Computer assistance in the analysis of high-resolution NMR spectra by P. Diehl, H. Kellerhals and E. Lustig, Springer-Verlag, 96 pages, DM 48, US. $15.00. Computer Physics Communications, 4(1):150, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900458.

Stanev:1979:SSS

[SV79] T. Stanev and Ch. Vankov. A set of subroutines for simulation of electron-photon cascades. Computer Physics Communications, 16(3):363–372, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900420.

Sommer:1979:NBT

[SZ79] B. Sommer and J. G. Zabolitzky. On numerical Bessel transformation. Computer Physics Communications, 16(3):383–387, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900444.

Torsti:1972:SPA

[TA72] J. J. Torsti and A. M. Aurela. Some physical applications of the solution of ill-conditioned systems of linear equations. Computer Physics Communications, 4(1):27–32, October 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900276.

Taff:1975:SRT

[Taf75a] L. M. Taff. Spying on real-time computers to improve performance. Computer Physics Communications, 10(5):262–263, November 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557590096X.

Taff:1975:TLC

[Taf75b] L. M. Taff. Two-level CAMAC handler software. Computer Physics Communications, 9(5):271–281, May 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557590079X.

Tallents:1976:CCC

[Tal76] G. J. Tallents. Collrad: a code for calculating the quasisteady state population densities of excited states of hydrogenlike ions. Computer Physics Communications, 12(2):205–212, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900692.

Tamura:1970:AMC

[Tam70] Taro Tamura. Angular momentum coupling coefficients. Computer Physics Communications, 1(5):337–342, September 1970. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465570900342.

Tamura:1971:DWB

[TCR71] T. Tamura, W. R. Coker, and F. Rybicki. Distorted wave Born approximation for nuclear reactions. Computer Physics Communications, 2(2):94–106, February/March 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900191.

Tellinghuisen:1973:FQM

[Tel73] Joel Tellinghuisen. A fast quadrature method for computing diatomic RKR potential curves. Computer Physics Communications, 6(5):221–228, November 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900933.

Thomas:1972:CPA

[Tho72] H.-C. Thomas. Computer programs in astrophysics. Computer Physics Communications, 3(S1):151–156, September 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572901233.

Thomson:1979:TPD

[Tho79] R. M. Thomson. Transport properties of dilute gas mixtures. Computer Physics Communications, 18(1):123–132, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900298.

Tamura:1974:EFR

[TL74] T. Tamura and K. S. Low. Exact finite range DWBA calculations for heavy-ion induced nuclear reactions. Computer Physics Communications, 8(5):349–376, December 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900125.

Tobin:1979:ASC

[Tob79] Frank L. Tobin. An algorithm for the storage of Condon– Shortley coefficients. Computer Physics Communications, 16(3): 283–284, March 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900353.

Tomasin:1973:CSA

[Tom73] Alberto A. Tomasin. A computer simulation of the Adriatic Sea for the study of its dynamics and for the forecasting of floods in the town of Venice. Computer Physics Communications, 5(1):51– 55, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900106.

Torrens:1973:CSR

[Tor73] Ian M. Torrens. Computer simulation in radiation damage studies. Computer Physics Communications, 5(1):32–43, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900076.

Trotman:1973:PCO

[Tro73] J. V. Trotman. A program for closed orbit minimization by analytic technique. Computer Physics Communications, 5(1):56– 63, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900118.

Thakkar:1975:ROC

[TS75a] Ajit J. Thakkar and Vedene H. Smith, Jr. Ring and other contributions to the higher virial coefficients. Computer Physics Communications, 10(2):80–85, August 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900740.

Thakkar:1975:SNE

[TS75b] Ajit J. Thakkar and Vedene H. Smith, Jr. A strategy for the numerical evaluation of Fourier sine and cosine transforms to controlled accuracy. Computer Physics Communications, 10(2):73– 79, August 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900739.

Thomas:1978:IMU

[TS78] Samuel Thomas and M. T. Sunny. Inner multiplicity of unitary groups. Computer Physics Communications, 14(3–4):267– 272, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900206.

Tscharnuter:1976:FP

[Tsc76] W. Tscharnuter. The formation of protostars. Computer Physics Communications, 12(1):1–7, September/October 1976. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465576900059.

Thomson:1976:BCS

[TSD76] R. M. Thomson, Kenneth Smith, and A. R. Davies. Boltz: a code to solve the transport equation for electron distributions and then calculate transport coefficients and vibrational excitation rates in gases with applied fields. Computer Physics Communications, 11(3):369–383, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900242.

Tran:1978:FEC

[TT78] T. M. Tran and F. Troyon. Finite element calculation of the anomalous skin effect in a homogeneous, unmagnetized cylindrical plasma column. Computer Physics Communications, 16(1):51– 56, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557890108X.

Takemasa:1979:CAM

[TTW79a] T. Takemasa, T. Tamura, and H. H. Wolter. Complex angular momentum methods for elastic scattering with an optical potential. Computer Physics Communications, 18(3):427–440, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900122.

Takemasa:1979:CFC

[TTW79b] T. Takemasa, T. Tamura, and H. H. Wolter. Coulomb functions with complex angular momenta. Computer Physics Communications, 17(4):351–355, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900973.

Tamura:1979:EFR

[TUWA79] T. Tamura, T. Udagawa, K. E. Wood, and H. Amakawa. Exact finite range DWBA form factor for heavy-ion induced nuclear reactions. Computer Physics Communications, 18(1):163– 169, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465579900341.

Tscharnuter:1979:MCS

[TW79] W. M. Tscharnuter and K.-H. Winkler. A method for computing selfgravitating gas flows with radiation. Computer Physics Communications, 18(2):171–199, October/November 1979. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465579901115.

Thomas:1973:ANP

[TYCT73] Lowell D. Thomas, Bhagat S. Yarlagadda, Gyorgy Csanak, and Howard S. Taylor. Analytical and numerical procedures in the application of many-body Green’s function methods to electron-atom scattering problems. Computer Physics Communications, 6(6):316–330, December 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465573900404.

Ugniewski:1978:RDL

[Ugn78] Stanislaw Ugniewski. Reduction of data from line, differential and surface probes in axially symmetric experiments. Computer Physics Communications, 14(3–4):155–167, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900115.

Vaccari:1974:DCG

[VD74] E. Vaccari and W. Delaney. Design criteria for a general purpose high-energy physics experiment simulation program. Computer Physics Communications, 7(3):135–144, March 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465574900034.

vandenBos:1974:PPC

[vdB74] J. van den Bos. A programming package for the calculation of cross-sections and probabilities for charge-exchange processes. Computer Physics Communications, 7(3):163–171, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557490006X.

vanderMeulen:1972:TCR

[vdM72] A. van der Meulen. Trajectory calculations for the reaction K + HBr → KBr + H in the eV-region. Computer Physics Communications, 3(1):42–52, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900537.

Virtue:1978:IPN

[VDM78] C. J. Virtue, R. J. Douglas, and B. T. A. McKee. Interactive positronfit: a new version of a program for analysing positron lifetime spectra. Computer Physics Communications, 15(1–2):97– 105, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557890084X.

Schee:1973:MPI

[vdSVSH73] E. v. d. Schee, B. P. Th. Veltman, L. Speet, and S. De Haan. Manyparticle interaction problems computed by special purpose processors, with emphasis on the Ising model. Computer Physics Communications, 5(2):104–114, February 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557390026X.

Veltman:1972:AT

[Vel72] M. Veltman. Algebraic techniques. Computer Physics Communications, 3(S1):75–78, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572901154.

VonMeerwall:1974:ASA

[VG74] E. D. Von Meerwall and M. D. Gawlik. Automatic spectrum analysis on minicomputers: Doublet resolution by Gaussian fit. Computer Physics Communications, 7(3):115–120, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465574900010.

Violino:1972:PCL

[Vio72] Paolo Violino. A program for computing level crossings and the Back–Goudsmit effect. Computer Physics Communications, 4(1):128–137, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900410.

vonMeerwall:1975:GPR

[vM75a] E. von Meerwall. A general-purpose routine for the analysis of spectroscopic peak shapes. Computer Physics Communications, 10(3):145–154, September 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900028.

vonMeerwall:1975:LSS

[vM75b] E. von Meerwall. A least-squares spectral curve fitting routine for strongly overlapping Lorentzians or Gaussians. Computer Physics Communications, 9(2):117–128, February 1975. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465575900284.

vonMeerwall:1975:FCA

[vM75c] E. D. von Meerwall. A Fortran code for automatic spectrum analysis on medium-scale computers. Computer Physics Communications, 9(6):351–359, June 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900156.

vonMeerwall:1976:FAP

[vM76a] E. von Meerwall. A flexible, all-purpose curve-fitting program. Computer Physics Communications, 11(2):211–219, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900539.

vonMeerwall:1976:SFP

[vM76b] E. D. von Meerwall. A simple Fortran program to interpret cubic X-ray powder diffraction data. Computer Physics Communications, 11(3):331–337, June/August 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465576900205.

vonMeerwall:1977:FPS

[vM77] E. D. von Meerwall. A Fortran program to simulate quadrupoledistorted NMR powder patterns. Computer Physics Communications, 13(2):107–115, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465577900364.

vonMeerwall:1978:FPR

[vM78a] E. D. von Meerwall. A Fortran program for routine analysis of magnetic susceptibility data. Computer Physics Communications, 15(3–4):237–245, October 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900942.

vonMeerwall:1978:FPC

[vM78b] E. D. von Meerwall. A Fortran program to collect histograms of microscopic scalar interactions. Computer Physics Communications, 13(5–6):421–427, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900395.

vonMeerwall:1979:APC

[vM79a] E. D. von Meerwall. An all-purpose curve-fitting program for functions of several variables. Computer Physics Communications, 18(3):411–416, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900109.

vonMeerwall:1979:FPP

[vM79b] E. D. von Meerwall. A FORTRAN program to perform signal averaging, multichannel scaling, and pulse-height analysis. Computer Physics Communications, 18(3):417–426, December 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900110.

vonMeerwall:1973:APA

[vMG73] E. von Meerwall and M. D. Gawlik. Automatic peak analysis on minicomputers. Computer Physics Communications, 5(5):309– 313, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465573900593.

Warburton:1973:BRB

[War73] A. E. A. Warburton. Book review: Padé approximants and their applications: ed. P. R. Graves-Morris, Academic Press £5.50. Computer Physics Communications, 6(2):96, August 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900568.

Warburton:1978:DVP

[War78] W. K. Warburton. DBLCON: a version of positronfit with non-Gaussian prompt for analysing position lifetime spectra. Computer Physics Communications, 13(5–6):371–379, January/ February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900346.

Wood:1974:NRS

[WB74] John H. Wood and Michael Boring. A non-relativistic SCF atomic program to compute one-electron energies, total energies, and Slater integrals. Computer Physics Communications, 7(2):73– 84, February 1974. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465574900423.

Weber:1979:ATD

[WBG79] Wim J. Weber, J. P. Boris, and J. H. Gardner. Alfven — A two-dimensional code based on shasta, solving the radiative, diffusive MHD equations. Computer Physics Communications, 16 (2):243–265, January/February 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900912.

Wang:1978:BPP

[WC78] C. S. Wang and J. Callaway. BNDPKG. A package of programs for the calculation of electronic energy bands by the LCGO method. Computer Physics Communications, 14(5–6):327–365, July/August 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900036.

Wolf:1977:CSC

[WD77] D. Wolf and K. Differt. Computer simulation of correlated self-diffusion via randomly migrating vacancies in cubic crystals. Computer Physics Communications, 13(3):167–182, September/ October 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046557790011X.

Wolf:1977:DCF

[WDM77] D. Wolf, K. Differt, and H. Mehrer. Determination of correlation factor and NMR diffusion parameters from the computer-simulated random motion of vacancies in cubic crystals. Computer Physics Communications, 13(3):183–191, September/October 1977. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465577900121.

Weber:1973:UCE

[Web73] C. Weber. The use of computers in electron- and ion-gun design. Computer Physics Communications, 5(1):44–47, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900088.

Weiner:1972:NST

[Wei72] J. H. Weiner. Numerical solution of the time-dependent Schrödinger equation. Computer Physics Communications, 4(1): 10, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900239.

Wesson:1976:NLB

[Wes76] J. A. Wesson. Non-linear behaviour of hydromagnetic instabilities. Computer Physics Communications, 12(1):53–65, September/October 1976. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465576900102.

Williams:1970:PFT

[Wil70] I. R. Williams. Program for fitting transition energies into a level scheme according to the combination principle. Computer Physics Communications, 1(6):465–467, December 1970. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465570900226.

Wills:1971:EAM

[Wil71] J. G. Wills. On the evaluation of angular momentum coupling coefficients. Computer Physics Communications, 2(6):381–382, October/November 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900300.

Wills:1972:AMC

[Wil72] R. D. Wills. Application of the Monte Carlo method to the calibration of detectors for gamma-ray astronomy. Computer Physics Communications, 4(1):51–58, October 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900318.

Wilson:1978:DMB

[Wil78] Stephen Wilson. Diagrammatic many-body perturbation expansion for atoms and molecules: III. Third-order ring energies. Computer Physics Communications, 14(1–2):91–98, March/April 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465578900528.

Wolfram:1971:KPN

[WMC71] Amy Wolfram, C. Fred Moore, and W. Rory Coker. Kinematical parameters of nuclear reactions. Computer Physics Communications, 2(7):443–448, December 1971. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465571900373.

Worlton:1972:EMM

[Wor72] T. G. Worlton. External modes of molecular crystals. Computer Physics Communications, 4(2):249–256, November 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900161.

Worlton:1973:IMR

[Wor73] T. G. Worlton. Irreducible multiplier representations. Computer Physics Communications, 6(3):149–155, September 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900519.

Wong:1973:CGF

[WP73] H. C. Wong and J. Paldus. Computer generation of Feynman diagrams for perturbation theory II. Program description. Computer Physics Communications, 6(1):9–16, July 1973. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465573900179.

Wilson:1974:FLS

[WS74] W. Wilson and L. J. Swartzendruber. A flexible least squares routine for general Mössbauer effect spectra fitting. Computer Physics Communications, 7(3):151–162, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900058.

Wilson:1979:DMB

[WS79] Stephen Wilson and David M. Silver. Diagrammatic manybody perturbation expansion for atoms and molecules: IV. Fourth-order linked diagrams involving quadruply-excited states. Computer Physics Communications, 17(1–2):47–50, April/May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900687.

Worlton:1972:GTA

[WW72] T. G. Worlton and J. L. Warren. Group-theoretical analysis of lattice vibrations. Computer Physics Communications, 3(2):88– 117, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465572900586.

Warren:1974:IVG

[WW74] J. L. Warren and T. G. Worlton. Improved version of grouptheoretical analysis of lattice dynamics. Computer Physics Communications, 8(2):71–84, September 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900575.

Yates:1971:PCR

[Yat71] A. C. Yates. A program for calculating relativistic elastic electronatom collision data. Computer Physics Communications, 2(4):175– 179, May/June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900506.

Yoshiki:1978:NAC

[Yos78] H. Yoshiki. A new approach to cross-section calculations. Computer Physics Communications, 16(1):43–49, December 1978. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465578901078.

Yurtsever:1978:IPO

[Yur78] E. Yurtsever. An integral package for one-center integrals over Slater-Transform–Preuss functions. Computer Physics Communications, 16(1):65–71, December 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578901108.

Zabusky:1973:SET

[Zab73] Norman J. Zabusky. Solitons and energy transport in nonlinear lattices. Computer Physics Communications, 5(1):1–10, January 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465573900027.

Zacharov:1972:DCS

[Zac72] B. Zacharov. Development of computer systems in physics. Computer Physics Communications, 3(S1):50–62, September 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572901130.

Zacharov:1974:BRB

[Zac74] B. Zacharov. Book review: Minicomputer interfacing. The symposium proceedings: eds. Y. Paker, G. Cain and P. Morse (Miniconsult, London, 1973), £7.00. Computer Physics Communications, 7(3):175–176, March 1974. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465574900095.

Zanella:1972:MRP

[Zan72] P. Zanella. Machine recognition of patterns in particle physics. Computer Physics Communications, 3(S1):63–74, September 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465572901142.

Zakrzewska:1978:NCM

[ZDN78] K. Zakrzewska, J. Dudek, and N. Nazarewicz. A numerical calculation of multidimensional integrals. Computer Physics Communications, 14(3–4):299–309, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465578900231.

Zhidkov:1972:NMM

[Zhi72] E. P. Zhidkov. New mathematical methods for analysing some nonlinear physics problems with the aid of a computer. Computer Physics Communications, 4(2):182–185, November 1972. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465572900057.

Zaikin:1979:PCT

[ZKU79] P. N. Zaikin, V. G. Kritsky, and M. V. Ufimtcev. The problem of comparison of two spectra. Computer Physics Communications, 18(3):327–329, December 1979. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557990002X.

Zlokazov:1978:USO

[Zlo78] V. B. Zlokazov. UPEAK — spectro-oriented routine for mixture decomposition. Computer Physics Communications, 13(5– 6):389–398, January/February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557890036X.

Zlokazov:1979:DPA

[Zlo79] V. B. Zlokazov. Domus — a program for the analysis of twodimensional spectra. Computer Physics Communications, 18(2): 281–286, October/November 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046557990119X.

Zohni:1972:FPC

[Zoh72] O. Zohni. A FORTRAN program for the computation of the generalized Talmi coefficients. Computer Physics Communications, 3(1):61–68, January/February 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465572900550.

Zurek:1977:PRM

[ZS77] Wojciech H. Zurek and William C. Schieve. Pattern recognition in molecular dynamics. Computer Physics Communications, 13(2): 75–79, July 1977. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465577900327.